### v1

# A Survey of LLM-based Automated Program Repair: Taxonomies, Design Paradigms, and Applications

Abstract.
Large language models (LLMs) are reshaping automated program repair (APR). We categorize the recent 63 LLM-based APR systems published from January 2022 to June 2025 into four paradigms, and show how retrieval- or analysis-augmented contexts strengthen any of them. This taxonomy clarifies key trade-offs: fine-tuning delivers strong task alignment at high training cost; prompting enables rapid deployment but is limited by prompt design and context windows; procedural pipelines offer reproducible control with moderate overhead; agentic frameworks tackle multi-hunk or cross-file bugs at the price of increased latency and complexity. Persistent challenges include verifying semantic correctness beyond test suites, repairing repository-scale defects, and lowering the costs of LLMs. We outline research directions that combine lightweight human feedback, repository-aware retrieval, code analysis, and cost-aware planning to advance reliable and efficient LLM-based APR.
*Corresponding author.
copyright: none
1. Introduction
Figure 1. Taxonomies of LLM-based APR
LLM-based APR has become a promising research direction in software engineering (Zhang et al., 2024a; Huang et al., 2024a; Zhang et al., 2023b; Fan et al., 2023b). Recent code-focused LLMs, such as OpenAI’s Codex (Chen et al., 2021), CodeLlama (Roziere et al., 2023), and DeepSeek-Coder (Guo et al., 2024; Zhu et al., 2024), have demonstrated capabilities to understand and generate code, enabling them to fix bugs in programs with minimal human input. Early studies have shown that LLM-based APR techniques outperform traditional repair methods on repair benchmarks, achieving state-of-the-art program repair success rates (Sobania et al., 2023; Prenner et al., 2022; Xia and Zhang, 2024; Yang et al., 2024c; Jiang et al., 2023a). These breakthroughs have drawn intense interest from both researchers and practitioners, transforming what was once a niche topic into a vibrant field (Le Goues et al., 2025; Zubair et al., 2024). Indeed, new LLM-driven repair approaches are appearing at a remarkable pace, reflecting a broad optimism that this paradigm can significantly improve automated debugging and maintenance in practice (Liu et al., 2024a; Li et al., 2024a, 2025).
Surveys conducted before the LLM era provided a taxonomy of classical APR techniques but naturally did not include approaches based on LLMs (Le Goues et al., 2019). More recent surveys diverge in scope: Huang et al. categorize repair into search, constraint, template, and learning while relegating LLMs to future work (Huang et al., 2023b); Winter et al. focus on developer perceptions rather than technical structure (Winter et al., 2022); Zhang et al. adopt a localization-generation-ranking-validation pipeline view (Zhang et al., 2023a); Haque et al. review tuning strategies but not full workflows (Haque et al., 2025a). Several niche perspectives have emerged: He et al. analyze the LLM-as-judge role for patch assessment rather than generation (He et al., 2025); Zhou et al. take a security-centric vulnerability taxonomy survey but exclude general program repair systems and newly agentic systems (Zhou et al., 2024a); Dikici et al. group forty-one tools under template, machine-learning, and deep-learning labels, implicitly folding GPT-style repair into a broad deep-learning bin (Dikici and Bilgin, 2025); Anand et al. survey LLM-based APR but cover only a small set of papers and propose no dedicated taxonomy (Anand et al., 2024). Overall, these surveys do not provide a unified view that encompasses prompt-driven, retrieval-augmented, procedural, and fully agentic repair workflows, nor do they address evaluation issues such as data leakage, weak oracles, or inconsistent success criteria. Without such a synthesis, practitioners cannot easily compare the strengths of competing strategies or identify shared research gaps.
To address this need, we present a comprehensive survey that maps the design space across the four paradigms (fine-tuning, prompting, procedural pipelines, and agentic frameworks, as shown in Figure 1), incorporates cross-cutting Retrieval Augmented Generation (RAG) and Analysis Augmented Generation (AAG) enhancements, and highlights challenges in methodology and evaluation (surveyed papers: Jan 2022 - June 2025). As Figure 1 shows, parameters versus control flow define four distinct paradigms; the subsequent sections unpack their trade-offs. Notably, we treat certain techniques as cross-cutting enhancements rather than separate categories in this taxonomy. In particular, Retrieval-Augmented Generation (RAG) refers to augmenting the LLM’s input with external knowledge (such as relevant code snippets, documentation, or historical fixes) (Lewis et al., 2021; Mansur et al., 2024; Yang et al., 2025a). Analysis-Augmented Generation (AAG) involves incorporating program static or dynamic analysis results (such as error traces or data-flow facts) into the prompt or constraining the model’s output (Dolcetti et al., 2025; Xiao et al., 2025; Li et al., 2024a; Jain et al., 2023b). Figure 2 juxtaposes the four paradigms in parallel columns and overlays Retrieval- and Analysis-Augmented contexts across the top bar, making explicit both the sequential flow within each paradigm and the cross-cutting enhancements that any paradigm can adopt.
Table 1. Landscape of LLM-based APR by sub-paradigm. Columns specify the Training objective, whether generation is Single or Multiple, who Control the next step of the workflow, and the Representative Systems (scope: Jan 2022 to June 2025).
Sub-paradigm Training Generation Control Representative Systems
Fine-Tuning
Full FT fine-tuning single scripted by human Huang et al. (2023), Jiang et al. (2023), VulMaster (2023), RepairCAT (2024)
PEFT fine-tuning single scripted by human RepairLLaMA (2023), MORepair (2024), Luo et al. (2024), Li et al. (2024), Ruiz et al. (2025)
Knowledge Distilling fine-tuning single scripted by human KNOD (2023), DistiLRR (2024), NARRepair (2024)
RLFT fine-tuning single scripted by human RePair (2024), SecRepair (2024), SWE-RL (2025), AdaPatcher (2025), Smart-LLaMA-DPO (2025)
Context-enriched FT fine-tuning single scripted by human TraceFixer (2023), InferFix (2023), PyTy (2024), NTR (2024)
Prompting
Zero-shot none single scripted by human AlphaRepair (2022), Prenner et al. (2022), Fan et al. (2023), Tian et al. (2023)
Few-shot none single scripted by human Xia et al. (2023), Gao et al. (2023), Ahmed et al. (2023), CEDAR (2023)
+ RAG none single scripted by human T-RAP (2024), RLCE (2024), DSrepair (2025)
+ AAG none single scripted by human Appatch (2024), D4C (2024), TracePrompt (2025)
Procedural
Test-in-the-Loop none multiple scripted by human ChatRepair (2024), ThinkRepair (2024), REx (2024), ContrastRepair (2024)
Human-in-the-Loop none multiple scripted by human CREF (2024), HULA (2024), DRCodePilot (2024)
+ RAG none multiple scripted by human Agentless (2024), RAGFix (2024), KGCompass (2025)
+ AAG none multiple scripted by human Repilot (2023), FormHe (2025), PredicateFix (2025), LLM4CVE (2025)
Agentic
Tool-Augmented Agents none multiple decided by LLM SWE-Agent (2024), AutoCodeRover (2024), RepairAgent (2024), LANTERN (2025), VulDebugger (2025), SWE-Agent M (2025), OpenHands (2025)
LLM-as-Judges none multiple decided by LLM VulSifter (2024), LLM4PatchCorrect (2024), Execution-free Critic (2025), PatchDiff (2025)
Self-Controlled System none multiple decided by LLM MAGIS (2024), SWE-Search (2024), Learn-by-Interact (2025)
Figure 2. Overview of LLM-based program repair techniques
Table 2. Annual distribution of LLM-based APR applications by paradigm
Year Fine-Tuning Prompting Procedural Agentic Total
2022 0 2 0 0 2
2023 7 6 1 0 14
2024 10 4 9 7 30
2025 (Jan - June) 4 2 4 7 17
Total 21 14 14 14 63
Our survey distills several insights into the trade-offs among the four LLM-based APR paradigms. Fine-tuning aligns a model tightly with the repair task, producing highly accurate fixes for defects similar to those in its training data; however, it demands large bug-fix corpora and considerable GPU time. Prompting requires no additional training and supports rapid deployment, yet its effectiveness hinges on careful prompt engineering and is bounded by the model’s fixed knowledge and context window. Both paradigms can be strengthened with retrieval-augmented or analysis-augmented prompts that inject external knowledge or program-analysis feedback, thereby improving semantic correctness and reducing invalid patches. Procedural pipelines embed one or more LLM calls in deterministic multi-step workflows, offering reproducible control at moderate runtime cost. Agentic frameworks delegate planning and tool invocation to the LLM itself, enabling multi-hunk and cross-file repair at the price of increased engineering complexity, latency, and resource consumption. This community shift from single-turn generation to iterative, tool-assisted workflows reflects a drive to boost success rates on complex defects. Nonetheless, three challenges remain: verifying semantic correctness beyond limited test suites, repairing bugs that span multiple modules or exceed context limits, and mitigating the substantial computational cost of large-model pipelines. Addressing these issues is crucial for making LLM-based APR reliable and efficient in continuous integration settings.
Our contributions.
  1. (i)
    We present the first unified taxonomy that covers four repair paradigms and two augmentation layers for LLM-based APR.
  2. (ii)
    We review 63 recent state-of-the-art APR systems published between Jan 2022 and June 2025, summarizing their design trade-offs, common trends, and open questions.
  3. (iii)
    We maintain a living bibliography at https://github.com/GLEAM-Lab/ProgramRepair.
2. Fine-Tuning Approaches
Fine-tuning adapts a pre-trained LLM for APR by updating its weights using bug-fix data. Research in this area now revolves around five distinct strategies. Fully supervised fine-tuning updates every parameter of the backbone and achieves strong accuracy across diverse benchmarks (Jiang et al., 2023a; Zhou et al., 2024b; Jiang and Wu, 2024). Parameter-efficient fine-tuning freezes most of the backbone and learns small adapter modules, which lowers memory cost while keeping competitive repair performance (Silva et al., 2023; Yang et al., 2025b; Luo et al., 2024; Li et al., 2024c; Ruiz et al., 2025). Reinforcement learning fine-tuning trains the model with execution or security rewards so that it optimizes for functional correctness instead of log likelihood (Zhao et al., 2024a; Islam et al., 2024; Wei et al., 2025; Dai et al., 2025; Yu et al., 2025a). Knowledge distillation transfers bug-fixing skill from a large teacher or rule set to a smaller student, cutting inference latency without much accuracy loss (Jiang et al., 2023b; Wong et al., 2024; Yang et al., 2024d). Context-enriched fine-tuning adds static-analysis results, execution traces, or retrieved patches to the input so the model can reason beyond the local snippet (Bouzenia et al., 2023; Jin et al., 2023; Chow et al., 2024; Huang et al., 2024b).
2.1. Full Fine-Tuning
With the emergence of code-target language models based on transformers (Vaswani et al., 2017) such as GPT-2 (Radford et al., 2019), CodeT5 (Wang et al., 2021), and CodeLlama (Roziere et al., 2023), researchers have pursued full-model supervised fine-tuning on program repair scenarios. Recent work explores several variants of this strategy. Large-scale benchmark adaptation fine-tunes multiple backbone LLMs across standard datasets to quantify the effects of size and data (Jiang et al., 2023a). Security-specific pipelines inject vulnerability patterns and CWE knowledge during full tuning to target secure patching (Zhou et al., 2024b). Domain-focused fine-tuning demonstrates that even 1B-parameter LLMs can learn high-yield repair heuristics when specialized data are available (Jiang and Wu, 2024). These efforts illustrate the breadth of full-parameter supervised fine-tuning for APR.
Huang et al. conducted an empirical study on fine-tuning LLMs (including CodeBERT (Feng et al., 2020), GraphCodeBERT (Guo et al., 2021), PLBART (Ahmad et al., 2021), CodeT5 (Wang et al., 2021), and UniXcoder (Guo et al., 2022)) for program repair tasks (Huang et al., 2023a). Experimental results show that LLMs in the fine-tuning paradigm can significantly outperform previous state-of-the-art APR tools on both single-hunk and multi-hunk bugs within Defects4J (Just et al., 2014). However, they reveal that LLMs usually fail to generate correct patches due to the input/output length limit of LLMs. On the other hand, the generated candidate patches are significantly limited, as hardware resources often fail to meet the high computational demands of LLMs for massive candidate patches.
Jiang et al. conducted a large-scale study that fully fine-tunes ten pre-trained code language models across four APR benchmarks to quantify the effect of supervised adaptation on repair performance (Jiang et al., 2023a). Fine-tuning raised fix rates by up to an order of magnitude and showed a clear positive correlation between model size and patch success. The work also revealed heavy GPU demand and frequent overfitting to the suspected bug location, highlighting concerns about efficiency and generalization.
VulMaster fully fine-tunes a CodeT5 (Wang et al., 2021) backbone on a security-focused corpus that blends complete vulnerable functions, abstract-syntax-tree structure, and CWE knowledge, then enlists ChatGPT at inference time to supply any missing contextual hints (Zhou et al., 2024b). On 1,754 real C/C++ projects with 5800 vulnerable functions, it doubles end-to-end effectiveness: exact-match rises from 10.2% to 20.0%, BLEU from 21.3% to 29.3%, and CodeBLEU from 32.5% to 40.9%. Strengths are its multi-source inputs and multi-LLM collaboration, but training demands $\sim$30 GB of GPU memory, and its gains taper on ultra-long files despite the enlarged context window.
RepairCAT is a fully fine-tuned 1B-parameter StarCoder (Li et al., 2023; Lozhkov et al., 2024), trained on faults generated by AI code synthesizers to learn characteristic correction patterns (Jiang and Wu, 2024). Targeted training addresses a significant portion of such failures at a modest computational cost, demonstrating that small open models can benefit from full-model supervision in a specialized domain. Limited capacity and a narrow data distribution, however, curtail its effectiveness in addressing complex or conventional software defects.
Collectively, these methods demonstrate that full-parameter supervised fine-tuning can achieve state-of-the-art repair accuracy while markedly reducing reliance on elaborate prompt engineering. Yet, practical challenges persist: high computational cost and, as large-scale benchmark studies indicate, a risk of catastrophic forgetting, dependence on the quality of auxiliary data, and susceptibility to overfitting when training data or evaluation suites are narrow. Tackling these issues remains essential for deploying large-scale, fine-tuned repair models in continuous integration environments.
2.2. Parameter-Efficient Fine-Tuning
Parameter-efficient fine-tuning (PEFT) alleviates the memory and compute burden of adapting LLMs for program repair by updating only a compact set of task-specific parameters. Current PEFT strategies for repair fall into three families: (i) structural adapters, including adapter layers (Houlsby et al., 2019), prefix vectors (Lester et al., 2021), and low-rank matrices such as LoRA (Hu et al., 2021); (ii) quantization-aware adapters typified by QLoRA, which merges 4-bit quantization with LoRA (Dettmers et al., 2023); and (iii) noise-regularized tuning exemplified by NEFTune, which injects calibrated embedding noise (Jain et al., 2023a). These designs shrink trainable parameters by up to three orders of magnitude, enable single-GPU training, and keep the frozen backbone reusable for other tasks. Efficiency-oriented variants demonstrate promising effectiveness in program and vulnerability repair, broadening PEFT’s applicability to larger models and consumer-grade hardware (Silva et al., 2023; Yang et al., 2025b; Luo et al., 2024). Nevertheless, attaining strong repair accuracy still requires judicious adapter placement, appropriate rank selection, and high-quality bug-fix data, because limited adapter capacity can miss semantic links that span multiple files (Li et al., 2024c; Ruiz et al., 2025).
RepairLLaMA augments a frozen CodeLlama-7B with rank-8 LoRA matrices (0.06% of parameters) and fine-tunes them on 4,000 curated bug-fix pairs (Silva et al., 2023). Training on a single GPU fixes 144 Defects4J-V2 (Just et al., 2014) bugs, 109 HumanEval-Java tasks (Jiang et al., 2023a), and 20 GitBug-Java issues (Silva et al., 2024), outperforming larger frozen models. Limited adapter capacity and reliance on static fault localization reduce accuracy for multi-hunk or cross-file defects.
MORepair applies QLoRA-style 4-bit quantization and rank-32 LoRA adapters, augmented with NEFTune noise, to multiple CodeLlama and other LLMs’ backbones while jointly optimizing patch correctness and repair rationale (Yang et al., 2025b). This configuration fixes the most bugs compared to standard fine-tuning on EvalRepair-Java and EvalRepair-C++, and in-depth case studies indicate the accuracy of both fixed code and generated rationale. However, performance still depends on rationale-annotated data, where the quality of rationale may influence the effectiveness of MORepair.
Luo et al. apply LoRA adapters in a federated setting where each organization trains locally on proprietary code and a server aggregates updates, then evaluates on EvalRepair-Java (Luo et al., 2024). The protocol preserves privacy and matches centralized training within two percentage points while exchanging only megabytes of adapter weights per round. Convergence slows under highly heterogeneous client data, and a secure aggregation infrastructure is required.
Li et al. benchmark LoRA, prefix-tuning, adapters, and (IA)3 (Liu et al., 2022) on the APR-INSTRUCTION corpus with CodeT5+ and StarCoder backbones (Li et al., 2024c). The strongest adapter (IA)3 resolves 58% more bugs than earlier LLM baselines while keeping memory usage low, but its gains plateau once adapter size or data volume exceeds moderate thresholds, so careful hyper-parameter tuning remains essential.
Ruiz et al. lightly apply LoRA to DeepSeek-Coder-Instruct (Guo et al., 2024) and CodeLlama-Instruct (Roziere et al., 2023), then adopt an iterative generate-test-refine loop capped at ten patches per bug (Ruiz et al., 2025). With only 1000 training examples and two refinement rounds, this strategy boosts plausible fixes by 78% on HumanEval-Java. It narrows much of the gap to GPT-4, although the extra iterations add runtime overhead and yield diminishing returns on simpler bugs or very large datasets.
Overall, existing evidence suggests that PEFT can recover a substantial share of the repair accuracy reported for full fine-tuning in many settings, while using only a fraction of the computational and storage costs required for full fine-tuning. Effective practice requires balancing adapter capacity with data diversity, enriching supervision signals (for example, federated or rationale-aware training), and pairing lightweight models with search or analysis feedback to address multi-file and semantic defects.
2.3. Knowledge Distilling
Knowledge distillation transfers bug-fixing skills from a high-capacity teacher or knowledge source to a smaller student, allowing inference costs to drop without compromising accuracy. Recent studies have applied classical teacher-student distillation (including soft-target and rationale distillation) (Wong et al., 2024), dataset distillation (Yang et al., 2024d), and rule-guided distillation (Jiang et al., 2023b) to APR, yielding strong empirical gains. These techniques either compress large teachers or replace them with structured knowledge, inject structural hints, restrict generation using structured rules, or enforce domain constraints while keeping the student lightweight. The following paragraphs review three representative methods that illustrate the breadth of distillation for program repair.
KNOD distills formal syntax and semantic rules into a student decoder by aligning its token probabilities with a rule-derived teacher distribution (Jiang et al., 2023b). The method repaired 72 Defects4J v1.2 bugs, 25 QuixBugs, and 50 Defects4J v2.0 bugs, establishing new benchmark records. By restricting token choices to grammar-compliant options, KNOD greatly improves compile success and patch validity. Rule coverage is limited to statically checkable properties and offers less help for logical faults that do not violate syntax or types.
DistiLRR applies teacher-student distillation in which GPT-3.5 or GPT-4 produces bug fixes and natural-language rationales that supervise a CodeLlama-7B student (Wong et al., 2024). The distilled student nearly doubled the pass@1 success rate on low-resource languages such as Perl and Golang and matched the teacher on high-resource languages. Its main strength is the use of rich rationales, which give the student detailed guidance and boost repair in under-represented domains. The approach depends on high-quality synthetic rationales and yields smaller gains when the student already has ample native training data.
NARRepair is a non-autoregressive APR model that first distills a CodeT5-large (Wang et al., 2021) teacher to build a cleaner student training set, then generates patches in parallel using action guidance and AST dependencies (Yang et al., 2024d). It repairs 69, 45, and 23 bugs on Defects4J (Just et al., 2014) v1.2, v2.0, and QuixBugs (Lin et al., 2017) while delivering 5.4-18.6 times faster inference than autoregressive baselines. The combination of dataset distillation and two-stage decoding bridges the usual quality gap of non-autoregressive generation without sacrificing speed. However, all evaluations are Java-only, so cross-language generality is unverified.
Distillation thus offers a practical route to simplify repair models while retaining high accuracy, and, in non-autoregressive settings, can markedly shorten decoding time. Classical teacher-student, dataset distillation, and rule-guided strategies each mitigate specific weaknesses of direct fine-tuning. Together, they show that transferring behavior, structural priors, or domain rules can deliver large gains without the runtime cost of very large models. Future work could blend these signals, including rationales and dataset curation, and extend distillation to reasoning traces, allowing lightweight students to tackle complex multi-file defects.
2.4. Reinforcement Learning Fine-Tuning
Recently, policy-gradient algorithms have been widely used in LLM-based research, including PPO (Schulman et al., 2017), DAPO (Yu et al., 2025b), and Group Relative Policy Optimization GRPO (Shao et al., 2024). In this section, we also include preference-based tuning, such as DPO (Rafailov et al., 2023), which is not strictly an RL algorithm. Recent RL-based systems include RePair (Zhao et al., 2024a), SecRepair (Islam et al., 2024), and SWE-RL (Wei et al., 2025). A preference-based alternative, AdaPatcher (Dai et al., 2025) and Smart-LLaMA-DPO (Yu et al., 2025a), adopt Direct Preference Optimization (DPO) to refine their patch policy. In program-repair settings, the reward most often combines execution signals, such as compile-and-test success, with static cues like vulnerability removal or diff similarity to reduce sparsity (Shojaee et al., 2023; Ye et al., 2025). Some recent works adopt a two-stage pipeline that performs supervised fine-tuning to obtain a stable repair model and then applies RLFT for further alignment (Dai et al., 2025). Compared with purely supervised fine-tuning, RLFT can push the model beyond the training distribution and align it directly with functional correctness; however, it incurs heavy computation because each gradient update may require running tests, rewards are often sparse, and poorly shaped rewards can lead to over-fitting weak test oracles (Dou et al., 2024; Liu et al., 2023; Ye et al., 2025).
RePair starts from a fully supervised StarCoder-15B model and adds an iterative RL loop miming human debugging (Zhao et al., 2024a). A smaller critic model, trained to distinguish good from bad patches based on compile and test outcomes, supplies rewards to the actor via PPO. The actor may propose several successive fixes for one bug, receiving intermediate feedback at each step. This process enhances repair success on competitive-programming bugs to a level comparable to that of much larger proprietary LLMs. The trade-off is higher latency and the need to guard against infinite repair loops on hard-to-fix defects.
SecRepair targets vulnerability repair. It slices the code around each flaw, then fine-tunes a CodeGen2 (Nijkamp et al., 2023) actor with PPO where the reward combines unit-test outcomes, semantic-similarity checks, and a security bonus that favors patches eliminating unsafe patterns (Islam et al., 2024). Trained on their built 36k-example InstructVul dataset and evaluated on zero- and N-day bugs from 6 open-source IoT operating systems, SecRepair fixes 58% of vulnerable functions and generates concise commit messages that describe the flaw, outperforming a supervised baseline by up to 22 percentage points F1 while adding only 9% runtime overhead. Strengths include security-aware rewards and explicit vulnerability descriptions; limitations are reliance on known flaw locations and increased token usage for descriptive comments.
SWE-RL fine-tunes a 70B-parameter Llama 3 backbone using policy-gradient RL on a large corpus of real software-evolution traces (GitHub issues, pull requests, and their ground-truth patches) (Wei et al., 2025). Its reward is a patch-level diff-similarity score, which eliminates costly execution while still offering fine-grained textual feedback. Training on authentic pull-request edits enables the model to capture developer reasoning, achieving a 41.0% solve rate on SWE-bench Verified (Jimenez et al., 2024), state-of-the-art among open-source LLMs under 100B, and transferring positively to five out-of-domain reasoning benchmarks. The approach scales to genuine project history at low cost, although its rule-based reward may undervalue semantically correct but stylistically different patches.
AdaPatcher adopts a strict two‑stage pipeline: it first performs supervised fine‑tuning on CodeLlama‑Instruct‑7B to train a Bug Locator that pinpoints faulty lines, then applies reinforcement learning, in the form of Direct Preference Optimization (DPO) (Rafailov et al., 2023), to a Program Modifier that refines the patch policy using preference pairs derived from unit‑test outcomes and a “fewer‑changes” reward, achieving 62% consistency and 67.57% accuracy on their newly released ACPR benchmark, a clear lead over purely supervised or prompt‑based baselines (Dai et al., 2025). Its strength lies in marrying location‑aware supervised fine-tuning with preference‑based RLFT: the first stage supplies reliable edit anchors, and the second stage’s lightweight reward (diff similarity + compile‑and‑test bonus) lets the model learn to generate minimal yet correct edits without the high runtime of full Monte‑Carlo execution.
Smart-LLaMA-DPO targets smart-contract vulnerability detection and explanation through a three-stage post-training pipeline on a LLaMA-3.1-8B backbone (Yu et al., 2025a). First, a continual pre-training (CPT) phase feeds millions of curated Solidity contracts to solidify domain knowledge. Second, supervised fine-tuning (SFT) optimizes a dual-objective loss that balances detection logits and explanation likelihoods on a hand-verified dataset covering four major vulnerability types and seven machine-unauditable (MU) categories. Finally, the model undergoes Direct Preference Optimization (DPO) (Rafailov et al., 2023), a lightweight, reward-free RL method that updates policy logits to favor expert-preferred explanations over suboptimal ones, using paired outputs crafted by security specialists. On the same evaluation suite, Smart-LLaMA-DPO raises F1 by 7.5%-11.1% and accuracy by up to 10.5% over the previous SOTA DMT (Qian et al., 2023), and beats iAudit (Ma et al., 2024) on seven MU vulnerability classes by +26% F1 and +18% accuracy, while human judges mark 81%, 84%, and 95% of its explanations as correct, thorough, and clear, respectively. However, its current scope is limited to Solidity and single-file contracts, and scaling the preference dataset remains an open challenge.
Existing RLFT studies confirm that execution- and security-based rewards can steer language models toward semantically correct patches, yet several obstacles remain. Many current methods still depend, to varying extents, on dynamic test execution, which can significantly increase training costs when no proxy reward is available. Promising directions include hybrid reward functions that combine static cues with selective dynamic tests to reduce runtime, curriculum schedules to alleviate sparsity, and hierarchical planner-critic architectures that break down repair into tractable sub-goals. Resolving these challenges will be essential before RL-enhanced repair models can be reliably integrated into continuous integration pipelines.
2.5. Context-Enriched Fine-Tuning
Context-enriched fine-tuning augments the input with auxiliary signals such as static-analysis reports, runtime traces, inter-procedural dependency graphs, and historical patches so the model can reason beyond the local snippet (Bouzenia et al., 2023; Jin et al., 2023; Chow et al., 2024; Huang et al., 2024b). Approaches along this line inject compiler diagnostics, execution stack traces, or retrieved historical patches into the encoder input, guiding the model toward the root cause and constraining its edits (Jin et al., 2023). The main challenge is scalability: computing static or dynamic context for every training instance is costly, and long sequences may exceed LLMs’ context limits (Tian et al., 2023). Additionally, noisy or irrelevant signals can confuse the model, so effective selection and encoding of context remain open problems (Parasaram et al., 2024; Holloway et al., 2024).
TraceFixer demonstrates the value of dynamic feedback by feeding a pre-trained CodeT5 model with partial execution traces that capture the concrete program state where behavior diverges from expectations (Bouzenia et al., 2023). The trace tokens are placed in a dedicated channel during fine-tuning, allowing the decoder to align state anomalies with the edits needed to repair them. This design repairs $13-20\%$ more bugs than a code-only baseline and excels on defects whose symptoms manifest only at runtime.
InferFix couples Meta’s Infer static analyzer with retrieval-augmented fine-tuning (Jin et al., 2023). For every alert, it retrieves analogous historical patches from a large corpus, concatenates the alert description, the buggy code, and the retrieved examples, and fine-tunes a 1.3B parameter transformer to emit a patch that clears the analyzer warning. The additional context leads to 59 unique repairs that baseline LLMs miss and reduces false positives by filtering candidates that still trigger Infer.
PyTy targets gradual-typing errors in Python by injecting the exact mypy diagnostic and nearby type annotations into the input sequence before fine-tuning a CodeT5 backbone (Chow et al., 2024). Cross-language pretraining on Java supplies a rich inductive bias, and the explicit type-error signal steers the model toward minimal, type-consistent edits. On their built PyTyDefects benchmark, the approach corrects 54.4% of real type errors and doubles the success rate of GPT-4 prompted without diagnostics.
NTR (Neural Template Repair) formalizes template-based repair as a conditional text-to-text task where the selected template identifier is injected as a control token (Huang et al., 2024b). A lightweight classifier first ranks candidate templates, then a billion-parameter decoder is fine-tuned to translate the buggy code and template hint into a concrete patch. An OtherTemplate token lets the model learn fixes beyond the known catalogs. NTR with StarCoder repairs 128 bugs on Defects4J-v2 and 129 on HumanEval-Java, surpassing the best prior work while remaining template-aware.
Overall, these recent studies confirm that richer semantic, dynamic, historical, and inter-procedural dependency graph contexts can materially improve LLM-based repair. Yet the cost of gathering and encoding such information remains significant, and ever-longer inputs challenge current context windows. Future work will need more selective retrieval, adaptive compression, and context-aware parameter-efficient fine-tuning strategies to retain the benefits of additional signals without compromising scalability or robustness.
2.6. Discussion and Outlook
Research on fine-tuning for APR reveals that full supervised adaptation yields the strongest patch accuracy, as it embeds rich code knowledge. However, this benefit comes at a high GPU cost and the potential erosion of broader coding abilities. Parameter-efficient tuning recovers much of that accuracy by updating small adapter modules and low-rank matrices. Still, its effectiveness hinges on sensible adapter design and diverse, well-curated training data. Reinforcement learning introduces execution rewards that guide models toward functional correctness, yet sparse signals and repeated test runs extend training time. Supplying static analysis, execution traces, or retrieved historical patches broadens the LLM’s perspective. It helps resolve multi-file or semantic defects, though collecting and encoding such context is expensive and can exceed current input limits. The most promising direction is to combine these strengths: coupling lightweight adapters with selective retrieval of high-value context, applying reinforcement learning only when tests provide clear feedback, and adopting staged curricula with locate-edit-verify workflows to shorten feedback loops. Privacy-preserving federated or continual updates can keep models current without exposing proprietary code. Advances across model architecture, training signals, and deployment practices will be needed before fine-tuned repair systems can deliver reliable, low-latency fixes in everyday continuous integration environments.
3. Prompting Approaches
Prompt-based APR treats the LLM as a frozen component that returns a complete patch after a single query, while every step of retrieval, example curation, or validation runs outside the model. Four practical prompt styles emerge, distinguished by the amount and source of extra context they add. Zero-shot prompting supplies only the buggy code and a brief fix instruction, counting on the LLM’s prior knowledge to generate a patch (Xia and Zhang, 2022; Prenner et al., 2022; Tian et al., 2023). Few-shot prompting precedes the defect with a handful of similar bug-fix pairs, enabling the model to imitate their edit pattern (Xia et al., 2023; Gao et al., 2023; Ahmed and Devanbu, 2023; Nashid et al., 2023). Retrieval-augmented prompting attaches code fragments, templates, or knowledge-graph facts fetched from external resources, grounding the patch in concrete project or domain information (Chen et al., 2024; Liu et al., 2024b; Ouyang et al., 2025). Analysis-augmented prompting injects diagnostics derived from static or dynamic analysis, such as failing-test logs or semantics-aware slices, to guide the model toward the root cause (Xu et al., 2024; Nong et al., 2024; Haque et al., 2025b). These approaches share the same single-turn interaction but trade off preparation effort, token budget, and repair accuracy in different ways.
3.1. Zero-shot Prompting
Zero-shot prompting keeps a pre-trained model frozen and asks it to generate a patch in a single turn by providing only the buggy code and a terse instruction such as “fix the bug”. Because the approach omits fine-tuning, retrieval, and multi-turn dialogue, deployment is rapid and incurs no additional training cost. AlphaRepair (Xia and Zhang, 2022) was the first study to demonstrate that LLMs can outperform state-of-the-art APR techniques with zero-shot. Empirical studies such as Prenner et al. (Prenner et al., 2022) and Fan et al. (Fan et al., 2023a) on Codex and Tian et al. (Tian et al., 2023) on ChatGPT, show that careful wording plus a sufficiently large model can already match or sometimes surpass several traditional repair pipelines. At the same time, they expose strong prompt sensitivity, but their robustness beyond single-hunk functions remains unclear.
AlphaRepair was the first LLM-based APR approach that performs repair in a zero-shot setting by exploiting the masked-language-modeling capability of CodeBERT without any task-specific fine-tuning (Xia and Zhang, 2022). AlphaRepair masks the tokens of the buggy lines and leverages CodeBERT to infill the correct code, thereby predicting the fixed codes directly instead of learning edit scripts. On Defects4J-v1.2/2.0, AlphaRepair repairs up to 3.3 times more bugs than the best baselines, and it also outperforms other neural and heuristic tools on QuixBugs, demonstrating that a frozen pre-trained LLM can surpass state-of-the-art APR techniques without further training.
Fan et al. conducted a large-scale evaluation of zero-shot Codex on competitive-programming tasks and then studied how APR can salvage its failures (Fan et al., 2023a). Their prompt contained only the LeetCode problem description and the target function’s signature; no exemplar fixes were provided. By generating 50 solutions per task, Codex solved 46 of the 113 tasks in their new LMDefects benchmark, while subsequent APR passes (TBar, Recoder, Codex-edit) raised the total by a further 9 tasks, showing that LLM with naive zero-shot prompting already covers a substantial fraction of real bugs and that lightweight repair can extend this reach.
Prenner et al. evaluated Codex (12B) (Chen et al., 2021) on all 40 QuixBugs (Lin et al., 2017) bugs in Python and Java (Prenner et al., 2022). They tried five one-shot prompt variants that differed only in the static text preceding the buggy function. With temperature 0 and max_tokens=1024, Codex generated ten candidates per bug; a patch was accepted if QuixBugs’ reference tests all passed. The best prompt repaired 23 Python and 14 Java bugs, outperforming three specialized neural APR systems despite Codex never having seen any examples. The study also showed that adding a simple “### bug is here” hint sometimes hurts accuracy, illustrating how small prompt edits can swing outcomes.
Tian et al. explored ChatGPT (GPT-3.5) as a programming assistant across code generation, summarization, and program repair (Tian et al., 2023). For repair, they built a benchmark of 1,783 incorrect student assignments covering five classic problems. A single-shot prompt that contained only the buggy function (or that function plus its textual assignment description) asked ChatGPT for a “corrected version”. Without any retries or analysis feedback, ChatGPT fixed 84% of the tasks at TOP-5 and averaged 60% correct among its five completions, rivaling the Refactory (Hu et al., 2019) semantic repair tool. Notably, providing the verbose assignment description reduced accuracy, confirming that concise prompts better focus the model’s attention.
In summary, single-turn zero-shot prompting already addresses a non-trivial portion of real bugs, ranging from small algorithmic errors to student-level assignments, and continued model scaling continues to raise this ceiling. Nevertheless, the approach remains prompt-sensitive and fragile, and its reliability with larger, multi-hunk patches or when extra but imprecise context is supplied remains uncertain.
3.2. Few-shot Prompting
Few-shot prompting keeps the language model frozen and places a small set of bug-fix exemplars before the target defect so that the model can infer an edit pattern in a single decoding pass. Compared with zero-shot requests, these demonstrations often stabilize generation and improve accuracy. However, they must fit within the context window. As shown by work on basic few-shot (Xia et al., 2023), similarity-based selection (Gao et al., 2023), explanation-rich exemplars (Ahmed and Devanbu, 2023), and retrieval-driven prompting such as CEDAR (Nashid et al., 2023), their composition requires careful curation.
Xia et al. performed the first large-scale study that applies nine modern LLMs to APR without fine-tuning (Xia et al., 2023). They evaluated three prompt settings—complete function generation, code infilling with prefix and suffix, and single-line edits, over 1,094 bugs drawn from Defects4J 1.2/2.0 (Java) (Just et al., 2014), QuixBugs (Java & Python) (Lin et al., 2017), and ManyBugs (C) (Le Goues et al., 2015). For few-shot generation, they prefixed two exemplar bug-fix pairs ahead of the target defect and requested up to 200 samples per bug. Under this setting, Codex repaired sixty-three single-function Java bugs on Defects4J, significantly better than the previous best APR tool. The study further reveals a clear scaling trend (larger models fix more bugs), highlights the value of including suffix context during infilling, and demonstrates that simple entropy scores enable frozen LLMs to rank patches effectively. These findings suggest that even a minimal number of few-shot prompts, when paired with sufficient sampling, can rival specialized learning-based systems. That prompt design should balance exemplar relevance, context length, and sample budget.
Gao et al. designed more than two hundred Codex prompts for Java repair (Gao et al., 2023). Each prompt began with exactly four bug-fix pairs that were retrieved by lexical similarity and copied verbatim, in the same order, followed by the new buggy snippet and the instruction “produce a fixed version”. Under this setting, exact-match accuracy more than doubled relative to random exemplars and tripled relative to zero-shot. The authors concluded that prompts should prioritize highly similar examples, keep their order stable, and stop adding demonstrations once the prompt nears the context limit; otherwise, gains vanish when no close fixes exist.
Ahmed et al. prefixed eight bug-fix demonstrations, each followed by its original commit message that briefly explained the change (Ahmed and Devanbu, 2023). After these exemplars, they placed the buggy JavaScript function and asked the frozen code-davinci-002 model to return a corrected version. They sampled fifty completions and selected the patch that appeared most often. The commit messages provided lightweight reasoning traces that helped the model generalize, and majority voting raised repair accuracy by roughly ten percentage points. This approach increases token usage and assumes that informative commit logs are available.
CEDAR automatically retrieves exemplar pairs from a large corpus using embedding or token-frequency similarity and inserts them into the prompt without any model fine-tuning (Nashid et al., 2023). On Java and Python repair tasks, four retrieved demonstrations yielded about 76.55% exact-match accuracy, significantly higher than task-specific baselines and competitive with fully fine-tuned models, showing that systematic retrieval can rival manual curation while keeping the model frozen.
These studies confirm that well-chosen exemplars, whether similarity-ranked, retrieval-selected, or explanation-enriched, can guide frozen LLMs to competitive one-step repairs. Open challenges include finding suitable examples for projects without a history, compressing prompts when defects span multiple lines, and controlling costs when majority-voting or other multi-sample strategies increase token usage.
3.3. Retrieval-Augmented Generation Enhanced Prompting
Retrieval-augmented generation (RAG) enriches a single-turn repair prompt with external artifacts fetched on demand. By retrieving repository structure information (Chen et al., 2024), security-fix templates (Liu et al., 2024b), or domain-specific API knowledge encoded in a graph (Ouyang et al., 2025), a frozen language model can ground its patch in concrete evidence instead of relying solely on parametric memory. Compared with plain zero- or few-shot prompts, RAG improves factual grounding and supplies missing context while deployment remains as light as a one-shot generation.
RLCE targets repository‑level bugs that span multiple files. A static parser builds a lightweight project map and then retrieves only those out‑of‑file definitions and usages that interact with the faulty code; the selected fragments are appended to the prompt that asks GPT‑3.5 or GPT‑4 to “produce the fixed version” (Chen et al., 2024). This single‑shot context expansion raises GPT‑3.5’s top-1 fix rate on their built 124‑bug RepoBugs suite from 22.6% to 56.0% and lifts GPT‑4 from 41.1% to 75.8%, more than doubling accuracy without extra inference rounds or model updates.
T-RAP focuses on vulnerability repair. It first abstracts historical security fixes into edit-script templates (Liu et al., 2024b), in a way similar to traditional repair techniques such as HDRepair (Le et al., 2016). Given a new bug, the system retrieves the most relevant template and its concrete instance from a patch corpus, appends both to the buggy code, and asks a frozen code LLM (demonstrated with CodeT5 and GPT family models) to instantiate the fix in one pass. Experimental results show that T-RAP repairs 56.8% more vulnerabilities than the previous state-of-the-art VulRepair (Fu et al., 2022) under identical evaluation settings. The unique ingredient is template guidance: retrieval supplies a high-level edit pattern that narrows the search space, enabling the LLM to generate correct patches for complex security flaws without any iterative feedback.
DSrepair retrieves structured API knowledge from a purpose‑built Data‑Science Knowledge Graph and injects the retrieved triples plus the localized buggy snippet into a single repair prompt for GPT‑3.5, GPT‑4, and two open-source code LLMs, DeepSeek-Coder (Guo et al., 2024) and Codestral (AI, 2024), without any fine‑tuning (Ouyang et al., 2025). On the DS‑1000 benchmark (Lai et al., 2023), it fixes 44.4%, 14.2%, 20.6%, and 32.1% more data‑science bugs than the best prior baseline while cutting prompt tokens by up to 34%, showing that knowledge‑graph evidence can outperform unstructured text retrieval in one‑shot repair.
These studies illustrate the versatility of RAG: knowledge-graph retrieval boosts domain-specific API fixes, repository-aware retrieval unlocks cross-file repair, and template-guided retrieval accelerates security patching, all while preserving the simplicity of a one-shot prompt and avoiding procedural or agentic orchestration.
3.4. Analysis-Augmented Generation Enhanced Prompting
Analysis-augmented generation (AAG) feeds a frozen language model with artifacts from program analysis, such as compiler diagnostics, fault localization scores, data flow constraints, failing test logs, or execution traces. These artifacts typically appear in three forms: failing-test evidence that exposes runtime failures, semantics-aware slices that prune unrelated code, and concise execution traces that embed dynamic behavior. Recent systems illustrate each form in turn: D4C (Xu et al., 2024) injects failing-test evidence, Appatch (Nong et al., 2024) applies semantics-aware slices, and TracePrompt (Haque et al., 2025b) contributes concise execution traces. Each patch is grounded in concrete evidence and is generated in a single pass. Compared to plain zero-shot prompts, AAG consistently improves repair accuracy and keeps edits focused on fault-relevant code, while ensuring that extra context remains within the model window.
D4C aligns the prompt with decoder-only models’ next-token objective and lets GPT-4 complete the whole function while embedding the failing test case and its output (Xu et al., 2024). Without any fault-localization hints, D4C fixes 180 of 437 single-function bugs in Defects4J, 10% more than the previous state of the art that assumed perfect localization, with only ten samples per bug. The approach relies on dynamic evidence, so its gains diminish when tests give little diagnostic detail.
Appatch applies semantics-aware scoping and adaptive exemplar prompting to vulnerability repair (Nong et al., 2024). It slices the program to include only code relevant to the vulnerability, mines an exemplar database, and selects examples on the fly to guide patch generation. In evaluations on 306 real vulnerabilities, it raised F1 from 44.9% to 57.2% across four LLMs, outpacing baseline prompting by up to 98.9% and surpassing state-of-the-art static approaches by as much as 65.4%. Appatch also fixed seven of eleven zero-day vulnerabilities that none of the baselines could handle. Strengths include reduced context, tailored exemplars, and ensemble validation across models; limitations are the need for a known vulnerability location and extra token overhead from exemplar insertion.
TracePrompt augments the buggy code and failing test with a structured execution trace produced by PySnooper (Rachum et al., 2019), then issues one repair request to GPT-4 (Haque et al., 2025b). Across Refactory (Hu et al., 2019), RunBugRun (Prenner and Robbes, 2023), and HumanEval-Java (Jiang et al., 2023a), trace-based prompts improve GPT-4’s correct-program accuracy by up to 6% over error-only prompts. However, the benefit vanishes for GPT-3.5 and for traces that exceed 200 lines, showing that runtime context helps powerful models only when it is concise.
These studies reveal three complementary AAG strategies: semantics-aware slices narrow the search space, failing-test evidence enables direct debugging without prior localization, and compact execution traces inject runtime semantics. All approaches preserve the simplicity of one-shot prompting and avoid the need for parameter fine-tuning. However, each must balance the need for extra context against the limited tokens and potential prompt bias. Future work should explore adaptive artifact selection and summarization to keep this balance in large-scale settings.
3.5. Discussion and Outlook
Prompt-based repair has evolved from succinct zero-shot instructions to prompts enriched with exemplar fixes, retrieved project fragments, and static or dynamic analysis outputs. Larger frozen language models consistently repair more bugs without requiring parameter updates, and they become even more effective when supplied with focused auxiliary information, such as highly similar bug-fix pairs, carefully selected repository or API snippets, or compact failing-test logs and execution traces. Conversely, verbose or loosely related context can distract the model and harm accuracy, underscoring the need to balance relevance against prompt length. Most current pipelines still rely on manual wording or heuristic retrieval, but evidence from CEDAR (Nashid et al., 2023) shows that automatic exemplar selection and prompt compression are already practical. Continued research on adaptive artifact gathering, scalable summarization, and systematic evaluation will be essential for making prompt-centric repair a dependable component of continuous-integration workflows and developer tooling.
4. Procedural Approaches
Procedural approaches treat APR as a fixed pipeline where the language model runs only at scripted checkpoints and never alters its weights or the surrounding logic. Every choice about search scope, context retrieval, validation order, and iteration budget is hard-coded, which yields reproducible outcomes and predictable cost. Within this umbrella, four recurring subclasses can be identified. Test-in-the-loop pipelines alternate generation with unit tests and take each failure as an oracle (Xia and Zhang, 2024; Yin et al., 2024; Tang et al., 2024; Kong et al., 2024). Human-in-the-loop pipelines thread developer insight into the loop through hints, plans, or reviews that shape the next prompt (Yang et al., 2024c; Takerngsaksiri et al., 2024; Zhao et al., 2024b). Retrieval-augmented pipelines, or RAG-in-the-loop, insert concise evidence fetched by static rules between calls (Xia et al., 2024; Mansur et al., 2024; Yang et al., 2025a). Analysis-augmented pipelines, or AAG-in-the-loop, invoke compilers, debuggers, or security analyzers to feed diagnostics or traces that steer regeneration toward the root cause (Brancas et al., 2025; Xiao et al., 2025; Fakih et al., 2025; Wei et al., 2023). These subclasses differ mainly in the source of the feedback they inject and in the auxiliary tools that collect it. At the same time, they all rely on scripted control to keep the repair deterministic.
4.1. Test-in-the-Loop Pipelines
Test-in-the-loop pipelines keep an LLM frozen and alternate patch generation with automatic test execution. Lightweight conversational loops, exemplified by ChatRepair (Xia and Zhang, 2024), feed failing assertions back to the LLM and regenerate patches in the loop. Reasoning-enhanced variants such as ThinkRepair (Yin et al., 2024) cache a chain of thought and partial passes. Bandit-guided search, represented by REx (Tang et al., 2024), allocates queries based on test rewards, whereas contrast-aware schemes like ContrastRepair (Kong et al., 2024) pair failing and passing tests to isolate the faulty logic. Because every step and data flow is predetermined by the script rather than the model, these pipelines differ from fully agentic frameworks while still turning the test suite into an iterative oracle that steers the LLM toward a correct fix.
ChatRepair prompts GPT-3.5 or GPT-4 to propose a patch, executes the full test suite, then feeds the exact failing assertion messages plus the buggy lines back to the model for up to five regeneration rounds (Xia and Zhang, 2024). Its conversational prompts are lightweight and cost-efficient, letting ChatGPT fix 162/337 Defects4J bugs at roughly \$0.42 per bug while requiring no auxiliary analysis. Strengths include minimal engineering effort and strong empirical performance on real Java defects. Weaknesses include prompt sensitivity and limited scalability to multi-hunk or cross-file bugs, as failure messages alone may not pinpoint the root cause. All loop conditions, timeouts, and data exchanges are hard-wired; the LLM only supplies patch text and never alters the surrounding script.
ThinkRepair first asks a frozen ChatGPT to write an explicit chain-of-thought that diagnoses the bug, stores any partially correct patches that pass some tests, then iteratively re-prompts the model with those traces plus the remaining failing tests (Yin et al., 2024). This two-phase Collect-Fix loop converts the test suite into a knowledge source, reusing previous reasoning to accelerate convergence. The explicit reasoning and knowledge pool significantly enhance repair success, lifting it to 98 fixed bugs on Defects4J, surpassing prior conversational methods without requiring additional tooling. Its drawback is longer prompts and higher token consumption, which may stress context limits on very large projects.
REx (Refine-Explore-Exploit) frames iterative program repair as a Thompson-sampled multi-armed bandit (Tang et al., 2024). It views every partial patch as an arm whose reward is the proportion of tests currently passing. Starting from a pool of seed fixes, the algorithm samples from each arm’s Beta posterior, chooses the most promising candidate, and prompts GPT-4 with its remaining failing tests. This adaptive explore-exploit policy reduces LLM calls by roughly 2- to 5-fold on challenging benchmarks compared to breadth-first, greedy, or fixed-width schedules. On the Nonlinear Loop Invariant dataset, REx solves 73.7 percent of tasks within 300 GPT-4 queries, surpassing the best specialized solver’s 60.5 percent record. Performance still depends on a well-calibrated Beta prior; if the prior is off, the search can miss rare but vital repair paths.
ContrastRepair augments conversation-based APR by supplying ChatGPT with a contrastive pair of nearly identical test cases (one failing and one passing) to focus the model on the specific code change that triggers the bug (Kong et al., 2024). It repairs 143 of 337 bugs on Defects4J (Just et al., 2014), and it also achieves the best results on QuixBugs (Lin et al., 2017) and HumanEval-Java (Jiang et al., 2023a). The authors demonstrate that selecting the passing test with the smallest Damerau-Levenshtein distance (Damerau, 1964; Levenshtein et al., 1966) from the failing test yields richer feedback and reduces the number of conversation rounds required for a correct patch. However, this method relies on the presence of at least one closely matching passing test and extra retrieval time, so its benefit diminishes when no closely matching passing case exists, when generating one by mutation exceeds the time budget, or when the synthetic passing test hides the real fault.
These representatives illustrate two consistent patterns. First, richer feedback yields higher fix rates: moving from raw assertion messages to contrastive test pairs, chain-of-thought traces, or adaptive candidate selection progressively boosts success without altering model weights. Second, feedback granularity trades off against resource usage. Lightweight conversations scale to hundreds of bugs at minimal cost, whereas trace-driven, contrast-driven, or search-intensive loops achieve greater accuracy at the expense of longer prompts, more LLM calls, and added engineering complexity. A promising future direction is an adaptive tiered pipeline that escalates from inexpensive cues to detailed analysis only when simpler signals fail, balancing efficiency with effectiveness across diverse defect types.
4.2. Human-in-the-Loop Pipelines
Earlier non-LLM studies have underscored the necessity of a human in the loop for APR (Geethal et al., 2023; Sobania et al., 2023). Human-in-the-loop pipelines embed frozen LLMs in a structured workflow that alternates between the LLMs’ inference and human insight. Recent work spans interactive tutoring for novice programmers (Yang et al., 2024c), planner-and-coder collaborations on industrial issues (Takerngsaksiri et al., 2024), and reviewer-driven patch refinement (Zhao et al., 2024b). For each human-in-the-loop approach, progress occurs when the LLM revises its patch after receiving new human feedback, such as tutor hints, input labels, design rationales, or reviewer advice.
CREF partners a programming tutor with a frozen GPT-4 in an interactive repair dialogue (Yang et al., 2024c). Each cycle shows the model the buggy function, failing tests, and the tutor’s natural-language hint. The tutor then reviews the candidate’s patch, supplying a fresh hint or approval that becomes part of the next prompt. Because new human guidance is injected at every turn, the dialogue converges quickly and lifts top-5 correctness to 76.6% on the 1,239-bug TutorCode (TutorCode, 2024) benchmark, 17-25% above no-tutor baselines.
HULA is a human-in-the-loop LLM agent framework, where an AI Planner and an AI Coding Agent work with a human engineer to localize files, draft a coding plan, and generate code patches for each issue (Takerngsaksiri et al., 2024). Experimental results on SWE-bench, the planner reached 86% file-recall and the coder attained 45% LLM-based code-similarity, while in an online study with 663 real Jira issues 82% of plans were approved and 56 pull requests were merged. The authors find that detailed issue descriptions and interactive human feedback are critical, since practitioner agreement rises when engineers enrich descriptions and review agent output. Nevertheless, HULA still struggles to guarantee functional correctness beyond unit tests and requires substantial context preparation.
DRCodePilot embeds frozen GPT-4 in a dialogue-style loop that repeatedly injects developer design rationales and reviewer-style feedback (Zhao et al., 2024b). Each cycle starts with DRMiner (Zhao et al., 2024c)-extracted solution information, letting GPT-4 locate the buggy segment and draft a patch; a tuned CodeT5P (Wang et al., 2023) then supplies a reference patch, while an identifier recommender offers renaming hints. GPT-4 reflects on this mixed human feedback, producing a refined fix. On a new 938-issue Flink and Solr benchmark, it yields 109 and 18 exact-match patches, respectively, 4.7x and 3.6x more than GPT-4 zero-shot, with CodeBLEU gains of 5.4% and 3.9%.
These studies illustrate complementary avenues for leveraging human insight into procedural repair loops, ranging from issue-tracker rationales to iterative tutor coaching, developer-supplied documents and hints, reviewer comments, and expert Socratic questioning. Across these settings, human feedback injected between LLM generations transforms a frozen model into an effective repair partner, while minimizing computational cost and workflow complexity.
4.3. RAG-in-the-Loop Pipelines
RAG-in-the-Loop pipelines keep the LLM frozen and insert a deterministic retrieval stage between successive generations so that each new prompt is enriched with fresh, highly targeted evidence. The control flow is scripted by the designer rather than chosen by the model, placing these methods squarely in the procedural paradigm. Within this paradigm, hierarchical repository pruning exemplified by Agentless (Xia et al., 2024), forum-augmented iteration as used in RAGFix (Mansur et al., 2024), and knowledge-graph expansion adopted in KGCompass (Yang et al., 2025a) all show that precise retrieval can raise repair accuracy while containing token cost.
Agentless follows a three-phase loop that first asks GPT-4o to rank suspicious files, then shows only their signatures so the model can pick concrete edit lines, and finally reveals full code for patch generation, followed by test validation (Xia et al., 2024). The hierarchical retrieval supplies exactly the required context at each stage and avoids exposing the entire repository up front. On SWE-bench Lite, the initial pipeline with GPT-4o repaired 32.0% of issues at about \$0.70 per bug, outperforming many tool-driven autonomous agents with much less tokens. More recently, Agentless has been adopted by various leading companies (including OpenAI, Meta, and DeepSeek) and inspired the design of various other recent software agents. While promising, the repository traversal can be inefficient or imprecise on very large projects, and the fixed three rounds may under-explore complex defects, suggesting adaptive depth or early-exit heuristics.
RAGFix targets Python errors by mining a six-million-post Stack Overflow vector knowledge graph: after an initial failure, it extracts the stack trace, retrieves the most relevant discussions, summarizes them, and prepends the advice to a follow-up GPT-3.5 or Llama-3 prompt (Mansur et al., 2024). The system iterates once or twice, each cycle combining feature extraction, retrieval, and regeneration, which markedly raises pass@1 accuracy on the HumanEvalFix (Muennighoff et al., 2023) benchmark, especially for small 8b-parameter models that lack strong built-in API knowledge. Leveraging real developer discussions provides rich, domain-specific hints without any parameter updates. Effectiveness depends on high-quality forum matches; therefore, projects that rely on proprietary APIs may require an in-house corpus or hybrid retrieval sources.
KGCompass converts a repository and its issue tracker into a heterogeneous knowledge graph, then, before each regeneration, expands from the buggy node to fetch the twenty most relevant code blocks along dependency edges (Yang et al., 2025a). These snippets and the current failing assertions are concatenated into the next GPT-4 prompt, providing high-signal structural context. The method fixes 46.0% of SWE-bench Lite issues while achieving 51.3% function-level localization accuracy at roughly \$0.20 per bug. On the SWE-bench Lite leaderboard, KGCompass achieves the best Claude 3.5 Sonnet score at 46.0% resolved and the best DeepSeek-V3 score at 36.7% resolved. Constructing and updating the knowledge graph incurs pre-processing overhead and can miss dynamic links, such as reflection. Incremental graph maintenance and lightweight dynamic traces are promising extensions.
These studies confirm that designer-defined retrieval, hierarchical pruning, forum mining, search-tree feedback, or graph expansion delivers concise, high-signal context that steers a frozen model toward correct edits while keeping low costs. Remaining challenges include pre-processing overhead, coverage of dynamic behavior, and balancing retrieval granularity against token economy; addressing these issues will further narrow the gap between procedural RAG pipelines and fully autonomous repair agents.
4.4. AAG-in-the-Loop Pipelines
Analysis-augmented generation (AAG) pipelines interleave an automated program analysis pass with LLM repair steps. Evidence can stem from logical inconsistency sets exploited by FormHe (Brancas et al., 2025), from semantics-based token pruning in Repilot (Wei et al., 2023), or static and hybrid security analyzers such as CodeQL (De Moor et al., 2007) adopted by PredicateFix (Xiao et al., 2025) and the exploit validators embedded in LLM4CVE (Fakih et al., 2025). These signals guide regeneration without relying solely on unit tests.
Repilot leverages a frozen LLM with a semantics-based Completion Engine—concretely, the incremental parser and code-completion API of Eclipse JDT Language Server (Foundation, 2023; Foundation and Wei, 2023), which inspects the partial AST, enumerates type-, scope-, and import-consistent continuations, prunes invalid tokens, and actively completes rare identifiers during autoregressive decoding (Wei et al., 2023). The workflow iterates token by token: the LLM supplies a probability list for the next token, the engine eliminates options statically infeasible, and occasionally inserts a full identifier when only one continuation type-checks, after which generation resumes. Evaluated on single-hunk Java bugs from Defects4J 1.2/2.0, Repilot fixes 66 and 50 bugs, respectively—27 % and 47 % more than prior state-of-the-art APR tools—and sustains a 66 % compilation rate at 30 samples per bug. Because static analysis is consulted for every partial program, Repilot delivers markedly higher validity and unique fixes but incurs roughly 7 % overhead for CodeT5-large (Wang et al., 2021) and negligible overhead for InCoder-6.7B (Fried et al., 2022). The authors argue that such fine-grained pruning is orthogonal to model choice and can be generalized to other generation tasks. However, scalability to multi-file projects and languages beyond Java remains an open question.
FormHe combines an ASP solver with a fine-tuned LLM for student assignments (Brancas et al., 2025). The solver first computes Minimal Strongly Inconsistent Correction Subsets and other hints, identifying the rules that violate the reference specification. These hints form a compact prompt for the LLM, which rewrites only the suspicious rules. If the repaired program still produces extra or missing answer sets, the solver extracts new MSICSs (Kullmann, 2009) and the loop repeats. In classroom data, the tool pinpointed faulty lines in $94\%$ of submissions and produced fully correct repairs for $58\%$ of the incorrect cases.
PredicateFix targets security alerts raised by CodeQL (De Moor et al., 2007) and GoInsight. Their analyzer inspects the rule predicates behind an alert, derives a bridging predicate that links the defect to missing sanitization, and then retrieves key examples that satisfy that predicate (Xiao et al., 2025). These examples, along with the original alert, are incorporated into a retrieval-augmented prompt, guiding GPT-4 or other models to generate a patch. After each proposal, the analyzer reruns; unresolved alerts trigger another round while already-fixed code is frozen. Across 6,027 CVE-based alerts, the pipeline raised the number of correct repairs by $27.1\%$-$72.5\%$ relative to plain RAG baselines.
LLM4CVE presents an iterative vulnerability-repair loop for real-world C and C++ functions (Fakih et al., 2025). The pipeline first asks an LLM for a candidate fix, then applies a combination of static heuristics, compilation checks, and custom proof-of-concept exploits to judge whether the vulnerability is eliminated. Failure feedback, including compiler errors and exploit traces, is summarized and fed back into the next prompt. Using Llama-3-70B (Grattafiori et al., 2024), three feedback rounds improved ground-truth similarity by $20\%$ and earned a human quality score of $8.51/10$.
Continuous analyzer feedback filters hallucinations, improves localization, and reduces overfitting to weak oracles. This benefit comes at the cost of longer running times and larger prompts, especially when many alerts remain. Future work should further cut overhead by reusing partial analysis results.
4.5. Discussion and Outlook
Procedural workflows provide predictable cost and reproducible behavior because the script predetermines every retrieval step, test execution, and human intervention. Results from ChatRepair, ThinkRepair, Agentless, and KGCompass indicate that repair accuracy rises as the feedback given to a frozen model becomes more precise; structured traces, knowledge-graph context, and reviewer comments offer clearer guidance than raw assertion text. Future work can improve efficiency by prioritizing feedback according to its information value, so that prompts contain only the most helpful facts, and by allowing the loop to increase its analytical depth only when simpler cues are insufficient. Integrating firm control of scripted pipelines with small autonomous routines for complex faults may extend coverage without incurring additional costs. These advances will support the reliable adoption of LLM-based repair in continuous integration practices.
5. Agentic Approaches
Agentic frameworks delegate the choice of each repair step to the language model, rather than prescribing a fixed sequence in code. The key distinction is straightforward: when handwritten logic selects the next action, the approach is procedural, whereas a model-driven choice makes it agentic. Within this category, autonomy rises along three tiers. Tool-augmented agents place one model at the center and let LLMs call repository utilities, build or test commands, and even perform graphical actions, while the outer loop stays static (Yang et al., 2024a, b; Wang et al., 2024; Bouzenia et al., 2024; Zhang et al., 2024b; Luo et al., 2025; Liu et al., 2025). LLM-as-judges detach generation from evaluation by adding a critic LLM that scores or ranks patches, guiding an external search without executing code (He et al., 2025; Zhou et al., 2024c; Li et al., 2024b; Yadavally et al., 2025). Self-controlled systems hand one or more LLMs full authority over planning, delegation, and termination, enabling them to tackle function-level or multi-file bugs at the cost of heavier coordination (Tao et al., 2024; Antoniades et al., 2024; Örwall, 2024; Su et al., 2025). The remainder of this section examines these three paradigms in depth.
5.1. Tool-Augmented Agents
Tool-augmented agents place a frozen LLM at the center of a fixed perceive-think-act loop and expose a curated toolbox to that loop. The skeleton which “observe the state, reason, select a tool, apply it, repeat” is handwritten and never reordered at run time; the LLMs are limited to deciding which provided tool to call or whether to stop. Systems that follow this pattern include SWE-Agent (Yang et al., 2024a), SWE-Agent M (Yang et al., 2024b), OpenHands (Wang et al., 2024), RepairAgent (Bouzenia et al., 2024), AutoCodeRover (Zhang et al., 2024b), LANTERN (Luo et al., 2025), and VulDebugger (Liu et al., 2025), all of which improve over one-shot prompting by letting LLMs choose among grep, compilation, testing, translation, or debugging actions while the outer control flow stays static.
SWE-Agent (Yang et al., 2024a) uses GPT-4 Turbo in a ReAct (Yao et al., 2023) loop that interleaves git_grep, targeted file edits, and unit-test runs to drive repair without any predefined call order. In large-scale experiments on SWE-bench (Jimenez et al., 2024) it fixed 12.47% of 2294 real issues, outperforming prior scripted methods while requiring no fine-tuning. The agent’s chief advantage is adaptability, as the model can dynamically choose the tool that seems most informative at each stage. Its drawback is cost: every reasoning step consumes new tokens and runtime, so the repair process is computationally expensive. Moreover, SWE-Agent cannot invent brand-new actions at run time; it can only select from the predefined toolkit registered by the authors.
SWE-Agent M extends SWE-Agent by adding a build_and_screenshot tool that compiles the project, launches a headless browser, captures the rendered UI, and feeds the image to GPT-4V for inspection (Yang et al., 2024b). On their built SWE-bench Multimodal benchmark, it increases the fix rate by 2.3 percentage points over the original agent while maintaining the same reasoning loop, indicating that lightweight visual cues can help resolve issues rooted in front-end behavior without redesigning the core architecture.
OpenHands broadens the action set by wrapping compiler options, package-manager queries, and build-system toggles as callable web APIs and guides GPT-4 through an explicit step-by-step plan (Wang et al., 2024). On HumanEval-Java (Jiang et al., 2023a), it cut the average iterations to reach a passing solution by about thirty percent relative to a comparable agent that lacked these extra tools. The wider tool menu and structured planning yield higher coverage and faster convergence on a diverse range of bugs. However, integrating and orchestrating many external actions adds engineering overhead, and the system still depends on the model’s judgment to apply each tool correctly.
AutoCodeRover views each failing test as a sub-goal (Zhang et al., 2024b). It navigates the abstract syntax tree to focus the LLM on suspicious functions, applies spectrum-based fault localization when tests exist, and retests immediately after every patch. The agent repaired about 19% of SWE-bench Lite issues at a mean cost of \$0.43 per bug. Structured context retrieval enhances search precision and minimizes tokens. Effectiveness drops when tests are weak or the project lacks a clear modular structure.
RepairAgent wraps an LLM inside a finite-state controller that exposes tools for reading files, running tests, compiling code, and editing sources (Bouzenia et al., 2024). The model chooses actions, observes outputs, and loops until tests pass or progress stalls, reaching 164 fixes on the Defects4J v1.2/2.0 benchmark. Tool freedom lets the agent mimic a developer and achieve state-of-the-art accuracy on this suite. Long exploratory sessions consume many tokens, and careful guardrails are required to avoid infinite loops.
LANTERN tackles low-resource languages by translating buggy code into a language where the LLM is stronger, repairing it there, and translating the patch back (Luo et al., 2025). A decision module picks the intermediate language, and iterative testing guides further rounds of translation and refinement. On the xCodeEval (Khan et al., 2024) benchmark, it raised Rust fix rates by over 22% and improved other under-represented languages as well. Cross-language transfer leverages the strengths of LLMs beyond the source language. Accurate round-trip translation is challenging, and the multi-stage process introduces latency and the risk of semantic drift.
VulDebugger embodies a “debug like a human” philosophy for security repair (Liu et al., 2025). Starting from a proof-of-concept crash, the agent launches the program under a debugger, compares the actual runtime state with expected crash-free constraints, and iteratively edits code until the constraints hold. Across 50 real-world C/C++ projects, VulDebugger fixed 60% of vulnerabilities, more than doubling prior state-of-the-art security agents. Dynamic state inspection pinpoints root causes that static retrieval misses, but the heavy reliance on platform-specific debuggers complicates deployment and slows throughput.
In summary, tool-augmented agents demonstrate that granting an LLM selective tool access significantly enhances fix rates compared to procedural scripts. However, every additional reasoning step and external call still follows the same static loop. Future work should consider caching tool outputs, incorporating lightweight heuristics, or adaptively adjusting the reasoning budget to maintain flexibility while reducing the efficiency gap with purely scripted pipelines.
5.2. LLM-as-Judges
LLMs can serve not only as code generators but also as judges that evaluate candidate patches. This discriminative role, highlighted by He et al. (He et al., 2025), positions them as scalable proxies for human reviewers when assessing whether a change fixes a defect or eliminates a vulnerability. Concrete judge modules now appear in 4 emerging lines of work: differential patch testing for benchmark validation (Wang et al., 2025), automatic patch correctness assessment (Zhou et al., 2024c), security-fix isolation (Li et al., 2024b), and reference-based execution-free critique (Yadavally et al., 2025). Each line integrates an LLM critic into an APR pipeline to filter, rank, or label patches.
LLM4PatchCorrect uses in-context learning with an LLM for code to judge whether a candidate patch is correct by jointly processing bug descriptions, failing tests, execution traces, test coverage, and previously labeled patches (Zhou et al., 2024c). Using StarCoder as the backbone and evaluated on public automatic patch correctness assessment (APCA) datasets (Defects4J (Just et al., 2014) and BEARS (Madeiral et al., 2019)), it achieved 84.4% accuracy and an 86.5% F1 score without requiring any labels from the target repair tool. In this way, LLM4PatchCorrect can reduce the manual labeling effort when building a model to automatically assess the correctness of generated patches of new APR tools. Both the training and test bugs originate from a limited set of projects, so the tool may not generalize well to other software.
VulSifter identifies the lines in a security-related commit that remove a vulnerability while ignoring cosmetic edits (Li et al., 2024b). It feeds the full diff and vulnerability description into an LLM, then applies handcrafted heuristics to discard incidental changes before accepting the patch. After scanning 5.3 million commits from 127 thousand projects, the workflow isolates 11,632 verified security fixes, and models trained on this purified data generalize better than those trained on raw commits. The LLM-rule hybrid yields high precision, but because it makes decisions after the commit, VulSifter cannot issue real-time alerts during an active repair loop.
Execution-Free Critic supplies instant feedback by comparing a candidate patch with a known ground-truth fix without running code or tests (Yadavally et al., 2025). The critic LLM inspects each modified region, reasons about type safety and semantics, and predicts whether the edited program will compile and behave as the reference solution. On a repository-level benchmark, it matched actual test outcomes in 84.8% of cases and achieved a token-level judgment F1 of 0.916, surpassing static metrics and prior heuristics by up to 72%. Its first advantage is speed: removing build and test execution slashes evaluation latency in multi-step repair loops. Second, token-level commentary highlights exactly which parts of the candidate diverge from the oracle, giving search algorithms fine-grained guidance. The method, however, presupposes access to a correct patch, so it mainly fits evaluation or teacher-student training scenarios rather than real-world debugging. Performance also drops when large, cross-file patches overflow the model context, exposing the usual length limits of transformer critics.
PatchDiff (Wang et al., 2025) pioneers call-trace-guided differential patch testing: it identifies functions touched by a candidate patch, prunes surrounding context, and prompts GPT-4o-mini to generate unit tests that pass under the oracle patch but fail under the candidate, thereby acting as an automated referee. Evaluated on the 500-task SWE-bench Verified benchmark and three leading agents (CodeStory (CodeStoryAI contributors, 2025), LearnByInteract (Su et al., 2025), and OpenHands (Wang et al., 2024)), running all developer tests already invalidates 7.8% of “plausible” patches. In contrast, PatchDiff’s tests expose behavioral divergence in another 29.6%, and manual inspection confirms 28.6% of those as incorrect, inflating leaderboard scores by approximately 6.4 percentage points. The study reveals that most false positives originate from weak test suites and subtle “similar-yet-divergent” implementations, indicating that LLM-driven differential testing can rapidly identify semantic errors without requiring the full execution of the regression suite.
5.3. Self-Controlled System
Self-controlled systems remove the handwritten skeleton altogether. One or more LLMs generate or revise the global plan at run time, spawn new sub-tasks when needed, and decide when to terminate, so the model, not the script, controls the entire outer loop. Representative examples are MAGIS (Tao et al., 2024), which distributes work to four specialized roles; SWE-Search (Antoniades et al., 2024; Örwall, 2024), which couples LLMs with Monte Carlo tree search to debate and apply patches; and Learn-by-Interact (Su et al., 2025), which lets a single model explore an environment, refine its strategy, and distill the trajectory into reusable instructions. With full control, these agents can tackle multi-file or cross-module defects but incur higher coordination costs and safety risks.
MAGIS models a software team with four role-specific agents: a Manager plans high-level steps, a Repository Custodian pinpoints relevant files, Developer agents propose edits, and a QA agent validates patches (Tao et al., 2024). The agents communicate through shared memory, iteratively refine the plan, and trigger new sub-tasks until all tests pass or a time limit expires. On SWE-bench Lite, it fixed 16.67% of real GitHub issues, an eight-fold gain over a single-agent GPT-4 baseline. Division of labor provides MAGIS with a strong repository coverage and internal quality checks. The price is heavy coordination overhead and modest absolute success, leaving many hard bugs unresolved.
SWE-Search combines Monte Carlo Tree Search with a multi-agent control loop in which GPT-4o iteratively searches the repository, plans edits, and debates candidate patches before applying a fix (Antoniades et al., 2024; Örwall, 2024). On SWE-bench Lite, the system raises the solve rate to 31% with GPT-4o and delivers a 23% relative gain across five LLMs, while the Claude 3.5 Sonnet achieves 39% fixes at 2.7 issues per dollar. However, the search depth scales with inference-time computation and has been noted as “probably pretty expensive,” pointing to future work on cost-aware depth control and quicker repository traversal.
Learn-by-Interact is a self-controlled agentic framework that lets an LLM explore a new software or web environment, plan its action sequence, execute those actions, and then retrospectively distill the trajectory into reusable instructions—so the model, not a script, decides every step of the outer loop, satisfying the 5.3 criterion for full autonomy (Su et al., 2025). They fine-tune a Codestral-22B backbone on the synthesized trajectories and also demonstrate a training-free variant in which Claude-3.5 Sonnet retrieves the distilled instructions at run time to guide its decisions. On SWE-bench and other benchmarks, Learn-by-Interact improves in-context learning accuracy by up to 12.2% and supervised fine-tuning performance by 19.5% over best baselines. However, it relies on detailed and accurate environmental documentation; when documentation is sparse or rapidly changing, the data-synthesis phase can yield low-quality trajectories and incur substantial computational overhead.
Self-controlled agents, therefore, deliver the highest flexibility and already outperform simpler agents on public benchmarks. Three lessons emerge. First, explicit role separation or self-generated sub-goals narrows the search space without sacrificing autonomy. Second, stable feedback signals such as full test suites or critical debates prevent blind exploration. Third, careful cost management remains crucial because token budgets, tool latency, and coordination overhead can offset accuracy gains. Continued progress on adaptive stopping rules, incremental context caching, and lightweight self-reflection is essential before these fully autonomous agents can meet the throughput and reliability demands of continuous integration.
5.4. Discussion and Outlook
Future research in agentic repair should emphasize cost-aware planning that caches intermediate results, prunes unproductive branches, and calls judgment modules only when their expected benefit justifies the extra computational resources, thereby reducing latency. Robust context management is also essential: agents must retain only the most relevant code fragments, tests, and analysis artifacts so that large repositories remain within the model’s context window without losing critical detail. Reliability will improve when lightweight static checks and policy guards intercept infinite loops or unsafe edits before they reach version control, satisfying continuous integration requirements. Finally, a tiered workflow that applies inexpensive heuristics to straightforward defects and escalates more complex cases to a fully autonomous agent can strike a balance between efficiency and autonomy. Sustained attention to these directions can elevate current prototypes into dependable tools for everyday software engineering practice.
Table 3. Detailed Information of LLM-based APR Systems
Sub-Paradigm System Training Dataset Representative Result Assumption
Base Model Benchmark Metric Result
Fine-Tuning Full Fine-Tuning Huang (2023) Recoder dataset CodeT5 Defects4J v1.2/2.0 pass@5 82/107 Perfect Fault Localization
Jiang (2023) Recoder dataset Incoder 7B Defects4J v1.2/2.0 pass@10 31/138 Perfect Fault Localization
VulMaster CVEFixes, Big-Vul CodeT5 CVEFixes full match 20% train/test split, requires CVE type
RepairCAT APPS StarCoder-1B APR Competition 2024 pass@1000 14%
Fine-Tuning PEFT RepairLlama Megadiff CodeLlama 7B HumanEval-Java pass@10 67.28% Perfect Fault Localization
MORepair TutorLLMCode CodeLlama 13B HumanEval-Java pass@10 77.90%
Luo (2024) TutorCode CodeLlama 13B HumanEval-Java pass@10 76.88%
Li (2024) APR Instruction DeepSeek-Coder 7B HumanEval-Java pass@10 68.10% Perfect Fault Localization
Ruiz (2025) Recoder dataset DeepSeek-Coder 7B HumanEval-Java pass@10 78.53% Perfect Fault Localization
Fine-Tuning Knowledge Distilling KNOD GitHub / Defects4J v1.2/2.0 pass@1000 122/837 Perfect Fault Localization
DistiLRR MBXP CodeLlama-7B HumanEval-Perl pass@1 42.10% train/test split
NARRepair selfAPR CodeT5 Defects4J v1.2/2.0 pass@1 110/127 Perfect Fault Localization
Fine-Tuning RLFT RePair CodeNet4Repair StarCoder-16B CodeNet4Repair pass@5 65.66% train/test split
SecRepair InstructVul CodeGen2-7B InstructVul F1 82 train/test split
SWE-RL Seed RL Llama3 70B SWE-bench Verified pass@1 41.00% with Agentless pipeline
AdaPatcher ACPR CodeLlama-7B ACPR pass@1 67.57% train/test split
Smart-LLaMA-DPO built trainset Llama3.1 8B test split F1 83.4 train/test split
Fine-Tuning Context-Enriched TraceFixer CodeNet CodeT5-small CodeNet pass@10 87%
InferFix InferredBugs CodeX 12B InferredBugs pass@1 76.80% train/test split
PyTy PyTyDefects CodeT5 PyTyDefects full match 54.40% train/test split
NTR TRANSFER CodeLlama-70B HumanEval-Java pass@100 83.44% Perfect Fault Localization
Prompting Zero-Shot AlphaRepair CodeBERT Defects4J v1.2/2.0 top@k (5h) 74/473 Perfect Fault Localization
Prenner (2022) CodeX-175B Quixbugs pass@1 37/80 Perfect Fault Localization
Fan (2023) CodeX-175B LMDefects pass@5 46/113
Tian (2023) GPT-3.5 Turbo LeetCode pass@5 105/120
Prompting Few-Shot Xia (2023) CodeX 12B Defects4J v1.2 pass@600 99/255 Perfect Fault Localization
Gao (2023) CodeX-175B B2F small full match 32.25%
Ahmed (2023) CodeX-175B MODIT pass@30 31.80%
CEDAR CodeX-175B ATLAS full match 76.55%
Prompting + RAG T-RAP CodeT5 Big-Fixes pass@10 25.80% Perfect Fault Localization
RLCE GPT-4 RepoBugs pass@1 81.45%
Dsrepair DeepSeek-Coder V2 DS-1000 pass@1 29.18%
Prompting + AAG D4C GPT-4 Defects4J v1.2/2.0 pass@10 180/437 single-function bugs
Appatch Claude-3.5 Sonnet Zero-Day F1 36.46
TracePrompt GPT-4 HumanEval-Java accuracy 71.30%
Procedural Test-in-the-Loop ChatRepair GPT-3.5 Defects4J v1.2/2.0 pass@500 162/337 single-function bugs
ThinkRepair GPT-3.5 Defects4J v1.2/2.0 pass@25 98/255 single-function bugs
Rex GPT-4 LoopInv pass@300 73.70%
ContrastRepair GPT-3.5 HumanEval-Java pass@40 84.05%
Procedural Human-in-the-Loop CREF GPT-4 TutorCode avg@5 76.60% requires tutor guidance
HULA GPT-4 SWE-bench Verified pass@1 31% requires human feedback
DRCodePilot CodeT5+ Flink full match 15.27% requires design-rationale logs
Procedural RAG-in-the-Loop Agentless GPT-4o SWE-bench Lite pass@1 32%
RAGFix Llama3 70B HumanEvalFix pass@1 78%
KGCompass Claude-3.5 Sonnet SWE-bench Lite pass@1 46%
Procedural AAG-in-the-Loop Repilot Incoder 7B Defects4J v1.2/2.0 pass@5000 116/273 single-hunk bugs
FormHe Gemma 2B Student Submissions pass@1 94.20%
PredicateFix Claude-3.5 Sonnet CVE-2020-13946 pass@1 81/117
LLM4CVE Llama3 70B CVEFixes CodeBLEU +20%
Agentic Tool-Augmented SWE-Agent GPT-4 Turbo SWE-bench Lite pass@1 18.00%
SWE-Agent M GPT-4o SWE-bench Multimodal pass@1 12.20%
OpenHands Claude-3.5 Sonnet SWE-bench Lite pass@1 26%
AutoCodeRover GPT-4 SWE-bench Lite pass@1 19.00%
RepairAgent GPT-3.5 Defects4J v1.2/2.0 pass@1 164/835
LANTERN DeepSeek-V3 Rust subset of xCodeEval pass@10 87.67%
VulDebugger GPT-4 50 of ExtractFix and ARVO pass@1 60%
Agentic LLM-as-Judges VulSifter GPT-4 CleanVul F1 82%
LLM4PatchCorrect StarCoder-7B Defects4J v1.2/2.0 AUC 76.9
Execution-free Critic Claude-3 Opus SWE-bench F1 91.60%
PatchDiff GPT-4o-mini SWE-bench-Verified accuracy 28.60%
Agentic Self-Controlled System MAGIS GPT-4 SWE-bench Lite pass@1 16.67%
SWE-Search Claude 3.5 Sonnet SWE-bench Lite pass@1 39.00%
Learn-by-Interact Claude-3.5 Sonnet SWE-bench pass@1 60.00%
6. Open Challenges
Despite rapid progress, LLM-based APR still faces several open problems that cover the four paradigms.
Training and deployment cost. Fine-tuning large checkpoints needs many GPUs, long wall-clock hours, and sizable energy budgets. Even token-efficient adapters must still process millions of tokens. During inference, procedural and especially agentic workflows frequently call the LLM, invoke external tools, and rerun tests. Each round consumes thousands of tokens; running the loop locally demands high-memory GPUs, while sending it to a commercial LLM incurs substantial token charges. These long, costly sessions often exceed the few-minute latency and budget limits typical of continuous-integration pipelines, which makes real-world deployment challenging.
Evaluation reliability. Benchmarks often overlap with pre-training data, which causes data leakage and inflated scores. Weak oracles let a patch pass if any test succeeds, and underspecified test suites can miss semantic faults. This leads to patch overfitting: the model produces a plausible patch that passes every available test but still contains hidden errors. Stricter suites, extra hidden tests, and leakage-excluded datasets frequently reveal large accuracy drops. Some studies attempt to compensate by having experts manually judge patch correctness; however, manual review introduces reviewer bias and does not scale due to the cost of human verification.
Workflow robustness. The repair process is highly sensitive to multiple factors. Minor changes in prompt wording, sampling temperature, or random seeds can change the generated patch and may turn a successful repair into a failure. In agentic workflows, this randomness accumulates across steps, so two runs of the same system can follow different tool sequences and yield different results. Language models may also produce invalid commands, ignore provided rules, or loop too long, which makes the total running time unpredictable. These issues make it hard to guarantee repeatable outcomes or keep running costs within fixed limits in real-world pipelines.
Retrieval and context management. Most RAG pipelines still rely on embedding-based similarity search, which often misses some of the key files or functions and gives no recall guarantee. Even when relevant snippets are found, the added context can push the prompt beyond the LLM’s context limitation, limiting the effectiveness of RAG. Recent adaptive-memory systems, A-MEM, store each interaction as a structured note, maintain a link graph, and supply only a compact top-$k$ context, thereby improving recall while staying within the token budget (Xu et al., 2025). AAG pipelines generate diagnostics and traces directly from program analysis; however, current LLMs treat this structured data as plain text and fail to exploit its logical structure in a meaningful way.
Security and trust. LLM-generated patches often change more code than needed: they may pass the available tests yet hide latent bugs, break style guides, or slow the program. Calls to commercial APIs can leak proprietary code, and agents that bypass static checks may leave vulnerabilities unfixed. In addition, LLMs usually provide a natural-language rationale for each patch; if this explanation is incomplete or inconsistent with the actual code change, it can mislead reviewers and hinder debugging. A central open problem is how to verify both the patch and its rationale quickly and accurately, ideally before deployment, using lightweight static analysis, selective dynamic checks, and automated rollback when risks are detected.
7. Conclusion
LLM-based APR has progressed from early prototypes to a diverse ecosystem that includes fine-tuned models, prompt-centric methods, scripted pipelines, and autonomous agents. Each paradigm balances accuracy, cost, and engineering effort in a distinct manner, and no single approach prevails across all repair scenarios. Future gains are more likely to come from disciplined evaluation, smarter context selection, deeper integration with existing developer workflows, and stronger safeguards that keep patches correct, secure, and maintainable.
Comprehensive Overview.
Table 3 consolidates 63 representative LLM-based APR systems, aligning each with its paradigm, training corpus, base model, benchmark, and headline result. This overview enables readers to see how design choices impact repair performance easily and sets the stage for the forward-looking discussion that follows.
LLM-based APR: A Turning Point.
Program repair has moved beyond single-shot prompts that fix simple bugs. Modern agentic systems can scan real repositories, interpret failing tests, and submit production-ready patches. These systems already significantly surpass traditional techniques on established benchmarks. They also open the door to repairing cross-file, multilingual, and security-critical defects that prior methods rarely addressed.
LLM-based APR: A Path Forward.
Turning promising prototypes into dependable development infrastructure will require several advances. Benchmarks must be free of data leakage and patch over-fitting; training pipelines must respect practical resource limits; and retrieval engines must reliably surface the most relevant context. Agentic planners need to reason efficiently and stay within predictable time and money costs. Meeting these challenges will elevate LLM-based APR from an emerging research topic to a routine part of modern software engineering.

### v2

@@
-A Survey of LLM-based Automated Program Repair: Taxonomies, Design Paradigms, and Applications
+A Survey of LLM-based Software Repair: Taxonomies, Design Paradigms, and Applications
 Abstract.
-Large language models (LLMs) are reshaping automated program repair (APR).
-We categorize the recent 63 LLM-based APR systems published from January 2022 to June 2025 into four paradigms, and show how retrieval- or analysis-augmented contexts strengthen any of them.
-This taxonomy clarifies key trade-offs: fine-tuning delivers strong task alignment at high training cost; prompting enables rapid deployment but is limited by prompt design and context windows; procedural pipelines offer reproducible control with moderate overhead; agentic frameworks tackle multi-hunk or cross-file bugs at the price of increased latency and complexity.
-Persistent challenges include verifying semantic correctness beyond test suites, repairing repository-scale defects, and lowering the costs of LLMs.
-We outline research directions that combine lightweight human feedback, repository-aware retrieval, code analysis, and cost-aware planning to advance reliable and efficient LLM-based APR.
+Large language models (LLMs) are reshaping automated program repair.
+We present a unified taxonomy that groups 62 recent LLM-based repair systems into four paradigms defined by parameter adaptation and control authority over the repair loop, and overlays two cross-cutting layers for retrieval and analysis augmentation.
+Prior surveys have either focused on classical software repair techniques, on LLMs in software engineering more broadly, or on subsets of LLM-based software repair, such as fine-tuning strategies or vulnerability repair.
+We complement these works by treating fine-tuning, prompting, procedural pipelines, and agentic frameworks as first-class paradigms and systematically mapping representative systems to each of these paradigms.
+We also consolidate evaluation practice on common benchmarks by recording benchmark scope, pass@k, and fault-localization assumptions to support a more meaningful comparison of reported success rates.
+We clarify trade-offs among paradigms in task alignment, deployment cost, controllability, and ability to repair multi-hunk or cross-file bugs.
+We discuss challenges in current LLM-based software repair and outline research directions.
+Our artifacts, including the representation papers and scripted survey pipeline, are publicly available at https://github.com/GLEAM-Lab/ProgramRepair.
+automated program repair, large language models, retrieval augmented generation, analysis augmented generation
 *Corresponding author.
-copyright: none
+journal: TOSEMccs: Software and its engineering Software defect analysisccs: Software and its engineering Software testing and debugging
 1. Introduction
-Figure 1. Taxonomies of LLM-based APR
-LLM-based APR has become a promising research direction in software engineering (Zhang et al., 2024a; Huang et al., 2024a; Zhang et al., 2023b; Fan et al., 2023b).
-Recent code-focused LLMs, such as OpenAI’s Codex (Chen et al., 2021), CodeLlama (Roziere et al., 2023), and DeepSeek-Coder (Guo et al., 2024; Zhu et al., 2024), have demonstrated capabilities to understand and generate code, enabling them to fix bugs in programs with minimal human input.
-Early studies have shown that LLM-based APR techniques outperform traditional repair methods on repair benchmarks, achieving state-of-the-art program repair success rates (Sobania et al., 2023; Prenner et al., 2022; Xia and Zhang, 2024; Yang et al., 2024c; Jiang et al., 2023a).
-These breakthroughs have drawn intense interest from both researchers and practitioners, transforming what was once a niche topic into a vibrant field (Le Goues et al., 2025; Zubair et al., 2024).
-Indeed, new LLM-driven repair approaches are appearing at a remarkable pace, reflecting a broad optimism that this paradigm can significantly improve automated debugging and maintenance in practice (Liu et al., 2024a; Li et al., 2024a, 2025).
-Surveys conducted before the LLM era provided a taxonomy of classical APR techniques but naturally did not include approaches based on LLMs (Le Goues et al., 2019).
-More recent surveys diverge in scope: Huang et al.
-categorize repair into search, constraint, template, and learning while relegating LLMs to future work (Huang et al., 2023b); Winter et al.
-focus on developer perceptions rather than technical structure (Winter et al., 2022); Zhang et al.
-adopt a localization-generation-ranking-validation pipeline view (Zhang et al., 2023a); Haque et al.
-review tuning strategies but not full workflows (Haque et al., 2025a).
-Several niche perspectives have emerged: He et al.
-analyze the LLM-as-judge role for patch assessment rather than generation (He et al., 2025); Zhou et al.
-take a security-centric vulnerability taxonomy survey but exclude general program repair systems and newly agentic systems (Zhou et al., 2024a); Dikici et al.
-group forty-one tools under template, machine-learning, and deep-learning labels, implicitly folding GPT-style repair into a broad deep-learning bin (Dikici and Bilgin, 2025); Anand et al.
-survey LLM-based APR but cover only a small set of papers and propose no dedicated taxonomy (Anand et al., 2024).
-Overall, these surveys do not provide a unified view that encompasses prompt-driven, retrieval-augmented, procedural, and fully agentic repair workflows, nor do they address evaluation issues such as data leakage, weak oracles, or inconsistent success criteria.
-Without such a synthesis, practitioners cannot easily compare the strengths of competing strategies or identify shared research gaps.
-To address this need, we present a comprehensive survey that maps the design space across the four paradigms (fine-tuning, prompting, procedural pipelines, and agentic frameworks, as shown in Figure 1), incorporates cross-cutting Retrieval Augmented Generation (RAG) and Analysis Augmented Generation (AAG) enhancements, and highlights challenges in methodology and evaluation (surveyed papers: Jan 2022 - June 2025).
-As Figure 1 shows, parameters versus control flow define four distinct paradigms; the subsequent sections unpack their trade-offs.
-Notably, we treat certain techniques as cross-cutting enhancements rather than separate categories in this taxonomy.
-In particular, Retrieval-Augmented Generation (RAG) refers to augmenting the LLM’s input with external knowledge (such as relevant code snippets, documentation, or historical fixes) (Lewis et al., 2021; Mansur et al., 2024; Yang et al., 2025a).
-Analysis-Augmented Generation (AAG) involves incorporating program static or dynamic analysis results (such as error traces or data-flow facts) into the prompt or constraining the model’s output (Dolcetti et al., 2025; Xiao et al., 2025; Li et al., 2024a; Jain et al., 2023b).
+Figure 1. Taxonomy of LLM-based Software Repair: four paradigms defined by parameter updates and control authority, with retrieval- and analysis-augmented contexts as cross-cutting layers.
+LLM-based software repair has become a promising research direction in software engineering (Zhang et al., 2024a; Huang et al., 2024a; Zhang et al., 2023b; Fan et al., 2023b).
+Recent code-focused LLMs, including CodeT5 (Wang et al., 2021, 2023), Codex (Chen et al., 2021), CodeLlama (Roziere et al., 2023), StarCoder (Li et al., 2023; Lozhkov et al., 2024), Qwen-Coder (Hui et al., 2024), and DeepSeek-Coder (Guo et al., 2024; Zhu et al., 2024), have demonstrated capabilities to understand and generate code, enabling them to fix bugs and vulnerabilities in programs with minimal or without human input.
+Prior studies have shown that LLM-based software repair techniques outperform traditional approaches on repair benchmarks, achieving state-of-the-art performance (Sobania et al., 2023; Prenner et al., 2022; Xia and Zhang, 2024; Yang et al., 2024b; Jiang et al., 2023a).
+The research frontier has quickly moved from zero-shot prompting on single-function benchmarks (Pearce et al., 2023; Xia and Zhang, 2022) to parameter-adapted (Silva et al., 2025; Yang et al., 2025b) or tool-augmented workflows that operate at repository-level and on real vulnerabilities (Xia et al., 2025; Yang et al., 2025a), making it timely to introduce a clear taxonomy that relates these emerging paradigms and their augmentation strategies.
+This rapid shift has also motivated several surveys.
+Zhang et al.
+provide a critical study of ChatGPT’s repair behavior on a contamination-resistant benchmark and later a systematic review that organizes LLM-based software repair by utilization mode and repair scenario rather than by control structure (Zhang et al., 2023b, 2024a).
+Zhou et al.
+provide a security-oriented overview focused on vulnerability detection and repair (Zhou et al., 2024a).
+He et al.
+focus on LLM-as-judge protocols for patch assessment rather than full repair pipelines (He et al., 2025).
+Haque et al.
+survey parameter-efficient tuning for large code models and only briefly touch on repair tasks (Haque et al., 2025a).
+Anand et al.
+discuss AI-driven software repair at a high level without a taxonomy centered on control paradigms (Anand et al., 2024).
+None of these surveys organizes LLM-based repair systems jointly by control authority and parameter adaptation, or provides a cell-wise comparison of evaluation practice across benchmarks.
+Taken together, these studies clarify many aspects of modern software repair but leave three gaps.
+First, none offers a unified design space that treats fine-tuning, prompting, procedural pipelines, and agentic frameworks as first-class control paradigms tied to who steers the repair loop.
+Second, retrieval-augmented and analysis-augmented repair are not modeled as orthogonal enhancement layers that can combine with any paradigm.
+Third, prior reviews largely summarize techniques and headline scores without consolidating benchmark scope, evaluation assumptions, and pass@k definitions into a common comparison grid, which makes it difficult to interpret performance trends across paradigms and scenarios.
+To address these gaps, we present a systematic survey of 62 LLM-based software repair systems published between January 2022 and October 2025.
+Rather than grouping systems only by how they invoke an LLM, we introduce a unified design framework that organizes the space along two axes, control authority over the repair loop and whether the base model is adapted or kept frozen, yielding four paradigms (fine-tuning, prompting, procedural pipelines, and agentic frameworks) as shown in Figure 1.
+Retrieval-Augmented Generation (RAG) and Analysis-Augmented Generation (AAG) are treated as orthogonal enhancement layers that can be combined with any paradigm.
+This control- and adaptation-centered view differs from prior utilization-mode or scenario-based taxonomies and makes the role of runtime orchestration explicit.
+RAG augments the model input with external knowledge such as relevant code snippets, documentation, or historical fixes (Lewis et al., 2021; Yang et al., 2025a).
+AAG incorporates static or dynamic program-analysis results, such as error traces or data-flow facts, into the prompt or uses them to constrain the model’s output (Dolcetti et al., 2025; Xiao et al., 2025; Li et al., 2024a; Jain et al., 2023b).
 Figure 2 juxtaposes the four paradigms in parallel columns and overlays retrieval- and analysis-augmented contexts across the top bar, making explicit both the sequential flow within each paradigm and the cross-cutting enhancements that any paradigm can adopt.
+Figure 2. Overview of LLM-based software repair techniques: parallel columns show workflow control within each paradigm, and the left bar overlays retrieval and analysis augmentation.
+Our contributions.
+  1. (i)
+    We propose a unified design framework for LLM-based program repair that organizes systems along control authority and parameter adaptation into four paradigms and treats retrieval and analysis augmentation as orthogonal layers rather than separate categories, going beyond prior surveys that classify systems only by utilization mode or coarse repair scenarios.
+  2. (ii)
+    We map 62 systems into this framework and, for each, record benchmark scope, pass@k definition, and fault localization assumptions, consolidating these attributes into comparative tables that make protocol differences explicit so that reported repair rates can be interpreted across paradigms and benchmarks in a way that earlier reviews do not support.
+  3. (iii)
+    We analyze how the paradigms behave in different scenarios, such as function-level benchmarks, repository-level repair, vulnerability patching, educational tutoring, and industrial continuous integration, and derive design guidelines that relate paradigm choices to defect scope, data availability, latency, and compute budgets.
+  4. (iv)
+    We maintain a living, versioned bibliography and release an end-to-end scripted pipeline that reproduces our literature retrieval, screening, and coding process, together with all extracted metadata, so that other researchers can audit our survey and reuse the artifact at https://github.com/GLEAM-Lab/ProgramRepair.
+2. Survey Methodology
+We conduct a systematic survey of LLM-based automated program repair.
+Our goal is to build a high-quality, representative corpus of LLM-based repair systems rather than an exhaustive census of every paper that mentions LLMs and bugs.
+Following the Goal Question Metric (GQM) approach (Caldiera and Rombach, 1994), we first articulate the overall issue, object, viewpoint, and purpose, then derive research questions and analysis procedures.
+Issue.
+Fragmented definitions of repair workflows and heterogeneous evaluation protocols make it hard to compare systems, understand their strengths, and accumulate comparable results.
+Object.
+We analyze design paradigms and augmentation layers for LLM-based software repair, evaluation protocols, benchmark families and defect scopes, and reported assumptions around correctness and scalability.
+Viewpoint.
+We adopt the perspective of researchers and practitioners in software engineering, program analysis, and ML systems.
+Purpose.
+We aim to map the design space, characterize how different paradigms and benchmarks are adopted in practice, and systematize evaluation assumptions, thereby highlighting open problems around semantic correctness, repository-scale, and multi-hunk repair.
+Throughout this paper, we use LLM-based software repair as an umbrella term for repair systems whose pipelines centrally invoke a pre-trained neural code model or large language model at inference time.
+2.1. Research Questions
+  • RQ1: How has the adoption of different LLM-based software repair paradigms evolved in recent years?
+Tracking adoption shows where effort concentrates and which workflows the LLM-based software repair community favors.
+We count studies per year and per paradigm within our 62 representative works and report the resulting distributions as descriptive trends rather than a complete census of all LLM-based software repair systems.
+  • RQ2: How can control paradigm and augmentation be used to organize the LLM-based software repair design space, and how are current systems distributed across this space and deployment scenarios?
+Prior work often conflates paradigms and context sources, which blurs scope and tradeoffs.
+We build a unified four-paradigm taxonomy with retrieval and analysis augmentation as orthogonal layers and map 62 systems to cells, annotating their defect scopes, deployment scenarios, and augmentation choices.
+  • RQ3: How do evaluation protocols differ across common benchmarks, and what systematic patterns and risk factors emerge for bias or data leakage?
+For each system, we extract the benchmarks it uses, the defect scope, and the reported metrics (such as pass@k or F1), record fault-localization assumptions, and then analyze how these choices co-vary across paradigms and benchmarks.
+  • RQ4: What bottlenecks remain for semantic correctness beyond tests and for repository-scale and multi-hunk repair, and what opportunities follow?
+We synthesize the limitations and open problems reported across the surveyed systems, and relate them to the observed adoption of paradigms, benchmarks, and evaluation protocols, outlining opportunities for future LLM-based repair research.
+Figure 3. Workflow of literature survey.
+2.2. Search Strategy
+We executed structured digital-library queries on ACM DL, IEEE Xplore, DBLP, and arXiv.
+To maximize recall, we first ran a broad query set and then refined it iteratively; the last full query across all sources was run on 31 October 2025.
+The exact query templates and per-source strings are archived in our replication package.
+We then applied a year filter ($\geq$2022) to align with the period in which LLM-based software repair became technically and practically feasible.
+Popular code-targeted LLMs such as Codex were publicly available only in late 2021, so work before 2022 rarely uses LLMs for program repair and mostly falls outside our scope.
+We also automatically crawled major SE, PL, and AI venues (for example, ICSE, TOSEM, NeurIPS, and USENIX Security) using the libraries above and the internet directly, so that influential work from adjacent venues enters the raw pool even when it is not in our final representative set.
+Our query templates combine (i) core software repair terms, (ii) LLM names or model terms, (iii) benchmark names, and (iv) concrete tool or system names.
+We formed concrete queries by conjoining at least one software repair term with one or more model, benchmark, or tool keywords.
+For each data source, we adapted the syntax to its search interface, for example, restricting matches to title, abstract, and keywords where supported.
+Table 1 summarizes the keyword families used in the search.
+After collecting the raw results, we applied three rule-based automatic filters, as shown in Figure 3.
+A repair-related filter retained records whose titles or abstracts mentioned core software repair terms, along with quantitative evaluation metrics such as pass@k, accuracy, or fix rate; at this step we also recorded venue information for later use in the eligibility criteria.
+An LLM-related filter then removed records that did not mention any LLM or model keyword from Table 1.
+A benchmark-related filter finally retained records that appeared to report empirical evaluation on a reproducible dataset, either by mentioning at least one established benchmark family or by using phrases that denote a clearly described defect dataset or issue suite.
+We treat this step as a high-recall prefilter that is subsequently validated during manual screening.
+To assess the impact of this filter on recall, we manually audited a random sample of records rejected by the benchmark-related rule and refined the regular-expression rules where necessary; the audit scripts and sampled identifiers are included in our replication package.
+The exact keyword lists and regular-expression rules for all filters are implemented and documented at https://github.com/GLEAM-Lab/ProgramRepair.
+We complemented this automated filtering with backward citation snowballing and manual full-text screening, as detailed next.
+Table 1. Representative keyword families used in our search.
+Family Keywords
+Core software repair terms program repair, automated program repair, automatic program repair, software repair, code repair, bug fixing, bug repair, patch generation, patch synthesis, vulnerability repair, vulnerability patching, security patch, debugging, bug localization, fault localization
+LLM / model terms large language model, code LLM, neural language model, pre-trained model, foundation model, generative AI, GPT-2, GPT-3, GPT-3.5, GPT-4, GPT-4o, Codex, ChatGPT, LLaMA, LLaMA 2, CodeLLaMA, StarCoder, StarCoder2, CodeGen, CodeGen2, Mistral, Gemini, PaLM, Qwen, DeepSeek, Phi, T5, BERT, RoBERTa, CodeBERT, GraphCodeBERT, PLBART, CodeT5, CodeT5+, UniXcoder, InCoder
+Benchmarks Defects4J (v1.x, v2.x), SWE-bench, SWE-bench Lite, QuixBugs, ManyBugs, IntroClass, Bugs.jar, Bears, BugsInPy, CodeFlaws, BugSwarm, CVEfixes, VulnFix, VJBench, HumanEval, MBPP, DS-1000
+Tools / systems AlphaRepair, ChatRepair, ThinkRepair, REx, ContrastRepair, DRCodePilot, CREF, HULA, PATCH, KGCompass, Agentless, SWE-agent, SWE-agent M, SWE-Search, AutoCodeRover, RepairAgent, MAGIS, OpenHands, LANTERN, VulDebugger, InferFix, TraceFixer, TracePrompt, KNOD, TSAPR, SecRepair, AdaPatcher, D4C, APPATCH, SAN2PATCH, PredicateFix, LLM4CVE, RepairLLaMA, MORepair, Vul-R2
+2.3. Eligibility Criteria
+Inclusion criteria.
+A record is retained if it satisfies all of the following:
+  1. (i)
+    Repair task.
+The primary technical contribution targets automated repair or modification of software artifacts (for example, source code, scripts, configuration files, or smart contracts) in response to bugs, test failures, or security vulnerabilities, and includes at least one concrete repair task rather than only detection, localization, or code understanding.
+  2. (ii)
+    Central use of LLMs.
+The study employs a large language model at inference time as a central component of the repair pipeline, such as generating patches, editing code, ranking candidate patches, validating or criticizing patches, or orchestrating multi-step repair workflows, rather than using LLMs only for data synthesis, annotation, or commentary.
+  3. (iii)
+    Empirical evaluation on reproducible benchmarks.
+The paper reports empirical results on at least one benchmark or dataset whose evaluation setting is documented well, including but not limited to popular benchmarks such as Defects4J, SWE-bench, and other open-sourced repository-level bug or vulnerability datasets, or carefully described industrial projects.
+The paper should provide quantitative repair metrics such as the number of correctly fixed bugs, pass@$k$, or patch correctness rate.
+  4. (iv)
+    Publication quality and time frame.
+The work is published between 2022 and October 2025 in a tier-1 SE, AI, or Security venue (as listed below), or is a recent arXiv preprint with clear research relevance, for example, cited or used by follow-up work.
+This restriction is intended to keep the corpus focused on studies that have undergone a comparable review process.
+  5. (v)
+    Accessibility, extractability, and basic reporting quality.
+The full text is available in PDF format and provides sufficient methodological detail to extract key attributes of the repair setting, including LLM family, repair paradigm, benchmarks, and evaluation metrics.
+When multiple versions of the same work exist, we retain the latest version.
+During extraction we also log simple quality indicators, such as whether code or data are publicly released and whether the evaluation setup is described in enough detail to support replication.
+Exclusion criteria.
+A record is excluded if it satisfies one of the following:
+  1. (i)
+    it is a duplicate of another repair work, for instance, an earlier workshop, technical report, or arXiv version of an included conference or journal paper, in which case we retain only the most complete and latest version;
+  2. (ii)
+    it does not concern automated software repair, for example, it only addresses bug or vulnerability detection or localization, program understanding, or generic code generation, without any patch generation or patch evaluation component;
+  3. (iii)
+    it does not employ a large language model at inference time as a central component to generate, rank, or judge patches, for instance, the LLM is used only for data synthesis, labeling, or commentary, but not in the repair pipeline itself;
+  4. (iv)
+    it does not report empirical results on at least one benchmark or dataset, or it relies exclusively on proprietary data while omitting sufficient detailed information and experimental settings.
+Tier-1 venues.
+  • Software engineering venues: ICSE, FSE, ASE, ISSTA, TSE, TOSEM.
+  • AI and NLP venues: NeurIPS, ICML, ICLR, AAAI, ACL.
+  • Security venues: USENIX Security.
+2.4. Screening Flow
+Figure 3 shows the screening flow, and Table 2 reports the number of records at each stage, from 1,581 initial hits to 62 representative works.
+In the first stage, records were filtered fully automatically using the repair-related, LLM-related, and benchmark-related rules described above, without any manual intervention.
+The benchmark-related filter produced 449 records that mentioned at least one known benchmark or a clearly described defect dataset.
+This filter is intentionally permissive and serves only to prioritize papers that already signal reproducible evaluation.
+We then applied backward citation snowballing from these 449 records, scanning their references for additional LLM-based software repair systems and adding papers that also passed the automatic filters.
+This step increased the pool to 462 records before manual screening, as shown in Table 2.
+In the second stage, two authors independently conducted a manual full-text assessment of all 462 records.
+For each candidate paper, they first verified that it satisfied all inclusion criteria and excluded records that addressed only detection or localization, used LLMs solely for data synthesis or commentary, or lacked a reproducible benchmark setting, correcting any misclassifications introduced by the automatic filters.
+Disagreements were resolved through discussion, and a third author was consulted for borderline cases; we measured inter-rater agreement using Cohen’s kappa and report detailed statistics, together with the coding guidelines, in our replication package.
+This process yielded 62 primary studies that form our final survey.
+Because our eligibility criteria restrict attention to tier-1 SE, AI, and Security venues and to recent arXiv papers with visible impact, we treat these 62 studies as a representative sample of current LLM-based software repair research at these venues rather than a complete survey of all existing systems.
+For each included paper, we then extracted a common set of information, including the primary control paradigm, the use of retrieval and analysis augmentation, the defect scope and benchmark families, the base model and adaptation strategy, and the reported evaluation metrics such as the number of correctly fixed bugs or pass@$k$.
+During extraction, we also recorded basic quality indicators described above.
+Per-stage record lists, the scripts for fetching and filtering papers from multiple sources, and analysis notebooks that regenerate our tables and figures are provided in our replication package.
+Table 2. Screening statistics for the filtering of representative works.
+Stage Records remaining Notes
+Search query hits 1581 Raw keyword queries
+Repair-related filter 994 Non-repair topics removed
+LLM-related filter 726 Papers without leveraging LLM removed
+Benchmark-related filter 449 Records without benchmarks removed
+Snowballing 462 Backward citation chasing and additions
+Representative works 62 Manually screened works meeting all criteria
+3. Selected Works Overview and Adoption by Paradigm
+3.1. Taxonomy at a Glance
+To make the taxonomy concrete, Table 3 summarizes sub-paradigms across training objectives, single versus multiple generation rounds, who controls the next step of the workflow, and representative systems.
+Each system is assigned a primary paradigm and a sub-paradigm according to the classification rules in the next subsection.
+We separate systems whose main repair results rely on task-specific adaptation of a base model (Fine-Tuning) from those that keep the base model frozen (Prompting, Procedural, and Agentic, distinguished by their control structure).
+Retrieval and analysis augmentation are treated as orthogonal tags that can apply to any of the four paradigms.
 Table 3.
-Landscape of LLM-based APR by sub-paradigm.
-Columns specify the Training objective, whether generation is Single or Multiple, who Control the next step of the workflow, and the Representative Systems (scope: Jan 2022 to June 2025).
+Landscape of LLM-based software repair by sub-paradigm.
+System names in blue denote program repair, orange denote vulnerability repair, and a mixed color denotes systems evaluated on both scopes.
 Sub-paradigm Training Generation Control Representative Systems
 Fine-Tuning
-Full FT fine-tuning single scripted by human Huang et al. (2023), Jiang et al. (2023), VulMaster (2023), RepairCAT (2024)
-PEFT fine-tuning single scripted by human RepairLLaMA (2023), MORepair (2024), Luo et al. (2024), Li et al. (2024), Ruiz et al. (2025)
-Knowledge Distilling fine-tuning single scripted by human KNOD (2023), DistiLRR (2024), NARRepair (2024)
-RLFT fine-tuning single scripted by human RePair (2024), SecRepair (2024), SWE-RL (2025), AdaPatcher (2025), Smart-LLaMA-DPO (2025)
-Context-enriched FT fine-tuning single scripted by human TraceFixer (2023), InferFix (2023), PyTy (2024), NTR (2024)
+Full FT fine-tuning single scripted by human Huang et al.
+(2023) (Huang et al., 2023a), Jiang et al.
+(2023) (Jiang et al., 2023a), VulMaster (2024) (Zhou et al., 2024b), RepairCAT (2024) (Jiang and Wu, 2024)
+PEFT fine-tuning single scripted by human Li et al.
+(2024) (Li et al., 2024c), RepairLLaMA (2025) (Silva et al., 2025), MORepair (2025) (Yang et al., 2025b), Luo et al.
+(2025) (Luo et al., 2025b), Ruiz et al.
+(2025) (Ruiz et al., 2025)
+Knowledge Distillation fine-tuning single scripted by human KNOD (2023) (Jiang et al., 2023b), DistiLRR (2024) (Wong et al., 2024), NARRepair (2024) (Yang et al., 2024c)
+RLFT fine-tuning single scripted by human RePair (2024) (Zhao et al., 2024a), SecRepair (2024) (Islam et al., 2024), SWE-RL (2025) (Wei et al., 2025), AdaPatcher (2025) (Dai et al., 2025), Vul-R2 (2025) (Wen et al., 2025)
+Context-enriched FT fine-tuning single scripted by human TraceFixer (2023) (Bouzenia et al., 2023), InferFix (2023) (Jin et al., 2023), PyTy (2024) (Chow et al., 2024), NTR (2024) (Huang et al., 2024b)
 Prompting
-Zero-shot none single scripted by human AlphaRepair (2022), Prenner et al. (2022), Fan et al. (2023), Tian et al. (2023)
-Few-shot none single scripted by human Xia et al. (2023), Gao et al. (2023), Ahmed et al. (2023), CEDAR (2023)
-+ RAG none single scripted by human T-RAP (2024), RLCE (2024), DSrepair (2025)
-+ AAG none single scripted by human Appatch (2024), D4C (2024), TracePrompt (2025)
+Zero-shot none single scripted by human AlphaRepair (2022) (Xia and Zhang, 2022), Prenner et al.
+(2022) (Prenner et al., 2022), Fan et al.
+(2023) (Fan et al., 2023a), Tian et al.
+(2023) (Tian et al., 2023)
+Few-shot none single scripted by human Xia et al. (2023) (Xia et al., 2023), Gao et al. (2023) (Gao et al., 2023), Ahmed et al. (2023) (Ahmed and Devanbu, 2023), CEDAR (2023) (Nashid et al., 2023)
++ RAG none single scripted by human Ehsani (2025) et al. (Ehsani et al., 2025), RLCE (2024) (Chen et al., 2024), DSrepair (2025) (Ouyang et al., 2025)
++ AAG none single scripted by human AppAtch (2025) (Nong et al., 2025), D4C (2024) (Xu et al., 2025a), TracePrompt (2025) (Haque et al., 2025b)
 Procedural
-Test-in-the-Loop none multiple scripted by human ChatRepair (2024), ThinkRepair (2024), REx (2024), ContrastRepair (2024)
-Human-in-the-Loop none multiple scripted by human CREF (2024), HULA (2024), DRCodePilot (2024)
-+ RAG none multiple scripted by human Agentless (2024), RAGFix (2024), KGCompass (2025)
-+ AAG none multiple scripted by human Repilot (2023), FormHe (2025), PredicateFix (2025), LLM4CVE (2025)
+Test-in-the-Loop none multiple scripted by human ChatRepair (2024) (Xia and Zhang, 2024), ThinkRepair (2024) (Yin et al., 2024), REx (2024) (Tang et al., 2024), ContrastRepair (2025) (Kong et al., 2025)
+Human-in-the-Loop none multiple scripted by human CREF (2024) (Yang et al., 2024b), HULA (2025) (Takerngsaksiri et al., 2025), DRCodePilot (2024) (Zhao et al., 2024b)
++ RAG none multiple scripted by human Agentless (2025) (Xia et al., 2025), PATCH (2025) (Zhang et al., 2025), KGCompass (2025) (Yang et al., 2025a)
++ AAG none multiple scripted by human Repilot (2023) (Wei et al., 2023), SAN2PATCH (2025) (Kim et al., 2025), PredicateFix (2025) (Xiao et al., 2025), LLM4CVE (2025) (Fakih et al., 2025)
 Agentic
-Tool-Augmented Agents none multiple decided by LLM SWE-Agent (2024), AutoCodeRover (2024), RepairAgent (2024), LANTERN (2025), VulDebugger (2025), SWE-Agent M (2025), OpenHands (2025)
-LLM-as-Judges none multiple decided by LLM VulSifter (2024), LLM4PatchCorrect (2024), Execution-free Critic (2025), PatchDiff (2025)
-Self-Controlled System none multiple decided by LLM MAGIS (2024), SWE-Search (2024), Learn-by-Interact (2025)
-Figure 2. Overview of LLM-based program repair techniques
-Table 2. Annual distribution of LLM-based APR applications by paradigm
+Tool-Augmented Agents none multiple decided by LLM SWE-Agent (2024) (Yang et al., 2024a), AutoCodeRover (2024) (Zhang et al., 2024b), RepairAgent (2025) (Bouzenia et al., 2025), LANTERN (2025) (Luo et al., 2025a), VulDebugger (2025) (Liu et al., 2025), SWE-Agent M (2025) (Yang et al., [n.
+d.]), OpenHands (2025) (Wang et al., 2024)
+LLM-as-Judges none multiple decided by LLM TSAPR (2025) (Hu et al., 2025), SpecRover (2025) (Ruan et al., 2025), Abstain and Validate (2025) (Cambronero et al., 2025)
+Self-Controlled System none multiple decided by LLM MAGIS (2024) (Tao et al., 2024), SWE-Search (2025) (Antoniades et al., [n. d.]), Learn-by-Interact (2025) (Hongjin et al., 2025)
+3.2. Classification rules
+Figure 2 sketches our taxonomy along two conceptual axes: whether the base model is adapted by fine-tuning or kept frozen, and who controls the outer repair loop and its iteration pattern.
+First, we check whether the main repair results rely on training an LLM on bug-fix data.
+When the paper reports task-specific supervised fine-tuning, parameter-efficient fine-tuning, reinforcement-learning fine-tuning, or distillation, we assign the primary paradigm Fine-Tuning, even if the adapted model is later wrapped inside a more complex procedure.
+For all remaining systems that keep the base model frozen, we determine their control paradigm from the repair loop structure described in the paper, including pseudocode and pipeline figures.
+If such a system issues at most one LLM call per bug, without conditionally invoking further calls based on intermediate feedback, we treat its runtime as a single generation and assign the Prompting paradigm.
+If a system issues several LLM calls whose content depends on test outcomes, tool outputs, or previous model responses, we treat its runtime as multiple generations and distinguish between Procedural and Agentic systems based on who selects the next control step.
+When the full control flow, including when to retrieve context, which tests or analyzers to run, and when to stop, is hard-coded by the experts and LLM outputs never decide which branch to explore next or whether to accept a candidate, we assign the Procedural paradigm.
+When at least one LLM instance is responsible at run time for selecting actions or branches from a predefined menu, for example, choosing which tool to call, which candidate patch to keep, or whether to continue searching, we assign the Agentic paradigm.
+This definition covers both tool-augmented agents, in which the LLM selects external tools within a fixed skeleton, and LLM-as-Judge setups, in which a critic model accepts or rejects candidate patches and thereby controls the progression of the repair loop.
+Then, we refine each work into a sub paradigm within its primary control paradigm, as summarized in Table 3.
+For systems in the Fine-Tuning column, we assign Full FT, PEFT, Knowledge Distillation, RLFT, or Context-enriched FT based on the dominant training objective and adaptation mechanism reported in the paper.
+Full fine-tuning updates all parameters of the backbone on repair data; PEFT methods update a small adapter subset; distillation methods train a smaller student from a larger teacher or rule system; RLFT methods optimize a repair policy against execution or security rewards; context-enriched fine-tuning injects traces, static analysis artifacts, or retrieved patches into the training input.
+For Prompting systems, we distinguish Zero-shot and Few-shot according to the presence of exemplar bug fix pairs in the prompt, and we mark + RAG or + AAG when the prompt is further enriched with retrieved project artifacts or analysis outputs in a single turn.
+For Procedural systems, we classify Test-in-the-Loop when the primary feedback comes from unit tests, Human-in-the-Loop when developer input is explicitly interleaved with model calls, + RAG when deterministic retrieval stages are inserted between calls, and + AAG when the script repeatedly consults static or dynamic analysis tools.
+For Agentic systems, we label Tool-Augmented Agents when an LLM chooses among external tools inside a fixed perceive–think–act skeleton, LLM-as-Judges when a dedicated LLM critic scores or filters candidate patches as its main role, and Self-Controlled System when LLMs also plan the high-level workflow, spawn sub-tasks, and decide when to terminate.
+Retrieval and analysis augmentation are treated as orthogonal tags.
+In other words, any system can be retrieval- or analysis-augmented, regardless of whether it belongs to the fine-tuning, prompting, procedural, or agentic category.
+We record these properties through the + RAG and + AAG sub-paradigms in Table 3 where retrieval or analysis are explicit design axes, and as additional attributes for all systems.
+This separation allows us to analyze how parameter adaptation and control authority shape the core paradigm while still capturing the cross-cutting effects of richer contextual evidence.
+Figure 4. Defect scope distribution of 62 selected works.
+Figure 4 summarizes the defect scope of the 62 systems: 53 systems target general program repair only, 8 systems focus exclusively on vulnerability repair, and a single system addresses both.
+3.3. Venues of selected works
+Figure 5. Venues of selected works.
+Figure 5 summarizes the publication venues of the 62 selected studies.
+Most systems appear in top-tier software engineering conferences and journals, with additional contributions from AI and security venues, showing that LLM-based software repair is already attracting attention across multiple research communities.
+3.4. Annual Adoption by Paradigm
+Table 4 summarizes how these 62 studies are distributed across years and paradigms.
+Fine-tuning and prompting dominated early work in 2023; 2024 saw an expansion of procedural pipelines and the first wave of agent-based frameworks; 2025 shows continued diversification, especially in privacy-aware fine-tuning and self-controlled agents.
+Adoption has shifted from single-turn prompting toward agentic workflows in recent years.
+Table 4. Annual distribution of LLM-based software repair applications by paradigm
 Year Fine-Tuning Prompting Procedural Agentic Total
 2022 0 2 0 0 2
-2023 7 6 1 0 14
-2024 10 4 9 7 30
-2025 (Jan - June) 4 2 4 7 17
-Total 21 14 14 14 63
-Our survey distills several insights into the trade-offs among the four LLM-based APR paradigms.
-Fine-tuning aligns a model tightly with the repair task, producing highly accurate fixes for defects similar to those in its training data; however, it demands large bug-fix corpora and considerable GPU time.
-Prompting requires no additional training and supports rapid deployment, yet its effectiveness hinges on careful prompt engineering and is bounded by the model’s fixed knowledge and context window.
-Both paradigms can be strengthened with retrieval-augmented or analysis-augmented prompts that inject external knowledge or program-analysis feedback, thereby improving semantic correctness and reducing invalid patches.
-Procedural pipelines embed one or more LLM calls in deterministic multi-step workflows, offering reproducible control at moderate runtime cost.
-Agentic frameworks delegate planning and tool invocation to the LLM itself, enabling multi-hunk and cross-file repair at the price of increased engineering complexity, latency, and resource consumption.
-This community shift from single-turn generation to iterative, tool-assisted workflows reflects a drive to boost success rates on complex defects.
-Nonetheless, three challenges remain: verifying semantic correctness beyond limited test suites, repairing bugs that span multiple modules or exceed context limits, and mitigating the substantial computational cost of large-model pipelines.
-Addressing these issues is crucial for making LLM-based APR reliable and efficient in continuous integration settings.
-Our contributions.
-  1. (i)
-    We present the first unified taxonomy that covers four repair paradigms and two augmentation layers for LLM-based APR.
-  2. (ii)
-    We review 63 recent state-of-the-art APR systems published between Jan 2022 and June 2025, summarizing their design trade-offs, common trends, and open questions.
-  3. (iii)
-    We maintain a living bibliography at https://github.com/GLEAM-Lab/ProgramRepair.
+2023 5 6 1 0 12
+2024 9 2 5 3 19
+2025 7 4 8 10 29
+Total 21 14 14 13 62
+Answer to RQ1 In our select works, adoption shifts from single-turn prompting toward procedural and agentic workflows between 2022 and October 2025.
+We observe growing use of retrieval and analysis augmentation across all paradigms, while fine-tuning remains common when high task alignment is required.
 4. Fine-Tuning Approaches
-Fine-tuning adapts a pre-trained LLM for APR by updating its weights using bug-fix data.
+Fine-tuning adapts a pre-trained LLM for software repair by updating its weights using bug-fix data.
 Research in this area now revolves around five distinct strategies.
 Fully supervised fine-tuning updates every parameter of the backbone and achieves strong accuracy across diverse benchmarks (Jiang et al., 2023a; Zhou et al., 2024b; Jiang and Wu, 2024).
-Parameter-efficient fine-tuning freezes most of the backbone and learns small adapter modules, which lowers memory cost while keeping competitive repair performance (Silva et al., 2023; Yang et al., 2025b; Luo et al., 2024; Li et al., 2024c; Ruiz et al., 2025).
+Parameter-efficient fine-tuning freezes most of the backbone and learns small adapter modules, which lowers memory cost while keeping competitive repair performance (Silva et al., 2025; Yang et al., 2025b; Luo et al., 2025b; Li et al., 2024c; Ruiz et al., 2025).
 Reinforcement learning fine-tuning trains the model with execution or security rewards so that it optimizes for functional correctness instead of log likelihood (Zhao et al., 2024a; Islam et al., 2024; Wei et al., 2025; Dai et al., 2025; Yu et al., 2025a).
-Knowledge distillation transfers bug-fixing skill from a large teacher or rule set to a smaller student, cutting inference latency without much accuracy loss (Jiang et al., 2023b; Wong et al., 2024; Yang et al., 2024d).
-Context-enriched fine-tuning adds static-analysis results, execution traces, or retrieved patches to the input so the model can reason beyond the local snippet (Bouzenia et al., 2023; Jin et al., 2023; Chow et al., 2024; Huang et al., 2024b).
+Knowledge distillation transfers bug-fixing skill from a large teacher or rule set to a smaller student, cutting inference latency without much accuracy loss (Jiang et al., 2023b; Wong et al., 2024; Yang et al., 2024c).
+Context-enriched fine-tuning injects static-analysis results, execution traces, or retrieved patches into the training input (Bouzenia et al., 2023; Jin et al., 2023; Chow et al., 2024; Huang et al., 2024b).
 4.1. Full Fine-Tuning
-With the emergence of code-target language models based on transformers (Vaswani et al., 2017) such as GPT-2 (Radford et al., 2019), CodeT5 (Wang et al., 2021), and CodeLlama (Roziere et al., 2023), researchers have pursued full-model supervised fine-tuning on program repair scenarios.
+With the emergence of code-target language models based on transformers (Vaswani et al., 2017) such as DeepSeek-Coder (Guo et al., 2024; Zhu et al., 2024), CodeT5 (Wang et al., 2021), and CodeLlama (Roziere et al., 2023), researchers have pursued full-model supervised fine-tuning on program repair scenarios.
 Recent work explores several variants of this strategy.
@@
 Domain-focused fine-tuning demonstrates that even 1B-parameter LLMs can learn high-yield repair heuristics when specialized data are available (Jiang and Wu, 2024).
-These efforts illustrate the breadth of full-parameter supervised fine-tuning for APR.
+These efforts illustrate the breadth of full-parameter supervised fine-tuning for software repair.
 Huang et al.
 conducted an empirical study on fine-tuning LLMs (including CodeBERT (Feng et al., 2020), GraphCodeBERT (Guo et al., 2021), PLBART (Ahmad et al., 2021), CodeT5 (Wang et al., 2021), and UniXcoder (Guo et al., 2022)) for program repair tasks (Huang et al., 2023a).
-Experimental results show that LLMs in the fine-tuning paradigm can significantly outperform previous state-of-the-art APR tools on both single-hunk and multi-hunk bugs within Defects4J (Just et al., 2014).
+Experimental results show that LLMs in the fine-tuning paradigm can significantly outperform previous state-of-the-art repair tools on both single-hunk and multi-hunk bugs within Defects4J (Just et al., 2014).
 However, they reveal that LLMs usually fail to generate correct patches due to the input/output length limit of LLMs.
@@
 Jiang et al.
-conducted a large-scale study that fully fine-tunes ten pre-trained code language models across four APR benchmarks to quantify the effect of supervised adaptation on repair performance (Jiang et al., 2023a).
+conducted a large-scale study that fully fine-tunes ten pre-trained code language models across four repair benchmarks to quantify the effect of supervised adaptation on repair performance (Jiang et al., 2023a).
 Fine-tuning raised fix rates by up to an order of magnitude and showed a clear positive correlation between model size and patch success.
@@
 These designs shrink trainable parameters by up to three orders of magnitude, enable single-GPU training, and keep the frozen backbone reusable for other tasks.
-Efficiency-oriented variants demonstrate promising effectiveness in program and vulnerability repair, broadening PEFT’s applicability to larger models and consumer-grade hardware (Silva et al., 2023; Yang et al., 2025b; Luo et al., 2024).
+Efficiency-oriented variants demonstrate promising effectiveness in program and vulnerability repair, broadening PEFT’s applicability to larger models and consumer-grade hardware (Silva et al., 2025; Yang et al., 2025b; Luo et al., 2025b).
 Nevertheless, attaining strong repair accuracy still requires judicious adapter placement, appropriate rank selection, and high-quality bug-fix data, because limited adapter capacity can miss semantic links that span multiple files (Li et al., 2024c; Ruiz et al., 2025).
-RepairLLaMA augments a frozen CodeLlama-7B with rank-8 LoRA matrices (0.06% of parameters) and fine-tunes them on 4,000 curated bug-fix pairs (Silva et al., 2023).
-Training on a single GPU fixes 144 Defects4J-V2 (Just et al., 2014) bugs, 109 HumanEval-Java tasks (Jiang et al., 2023a), and 20 GitBug-Java issues (Silva et al., 2024), outperforming larger frozen models.
+RepairLLaMA augments a frozen CodeLlama-7B with rank-8 LoRA matrices (0.06% of parameters) and fine-tunes them on 4,000 curated bug-fix pairs (Silva et al., 2025).
+Training on a single GPU fixes 144 Defects4J v2 (Just et al., 2014) bugs, 109 HumanEval-Java tasks (Jiang et al., 2023a), and 20 GitBug-Java issues (Silva et al., 2024), outperforming larger frozen models.
 Limited adapter capacity and reliance on static fault localization reduce accuracy for multi-hunk or cross-file defects.
@@
 Luo et al.
-apply LoRA adapters in a federated setting where each organization trains locally on proprietary code and a server aggregates updates, then evaluates on EvalRepair-Java (Luo et al., 2024).
+apply LoRA adapters in a federated setting where each organization trains locally on proprietary code and a server aggregates updates, then evaluates on EvalRepair-Java (Luo et al., 2025b).
 The protocol preserves privacy and matches centralized training within two percentage points while exchanging only megabytes of adapter weights per round.
@@
 Li et al.
-benchmark LoRA, prefix-tuning, adapters, and (IA)3 (Liu et al., 2022) on the APR-INSTRUCTION corpus with CodeT5+ and StarCoder backbones (Li et al., 2024c).
+benchmark LoRA, prefix-tuning, adapters, and (IA)3 (Liu et al., 2022) with CodeT5+ and StarCoder backbones (Li et al., 2024c).
 The strongest adapter (IA)3 resolves 58% more bugs than earlier LLM baselines while keeping memory usage low, but its gains plateau once adapter size or data volume exceeds moderate thresholds, so careful hyper-parameter tuning remains essential.
@@
 Effective practice requires balancing adapter capacity with data diversity, enriching supervision signals (for example, federated or rationale-aware training), and pairing lightweight models with search or analysis feedback to address multi-file and semantic defects.
-2.3. Knowledge Distilling
+4.3. Knowledge Distillation
 Knowledge distillation transfers bug-fixing skills from a high-capacity teacher or knowledge source to a smaller student, allowing inference costs to drop without compromising accuracy.
-Recent studies have applied classical teacher-student distillation (including soft-target and rationale distillation) (Wong et al., 2024), dataset distillation (Yang et al., 2024d), and rule-guided distillation (Jiang et al., 2023b) to APR, yielding strong empirical gains.
+Recent studies have applied classical teacher-student distillation (including soft-target and rationale distillation) (Wong et al., 2024), dataset distillation (Yang et al., 2024c), and rule-guided distillation (Jiang et al., 2023b) to software repair, yielding strong empirical gains.
 These techniques either compress large teachers or replace them with structured knowledge, inject structural hints, restrict generation using structured rules, or enforce domain constraints while keeping the student lightweight.
@@
 KNOD distills formal syntax and semantic rules into a student decoder by aligning its token probabilities with a rule-derived teacher distribution (Jiang et al., 2023b).
-The method repaired 72 Defects4J v1.2 bugs, 25 QuixBugs, and 50 Defects4J v2.0 bugs, establishing new benchmark records.
+The method repaired 72 Defects4J v1.2 bugs and 50 Defects4J v2.0 bugs, establishing new benchmark records.
 By restricting token choices to grammar-compliant options, KNOD greatly improves compile success and patch validity.
@@
 The approach depends on high-quality synthetic rationales and yields smaller gains when the student already has ample native training data.
-NARRepair is a non-autoregressive APR model that first distills a CodeT5-large (Wang et al., 2021) teacher to build a cleaner student training set, then generates patches in parallel using action guidance and AST dependencies (Yang et al., 2024d).
-It repairs 69, 45, and 23 bugs on Defects4J (Just et al., 2014) v1.2, v2.0, and QuixBugs (Lin et al., 2017) while delivering 5.4-18.6 times faster inference than autoregressive baselines.
+NARRepair first distills a CodeT5-large (Wang et al., 2021) teacher to build a cleaner student training set, then generates patches in parallel using action guidance and AST dependencies (Yang et al., 2024c).
+It repairs 69, 41, and 23 bugs on Defects4J (Just et al., 2014) v1.2, v2.0, and QuixBugs (Lin et al., 2017) while delivering 5.4-18.6 times faster inference than autoregressive baselines.
 The combination of dataset distillation and two-stage decoding bridges the usual quality gap of non-autoregressive generation without sacrificing speed.
@@
 In this section, we also include preference-based tuning, such as DPO (Rafailov et al., 2023), which is not strictly an RL algorithm.
-Recent RL-based systems include RePair (Zhao et al., 2024a), SecRepair (Islam et al., 2024), and SWE-RL (Wei et al., 2025).
-A preference-based alternative, AdaPatcher (Dai et al., 2025) and Smart-LLaMA-DPO (Yu et al., 2025a), adopt Direct Preference Optimization (DPO) to refine their patch policy.
+Recent RL-based systems include RePair (Zhao et al., 2024a), SecRepair (Islam et al., 2024), SWE-RL (Wei et al., 2025), and Vul-R2 (Wen et al., 2025), while a preference-based alternative, AdaPatcher (Dai et al., 2025), adopts Direct Preference Optimization (DPO) to refine its patch policy.
 In program-repair settings, the reward most often combines execution signals, such as compile-and-test success, with static cues like vulnerability removal or diff similarity to reduce sparsity (Shojaee et al., 2023; Ye et al., 2025).
@@
 Compared with purely supervised fine-tuning, RLFT can push the model beyond the training distribution and align it directly with functional correctness; however, it incurs heavy computation because each gradient update may require running tests, rewards are often sparse, and poorly shaped rewards can lead to over-fitting weak test oracles (Dou et al., 2024; Liu et al., 2023; Ye et al., 2025).
-RePair starts from a fully supervised StarCoder-15B model and adds an iterative RL loop miming human debugging (Zhao et al., 2024a).
+RePair starts from a fully supervised StarCoder-15B model and adds an iterative RL loop mimicking human debugging (Zhao et al., 2024a).
 A smaller critic model, trained to distinguish good from bad patches based on compile and test outcomes, supplies rewards to the actor via PPO.
@@
 SecRepair targets vulnerability repair.
-It slices the code around each flaw, then fine-tunes a CodeGen2 (Nijkamp et al., 2023) actor with PPO where the reward combines unit-test outcomes, semantic-similarity checks, and a security bonus that favors patches eliminating unsafe patterns (Islam et al., 2024).
+It slices the code around each flaw, then fine-tunes a CodeGen2 (Nijkamp et al., 2023) actor with PPO, where the reward combines unit-test outcomes, semantic-similarity checks, and a security bonus that favors patches eliminating unsafe patterns (Islam et al., 2024).
 Trained on their built 36k-example InstructVul dataset and evaluated on zero- and N-day bugs from 6 open-source IoT operating systems, SecRepair fixes 58% of vulnerable functions and generates concise commit messages that describe the flaw, outperforming a supervised baseline by up to 22 percentage points F1 while adding only 9% runtime overhead.
@@
 The approach scales to genuine project history at low cost, although its rule-based reward may undervalue semantically correct but stylistically different patches.
-AdaPatcher adopts a strict two‑stage pipeline: it first performs supervised fine‑tuning on CodeLlama‑Instruct‑7B to train a Bug Locator that pinpoints faulty lines, then applies reinforcement learning, in the form of Direct Preference Optimization (DPO) (Rafailov et al., 2023), to a Program Modifier that refines the patch policy using preference pairs derived from unit‑test outcomes and a “fewer‑changes” reward, achieving 62% consistency and 67.57% accuracy on their newly released ACPR benchmark, a clear lead over purely supervised or prompt‑based baselines (Dai et al., 2025).
-Its strength lies in marrying location‑aware supervised fine-tuning with preference‑based RLFT: the first stage supplies reliable edit anchors, and the second stage’s lightweight reward (diff similarity + compile‑and‑test bonus) lets the model learn to generate minimal yet correct edits without the high runtime of full Monte‑Carlo execution.
-Smart-LLaMA-DPO targets smart-contract vulnerability detection and explanation through a three-stage post-training pipeline on a LLaMA-3.1-8B backbone (Yu et al., 2025a).
-First, a continual pre-training (CPT) phase feeds millions of curated Solidity contracts to solidify domain knowledge.
-Second, supervised fine-tuning (SFT) optimizes a dual-objective loss that balances detection logits and explanation likelihoods on a hand-verified dataset covering four major vulnerability types and seven machine-unauditable (MU) categories.
-Finally, the model undergoes Direct Preference Optimization (DPO) (Rafailov et al., 2023), a lightweight, reward-free RL method that updates policy logits to favor expert-preferred explanations over suboptimal ones, using paired outputs crafted by security specialists.
-On the same evaluation suite, Smart-LLaMA-DPO raises F1 by 7.5%-11.1% and accuracy by up to 10.5% over the previous SOTA DMT (Qian et al., 2023), and beats iAudit (Ma et al., 2024) on seven MU vulnerability classes by +26% F1 and +18% accuracy, while human judges mark 81%, 84%, and 95% of its explanations as correct, thorough, and clear, respectively.
-However, its current scope is limited to Solidity and single-file contracts, and scaling the preference dataset remains an open challenge.
+AdaPatcher adopts a strict two-stage pipeline: it first performs supervised fine-tuning on CodeLlama-Instruct-7B to train a Bug Locator that pinpoints faulty lines, then applies reinforcement learning, in the form of Direct Preference Optimization (DPO) (Rafailov et al., 2023), to a Program Modifier that refines the patch policy using preference pairs derived from unit-test outcomes and a “fewer-changes” reward, achieving 62% consistency and 67.57% accuracy on their newly released ACPR benchmark, a clear lead over purely supervised or prompt-based baselines (Dai et al., 2025).
+Its strength lies in marrying location-aware supervised fine-tuning with preference-based RLFT: the first stage supplies reliable edit anchors, and the second stage’s lightweight reward (diff similarity + compile-and-test bonus) lets the model learn to generate minimal yet correct edits without the high runtime of full Monte-Carlo execution.
+Vul-R2 treats vulnerability repair as a reasoning task, training a Qwen-based model that produces explanations and patches for C/C++ code through a two-stage pipeline: supervised fine-tuning for reasoning first, then reinforcement learning from verifiable rewards (Wen et al., 2025).
+In its Curriculum-based Verifiable Rewarded Training phase, an RLVR objective uses a critic LLM’s fix-or-not judgment plus ground-truth similarity and a simple format reward to score each patch.
+On the PrimeVul (Ding et al., 2025) and SVEN (He and Vechev, 2023) benchmarks, this RL phase significantly improves exact match and CodeBLEU over SFT and CoT baselines, and these gains coincide with longer, more reflective repair traces, giving unusually direct evidence that RLFT can improve reasoning rather than simply overfitting tests.
 Existing RLFT studies confirm that execution- and security-based rewards can steer language models toward semantically correct patches, yet several obstacles remain.
@@
 InferFix couples Meta’s Infer static analyzer with retrieval-augmented fine-tuning (Jin et al., 2023).
-For every alert, it retrieves analogous historical patches from a large corpus, concatenates the alert description, the buggy code, and the retrieved examples, and fine-tunes a 1.3B parameter transformer to emit a patch that clears the analyzer warning.
+For every alert, it retrieves analogous historical patches from a large corpus, concatenates the alert description, the buggy code, and the retrieved examples, and fine-tunes Codex-12B to generate patches that clear the analyzer warning.
 The additional context leads to 59 unique repairs that baseline LLMs miss and reduces false positives by filtering candidates that still trigger Infer.
@@
 An OtherTemplate token lets the model learn fixes beyond the known catalogs.
-NTR with StarCoder repairs 128 bugs on Defects4J-v2 and 129 on HumanEval-Java, surpassing the best prior work while remaining template-aware.
+NTR with CodeLlama-70B repairs 139 bugs on Defects4J v2 and 136 on HumanEval-Java, surpassing the best prior work while remaining template-aware.
 Overall, these recent studies confirm that richer semantic, dynamic, historical, and inter-procedural dependency graph contexts can materially improve LLM-based repair.
@@
 Future work will need more selective retrieval, adaptive compression, and context-aware parameter-efficient fine-tuning strategies to retain the benefits of additional signals without compromising scalability or robustness.
-2.6. Discussion and Outlook
-Research on fine-tuning for APR reveals that full supervised adaptation yields the strongest patch accuracy, as it embeds rich code knowledge.
-However, this benefit comes at a high GPU cost and the potential erosion of broader coding abilities.
-Parameter-efficient tuning recovers much of that accuracy by updating small adapter modules and low-rank matrices.
-Still, its effectiveness hinges on sensible adapter design and diverse, well-curated training data.
-Reinforcement learning introduces execution rewards that guide models toward functional correctness, yet sparse signals and repeated test runs extend training time.
-Supplying static analysis, execution traces, or retrieved historical patches broadens the LLM’s perspective.
-It helps resolve multi-file or semantic defects, though collecting and encoding such context is expensive and can exceed current input limits.
-The most promising direction is to combine these strengths: coupling lightweight adapters with selective retrieval of high-value context, applying reinforcement learning only when tests provide clear feedback, and adopting staged curricula with locate-edit-verify workflows to shorten feedback loops.
-Privacy-preserving federated or continual updates can keep models current without exposing proprietary code.
-Advances across model architecture, training signals, and deployment practices will be needed before fine-tuned repair systems can deliver reliable, low-latency fixes in everyday continuous integration environments.
+Overview Research on fine-tuning for software repair shows a clear tradeoff between accuracy, cost, and generality.
+Full supervised adaptation still delivers the strongest patch accuracy, but is expensive and can erode broader coding abilities.
+Parameter-efficient tuning and distillation recover much of this accuracy at far lower training and inference cost, yet depend on careful adapter or student design and diverse, high-quality bug-fix data.
+Reinforcement- and preference-based objectives push models toward execution-level correctness but remain sensitive to reward design and test availability, while context-enriched training that incorporates analysis results, traces, or retrieved patches improves multi-file and semantic defects at the price of heavy pre-processing and context management.
+A promising direction is to combine these strengths in cost-aware curricula and privacy-preserving continual updates, so that fine-tuned repair models can offer reliable, low-latency fixes that fit continuous integration constraints.
 5. Prompting Approaches
-Prompt-based APR treats the LLM as a frozen component that returns a complete patch after a single query, while every step of retrieval, example curation, or validation runs outside the model.
+Prompt-based software repair treats the LLM as a frozen component that returns a complete patch after a single query, while every step of retrieval, example curation, or validation runs outside the model.
 Four practical prompt styles emerge, distinguished by the amount and source of extra context they add.
@@
 Few-shot prompting precedes the defect with a handful of similar bug-fix pairs, enabling the model to imitate their edit pattern (Xia et al., 2023; Gao et al., 2023; Ahmed and Devanbu, 2023; Nashid et al., 2023).
-Retrieval-augmented prompting attaches code fragments, templates, or knowledge-graph facts fetched from external resources, grounding the patch in concrete project or domain information (Chen et al., 2024; Liu et al., 2024b; Ouyang et al., 2025).
-Analysis-augmented prompting injects diagnostics derived from static or dynamic analysis, such as failing-test logs or semantics-aware slices, to guide the model toward the root cause (Xu et al., 2024; Nong et al., 2024; Haque et al., 2025b).
+Retrieval-augmented prompting attaches code fragments, templates, or knowledge-graph facts fetched from external resources, grounding the patch in concrete project or domain information (Chen et al., 2024; Ehsani et al., 2025; Ouyang et al., 2025).
+Analysis-augmented prompting injects diagnostics derived from static or dynamic analysis, such as failing-test logs or semantics-aware slices, to guide the model toward the root cause (Xu et al., 2025a; Nong et al., 2025; Haque et al., 2025b).
 These approaches share the same single-turn interaction but trade off preparation effort, token budget, and repair accuracy in different ways.
 5.1. Zero-shot Prompting
-Zero-shot prompting keeps a pre-trained model frozen and asks it to generate a patch in a single turn by providing only the buggy code and a terse instruction such as “fix the bug”.
+Zero-shot prompting keeps a pre-trained model frozen and asks it to generate a patch in a single turn, given only the buggy code and a terse instruction such as “fix the bug”.
 Because the approach omits fine-tuning, retrieval, and multi-turn dialogue, deployment is rapid and incurs no additional training cost.
-AlphaRepair (Xia and Zhang, 2022) was the first study to demonstrate that LLMs can outperform state-of-the-art APR techniques with zero-shot.
+AlphaRepair (Xia and Zhang, 2022) was the first study to demonstrate that LLMs can outperform state-of-the-art repair techniques with zero-shot.
 Empirical studies such as Prenner et al.
@@
 (Fan et al., 2023a) on Codex and Tian et al.
-(Tian et al., 2023) on ChatGPT, show that careful wording plus a sufficiently large model can already match or sometimes surpass several traditional repair pipelines.
+(Tian et al., 2023) on ChatGPT show that careful wording plus a sufficiently large model can already match or sometimes surpass several traditional repair pipelines.
 At the same time, they expose strong prompt sensitivity, but their robustness beyond single-hunk functions remains unclear.
-AlphaRepair was the first LLM-based APR approach that performs repair in a zero-shot setting by exploiting the masked-language-modeling capability of CodeBERT without any task-specific fine-tuning (Xia and Zhang, 2022).
-AlphaRepair masks the tokens of the buggy lines and leverages CodeBERT to infill the correct code, thereby predicting the fixed codes directly instead of learning edit scripts.
-On Defects4J-v1.2/2.0, AlphaRepair repairs up to 3.3 times more bugs than the best baselines, and it also outperforms other neural and heuristic tools on QuixBugs, demonstrating that a frozen pre-trained LLM can surpass state-of-the-art APR techniques without further training.
-Fan et al.
-conducted a large-scale evaluation of zero-shot Codex on competitive-programming tasks and then studied how APR can salvage its failures (Fan et al., 2023a).
-Their prompt contained only the LeetCode problem description and the target function’s signature; no exemplar fixes were provided.
-By generating 50 solutions per task, Codex solved 46 of the 113 tasks in their new LMDefects benchmark, while subsequent APR passes (TBar, Recoder, Codex-edit) raised the total by a further 9 tasks, showing that LLM with naive zero-shot prompting already covers a substantial fraction of real bugs and that lightweight repair can extend this reach.
+AlphaRepair (Xia and Zhang, 2022) masks the tokens of the buggy lines and leverages CodeBERT to infill the correct code, thereby predicting the fixed codes directly instead of learning edit scripts.
+On Defects4J-v1.2/2.0, AlphaRepair repairs up to 3.3 times more bugs than the best baselines, and it also outperforms other neural and heuristic tools on QuixBugs, demonstrating that a frozen pre-trained LLM can surpass state-of-the-art repair techniques without further training.
 Prenner et al.
@@
 With temperature 0 and max_tokens=1024, Codex generated ten candidates per bug; a patch was accepted if QuixBugs’ reference tests all passed.
-The best prompt repaired 23 Python and 14 Java bugs, outperforming three specialized neural APR systems despite Codex never having seen any examples.
+The best prompt repaired 23 Python and 14 Java bugs, outperforming three specialized neural repair systems despite Codex never having seen any examples.
 The study also showed that adding a simple “### bug is here” hint sometimes hurts accuracy, illustrating how small prompt edits can swing outcomes.
+Fan et al.
+conducted a large-scale evaluation of zero-shot Codex on competitive-programming tasks and then studied how software repair can salvage its failures (Fan et al., 2023a).
+Their prompt contained only the LeetCode problem description and the target function’s signature; no exemplar fixes were provided.
+By generating 50 solutions per task, Codex solved 46 of the 113 tasks on their new LMDefects benchmark, showing that LLM with naive zero-shot prompting already covers a substantial fraction of real bugs and that lightweight repair can extend this reach.
 Tian et al.
@@
 Xia et al.
-performed the first large-scale study that applies nine modern LLMs to APR without fine-tuning (Xia et al., 2023).
+performed the first large-scale study that applies nine modern LLMs to repair without fine-tuning (Xia et al., 2023).
 They evaluated three prompt settings—complete function generation, code infilling with prefix and suffix, and single-line edits, over 1,094 bugs drawn from Defects4J 1.2/2.0 (Java) (Just et al., 2014), QuixBugs (Java & Python) (Lin et al., 2017), and ManyBugs (C) (Le Goues et al., 2015).
 For few-shot generation, they prefixed two exemplar bug-fix pairs ahead of the target defect and requested up to 200 samples per bug.
-Under this setting, Codex repaired sixty-three single-function Java bugs on Defects4J, significantly better than the previous best APR tool.
+Under this setting, Codex repaired sixty-three single-function Java bugs on Defects4J, significantly better than the previous best repair tool.
 The study further reveals a clear scaling trend (larger models fix more bugs), highlights the value of including suffix context during infilling, and demonstrates that simple entropy scores enable frozen LLMs to rank patches effectively.
@@
 designed more than two hundred Codex prompts for Java repair (Gao et al., 2023).
-Each prompt began with exactly four bug-fix pairs that were retrieved by lexical similarity and copied verbatim, in the same order, followed by the new buggy snippet and the instruction “produce a fixed version”.
+Each prompt began with exactly four bug-fix pairs that were retrieved by lexical similarity and copied verbatim, in the same order, followed by the new buggy snippet and the instruction for producing a fixed version.
 Under this setting, exact-match accuracy more than doubled relative to random exemplars and tripled relative to zero-shot.
@@
 After these exemplars, they placed the buggy JavaScript function and asked the frozen code-davinci-002 model to return a corrected version.
-They sampled fifty completions and selected the patch that appeared most often.
+They sampled 30 completions and selected the patch that appeared most often.
 The commit messages provided lightweight reasoning traces that helped the model generalize, and majority voting raised repair accuracy by roughly ten percentage points.
@@
 Retrieval-augmented generation (RAG) enriches a single-turn repair prompt with external artifacts fetched on demand.
-By retrieving repository structure information (Chen et al., 2024), security-fix templates (Liu et al., 2024b), or domain-specific API knowledge encoded in a graph (Ouyang et al., 2025), a frozen language model can ground its patch in concrete evidence instead of relying solely on parametric memory.
+By retrieving repository structure information (Chen et al., 2024), layered knowledge injection (Ehsani et al., 2025), or domain-specific API knowledge encoded in a graph (Ouyang et al., 2025), a frozen language model can ground its patch in concrete evidence instead of relying solely on parametric memory.
 Compared with plain zero- or few-shot prompts, RAG improves factual grounding and supplies missing context while deployment remains as light as a one-shot generation.
-RLCE targets repository‑level bugs that span multiple files.
-A static parser builds a lightweight project map and then retrieves only those out‑of‑file definitions and usages that interact with the faulty code; the selected fragments are appended to the prompt that asks GPT‑3.5 or GPT‑4 to “produce the fixed version” (Chen et al., 2024).
-This single‑shot context expansion raises GPT‑3.5’s top-1 fix rate on their built 124‑bug RepoBugs suite from 22.6% to 56.0% and lifts GPT‑4 from 41.1% to 75.8%, more than doubling accuracy without extra inference rounds or model updates.
-T-RAP focuses on vulnerability repair.
-It first abstracts historical security fixes into edit-script templates (Liu et al., 2024b), in a way similar to traditional repair techniques such as HDRepair (Le et al., 2016).
-Given a new bug, the system retrieves the most relevant template and its concrete instance from a patch corpus, appends both to the buggy code, and asks a frozen code LLM (demonstrated with CodeT5 and GPT family models) to instantiate the fix in one pass.
-Experimental results show that T-RAP repairs 56.8% more vulnerabilities than the previous state-of-the-art VulRepair (Fu et al., 2022) under identical evaluation settings.
-The unique ingredient is template guidance: retrieval supplies a high-level edit pattern that narrows the search space, enabling the LLM to generate correct patches for complex security flaws without any iterative feedback.
-DSrepair retrieves structured API knowledge from a purpose‑built Data‑Science Knowledge Graph and injects the retrieved triples plus the localized buggy snippet into a single repair prompt for GPT‑3.5, GPT‑4, and two open-source code LLMs, DeepSeek-Coder (Guo et al., 2024) and Codestral (AI, 2024), without any fine‑tuning (Ouyang et al., 2025).
-On the DS‑1000 benchmark (Lai et al., 2023), it fixes 44.4%, 14.2%, 20.6%, and 32.1% more data‑science bugs than the best prior baseline while cutting prompt tokens by up to 34%, showing that knowledge‑graph evidence can outperform unstructured text retrieval in one‑shot repair.
+RLCE targets repository-level bugs that span multiple files.
+A static parser builds a lightweight project map and then retrieves only those out-of-file definitions and usages that interact with the faulty code; the selected fragments are appended to the prompt that asks GPT-3.5 or GPT-4 to produce the fixed version (Chen et al., 2024).
+This single-shot context expansion raises GPT-3.5’s top-1 fix rate on their built 124-bug RepoBugs suite from 22.6% to 56.0% and lifts GPT-4 from 41.1% to 81.5%, more than doubling accuracy without extra inference rounds or model updates.
+Ehsani et al.
+propose a layered knowledge injection framework for LLM-based software repair that incrementally augments prompts with Bug Knowledge, Repository Knowledge, and Project Knowledge automatically extracted from the buggy function, co-changing files, Git history, documentation, and previously resolved issues (Ehsani et al., 2025).
+On the BugsInPy (Widyasari et al., 2020) benchmark of 314 Python bugs, this framework with Llama 3.3 raises the fix rate to 79% by correctly repairing 250 bugs, a 23% improvement over prior work, with 65% fixed using only the Bug Knowledge Layer, a further 9% gain from adding the Repository Knowledge Layer, and an additional 5% gain from the Project Knowledge Layer.
+They find that retrieved context should be injected hierarchically and in a bug-type-aware way, rather than being concatenated all at once, because different bug types benefit from different kinds and depths of retrieved repository and project knowledge for effective repair.
+DSrepair retrieves structured API knowledge from a purpose-built Data-Science Knowledge Graph and injects the retrieved triples plus the localized buggy snippet into a single repair prompt for GPT-3.5, GPT-4, and two open-source code LLMs, DeepSeek-Coder (Guo et al., 2024) and Codestral (AI, 2024), without any fine-tuning (Ouyang et al., 2025).
+On the DS-1000 benchmark (Lai et al., 2023), it fixes 44.4%, 14.2%, 20.6%, and 32.1% more data-science bugs than the best prior baseline while cutting prompt tokens by up to 34%, showing that knowledge-graph evidence can outperform unstructured text retrieval in one-shot repair.
 These studies illustrate the versatility of RAG: knowledge-graph retrieval boosts domain-specific API fixes, repository-aware retrieval unlocks cross-file repair, and template-guided retrieval accelerates security patching, all while preserving the simplicity of a one-shot prompt and avoiding procedural or agentic orchestration.
@@
 These artifacts typically appear in three forms: failing-test evidence that exposes runtime failures, semantics-aware slices that prune unrelated code, and concise execution traces that embed dynamic behavior.
-Recent systems illustrate each form in turn: D4C (Xu et al., 2024) injects failing-test evidence, Appatch (Nong et al., 2024) applies semantics-aware slices, and TracePrompt (Haque et al., 2025b) contributes concise execution traces.
+Recent systems illustrate each form in turn: D4C (Xu et al., 2025a) injects failing-test evidence, Appatch (Nong et al., 2025) applies semantics-aware slices, and TracePrompt (Haque et al., 2025b) contributes concise execution traces.
 Each patch is grounded in concrete evidence and is generated in a single pass.
 Compared to plain zero-shot prompts, AAG consistently improves repair accuracy and keeps edits focused on fault-relevant code, while ensuring that extra context remains within the model window.
-D4C aligns the prompt with decoder-only models’ next-token objective and lets GPT-4 complete the whole function while embedding the failing test case and its output (Xu et al., 2024).
+D4C aligns the prompt with decoder-only models’ next-token objective and lets GPT-4 complete the whole function while embedding the failing test case and its output (Xu et al., 2025a).
 Without any fault-localization hints, D4C fixes 180 of 437 single-function bugs in Defects4J, 10% more than the previous state of the art that assumed perfect localization, with only ten samples per bug.
 The approach relies on dynamic evidence, so its gains diminish when tests give little diagnostic detail.
-Appatch applies semantics-aware scoping and adaptive exemplar prompting to vulnerability repair (Nong et al., 2024).
+Appatch applies semantics-aware scoping and adaptive exemplar prompting to vulnerability repair (Nong et al., 2025).
 It slices the program to include only code relevant to the vulnerability, mines an exemplar database, and selects examples on the fly to guide patch generation.
@@
 Future work should explore adaptive artifact selection and summarization to keep this balance in large-scale settings.
-3.5. Discussion and Outlook
-Prompt-based repair has evolved from succinct zero-shot instructions to prompts enriched with exemplar fixes, retrieved project fragments, and static or dynamic analysis outputs.
+Overview Prompt-based repair has evolved from succinct zero-shot instructions to prompts enriched with exemplar fixes, retrieved project fragments, and static or dynamic analysis outputs.
 Larger frozen language models consistently repair more bugs without requiring parameter updates, and they become even more effective when supplied with focused auxiliary information, such as highly similar bug-fix pairs, carefully selected repository or API snippets, or compact failing-test logs and execution traces.
@@
 6. Procedural Approaches
-Procedural approaches treat APR as a fixed pipeline where the language model runs only at scripted checkpoints and never alters its weights or the surrounding logic.
-Every choice about search scope, context retrieval, validation order, and iteration budget is hard-coded, which yields reproducible outcomes and predictable cost.
+Procedural approaches treat software repair as a fixed pipeline where the language model runs only at scripted checkpoints and never alters its weights or the surrounding logic.
+Every choice about search scope, context retrieval, validation order, and iteration budget is hard-coded, which yields reproducible outcomes and predictable costs.
 Within this umbrella, four recurring subclasses can be identified.
-Test-in-the-loop pipelines alternate generation with unit tests and take each failure as an oracle (Xia and Zhang, 2024; Yin et al., 2024; Tang et al., 2024; Kong et al., 2024).
-Human-in-the-loop pipelines thread developer insight into the loop through hints, plans, or reviews that shape the next prompt (Yang et al., 2024c; Takerngsaksiri et al., 2024; Zhao et al., 2024b).
-Retrieval-augmented pipelines, or RAG-in-the-loop, insert concise evidence fetched by static rules between calls (Xia et al., 2024; Mansur et al., 2024; Yang et al., 2025a).
-Analysis-augmented pipelines, or AAG-in-the-loop, invoke compilers, debuggers, or security analyzers to feed diagnostics or traces that steer regeneration toward the root cause (Brancas et al., 2025; Xiao et al., 2025; Fakih et al., 2025; Wei et al., 2023).
+Test-in-the-loop pipelines alternate generation with unit tests and take each failure as an oracle (Xia and Zhang, 2024; Yin et al., 2024; Tang et al., 2024; Kong et al., 2025).
+Human-in-the-loop pipelines thread developer insight into the loop through hints, plans, or reviews that shape the next prompt (Yang et al., 2024b; Takerngsaksiri et al., 2025; Zhao et al., 2024b).
+Retrieval-augmented pipelines, or RAG-in-the-loop, insert concise evidence fetched by static rules between calls (Xia et al., 2025; Fei et al., 2025; Yang et al., 2025a).
+Analysis-augmented pipelines, or AAG-in-the-loop, invoke compilers, debuggers, or security analyzers to feed diagnostics or traces that steer regeneration toward the root cause (Kim et al., 2025; Xiao et al., 2025; Fakih et al., 2025; Wei et al., 2023).
 These subclasses differ mainly in the source of the feedback they inject and in the auxiliary tools that collect it.
@@
 Reasoning-enhanced variants such as ThinkRepair (Yin et al., 2024) cache a chain of thought and partial passes.
-Bandit-guided search, represented by REx (Tang et al., 2024), allocates queries based on test rewards, whereas contrast-aware schemes like ContrastRepair (Kong et al., 2024) pair failing and passing tests to isolate the faulty logic.
+Bandit-guided search, represented by REx (Tang et al., 2024), allocates queries based on test rewards, whereas contrast-aware schemes, such as ContrastRepair (Kong et al., 2025), pair failing and passing tests to isolate the faulty logic.
 Because every step and data flow is predetermined by the script rather than the model, these pipelines differ from fully agentic frameworks while still turning the test suite into an iterative oracle that steers the LLM toward a correct fix.
-ChatRepair prompts GPT-3.5 or GPT-4 to propose a patch, executes the full test suite, then feeds the exact failing assertion messages plus the buggy lines back to the model for up to five regeneration rounds (Xia and Zhang, 2024).
+ChatRepair prompts GPT-3.5 or GPT-4 to propose a patch, executes the full test suite, then feeds the exact failing assertion messages plus the buggy lines back to the model for an average of 21.86 regeneration rounds (Xia and Zhang, 2024).
 Its conversational prompts are lightweight and cost-efficient, letting ChatGPT fix 162/337 Defects4J bugs at roughly \$0.42 per bug while requiring no auxiliary analysis.
@@
 Performance still depends on a well-calibrated Beta prior; if the prior is off, the search can miss rare but vital repair paths.
-ContrastRepair augments conversation-based APR by supplying ChatGPT with a contrastive pair of nearly identical test cases (one failing and one passing) to focus the model on the specific code change that triggers the bug (Kong et al., 2024).
+ContrastRepair augments conversation-based software repair by supplying ChatGPT with a contrastive pair of nearly identical test cases (one failing and one passing) to focus the model on the specific code change that triggers the bug (Kong et al., 2025).
 It repairs 143 of 337 bugs on Defects4J (Just et al., 2014), and it also achieves the best results on QuixBugs (Lin et al., 2017) and HumanEval-Java (Jiang et al., 2023a).
@@
 Second, feedback granularity trades off against resource usage.
-Lightweight conversations scale to hundreds of bugs at minimal cost, whereas trace-driven, contrast-driven, or search-intensive loops achieve greater accuracy at the expense of longer prompts, more LLM calls, and added engineering complexity.
+Lightweight conversations scale to hundreds of bugs at minimal cost.
+In contrast, trace-driven, contrast-driven, or search-intensive loops achieve greater accuracy at the expense of longer prompts, more LLM calls, and added engineering complexity.
 A promising future direction is an adaptive tiered pipeline that escalates from inexpensive cues to detailed analysis only when simpler signals fail, balancing efficiency with effectiveness across diverse defect types.
 6.2. Human-in-the-Loop Pipelines
-Earlier non-LLM studies have underscored the necessity of a human in the loop for APR (Geethal et al., 2023; Sobania et al., 2023).
-Human-in-the-loop pipelines embed frozen LLMs in a structured workflow that alternates between the LLMs’ inference and human insight.
-Recent work spans interactive tutoring for novice programmers (Yang et al., 2024c), planner-and-coder collaborations on industrial issues (Takerngsaksiri et al., 2024), and reviewer-driven patch refinement (Zhao et al., 2024b).
+Earlier non-LLM studies have underscored the necessity of a human in the loop for software repair (Geethal et al., 2023; Sobania et al., 2023).
+Human-in-the-loop pipelines embed frozen LLMs into a structured workflow that alternates between LLM inference and human insight.
+Recent work spans interactive tutoring for novice programmers (Yang et al., 2024b), planner-and-coder collaborations on industrial issues (Takerngsaksiri et al., 2025), and reviewer-driven patch refinement (Zhao et al., 2024b).
 For each human-in-the-loop approach, progress occurs when the LLM revises its patch after receiving new human feedback, such as tutor hints, input labels, design rationales, or reviewer advice.
-CREF partners a programming tutor with a frozen GPT-4 in an interactive repair dialogue (Yang et al., 2024c).
+CREF partners a programming tutor with a frozen GPT-4 in an interactive repair dialogue (Yang et al., 2024b).
 Each cycle shows the model the buggy function, failing tests, and the tutor’s natural-language hint.
@@
 Because new human guidance is injected at every turn, the dialogue converges quickly and lifts top-5 correctness to 76.6% on the 1,239-bug TutorCode (TutorCode, 2024) benchmark, 17-25% above no-tutor baselines.
-HULA is a human-in-the-loop LLM agent framework, where an AI Planner and an AI Coding Agent work with a human engineer to localize files, draft a coding plan, and generate code patches for each issue (Takerngsaksiri et al., 2024).
+HULA is a human-in-the-loop LLM agent framework, where an AI Planner and an AI Coding Agent work with a human engineer to localize files, draft a coding plan, and generate code patches for each issue (Takerngsaksiri et al., 2025).
 Experimental results on SWE-bench, the planner reached 86% file-recall and the coder attained 45% LLM-based code-similarity, while in an online study with 663 real Jira issues 82% of plans were approved and 56 pull requests were merged.
@@
 Each cycle starts with DRMiner (Zhao et al., 2024c)-extracted solution information, letting GPT-4 locate the buggy segment and draft a patch; a tuned CodeT5P (Wang et al., 2023) then supplies a reference patch, while an identifier recommender offers renaming hints.
-GPT-4 reflects on this mixed human feedback, producing a refined fix.
+GPT-4 reflects on this mixed human feedback and produces a refined fix.
 On a new 938-issue Flink and Solr benchmark, it yields 109 and 18 exact-match patches, respectively, 4.7x and 3.6x more than GPT-4 zero-shot, with CodeBLEU gains of 5.4% and 3.9%.
@@
 6.3. RAG-in-the-Loop Pipelines
-RAG-in-the-Loop pipelines keep the LLM frozen and insert a deterministic retrieval stage between successive generations so that each new prompt is enriched with fresh, highly targeted evidence.
+RAG-in-the-Loop pipelines keep the LLM frozen and insert a deterministic retrieval stage between successive generations, enriching each new prompt with fresh, highly targeted evidence.
 The control flow is scripted by the designer rather than chosen by the model, placing these methods squarely in the procedural paradigm.
-Within this paradigm, hierarchical repository pruning exemplified by Agentless (Xia et al., 2024), forum-augmented iteration as used in RAGFix (Mansur et al., 2024), and knowledge-graph expansion adopted in KGCompass (Yang et al., 2025a) all show that precise retrieval can raise repair accuracy while containing token cost.
-Agentless follows a three-phase loop that first asks GPT-4o to rank suspicious files, then shows only their signatures so the model can pick concrete edit lines, and finally reveals full code for patch generation, followed by test validation (Xia et al., 2024).
-The hierarchical retrieval supplies exactly the required context at each stage and avoids exposing the entire repository up front.
-On SWE-bench Lite, the initial pipeline with GPT-4o repaired 32.0% of issues at about \$0.70 per bug, outperforming many tool-driven autonomous agents with much less tokens.
-More recently, Agentless has been adopted by various leading companies (including OpenAI, Meta, and DeepSeek) and inspired the design of various other recent software agents.
+Within this paradigm, hierarchical repository pruning exemplified by Agentless (Xia et al., 2025), multi-stage retrieving as used in PATCH (Zhang et al., 2025), and knowledge-graph expansion adopted in KGCompass (Yang et al., 2025a) all show that precise retrieval can raise repair accuracy while containing token cost.
+Agentless follows a three-phase loop that first asks GPT-4o to rank suspicious files, then shows only their signatures so the model can pick concrete edit lines, and finally reveals full code for patch generation, followed by test validation (Xia et al., 2025).
+Hierarchical retrieval provides exactly the required context at each stage and avoids exposing the entire repository up front.
+On SWE-bench Lite, the initial pipeline with GPT-4o repaired 32.0% of issues at about \$0.70 per bug, outperforming many tool-driven autonomous agents with much fewer tokens.
 While promising, the repository traversal can be inefficient or imprecise on very large projects, and the fixed three rounds may under-explore complex defects, suggesting adaptive depth or early-exit heuristics.
-RAGFix targets Python errors by mining a six-million-post Stack Overflow vector knowledge graph: after an initial failure, it extracts the stack trace, retrieves the most relevant discussions, summarizes them, and prepends the advice to a follow-up GPT-3.5 or Llama-3 prompt (Mansur et al., 2024).
-The system iterates once or twice, each cycle combining feature extraction, retrieval, and regeneration, which markedly raises pass@1 accuracy on the HumanEvalFix (Muennighoff et al., 2023) benchmark, especially for small 8b-parameter models that lack strong built-in API knowledge.
-Leveraging real developer discussions provides rich, domain-specific hints without any parameter updates.
-Effectiveness depends on high-quality forum matches; therefore, projects that rely on proprietary APIs may require an in-house corpus or hybrid retrieval sources.
+PATCH is a stage-wise framework that augments the buggy code snippet with class and repository-level dependence context, commit message programmer intent, and retrieved bug-fixing demonstrations, then simulates tester, developer, and reviewer roles with three ChatGPT agents across bug reporting, bug diagnosis, patch generation, and patch verification (Zhang et al., 2025).
+On the BFP (Tufano et al., 2019) benchmark, PATCH reaches pass@1 33.97%, pass@3 37.08%, and pass@5 39.81% with a Levenshtein distance of 21.44, compared with GPT-4 at 19.96%, 24.42%, 26.00% and 26.07 respectively.
+For RAG in the loop program repair, the main lesson is that retrieval of dependence context, commit messages, and similar bug fixing pairs is most effective when combined with an explicit multi-role, multi-stage interaction loop that lets the LLM iteratively explain, reason, and review patches instead of treating repair as a single pass generation task.
 KGCompass converts a repository and its issue tracker into a heterogeneous knowledge graph, then, before each regeneration, expands from the buggy node to fetch the twenty most relevant code blocks along dependency edges (Yang et al., 2025a).
@@
 Analysis-augmented generation (AAG) pipelines interleave an automated program analysis pass with LLM repair steps.
-Evidence can stem from logical inconsistency sets exploited by FormHe (Brancas et al., 2025), from semantics-based token pruning in Repilot (Wei et al., 2023), or static and hybrid security analyzers such as CodeQL (De Moor et al., 2007) adopted by PredicateFix (Xiao et al., 2025) and the exploit validators embedded in LLM4CVE (Fakih et al., 2025).
-These signals guide regeneration without relying solely on unit tests.
+Evidence can stem from sanitizer logs analyzed by SAN2PATCH (Kim et al., 2025), from semantics-based token pruning in Repilot (Wei et al., 2023), or from static and hybrid security analyzers such as CodeQL (De Moor et al., 2007), adopted by PredicateFix (Xiao et al., 2025), and from exploit validators embedded in LLM4CVE (Fakih et al., 2025).
+A fixed script orchestrates the analysis stages and regeneration steps, keeping these pipelines within the procedural paradigm.
+The analysis signals guide regeneration without relying solely on unit tests.
 Repilot leverages a frozen LLM with a semantics-based Completion Engine—concretely, the incremental parser and code-completion API of Eclipse JDT Language Server (Foundation, 2023; Foundation and Wei, 2023), which inspects the partial AST, enumerates type-, scope-, and import-consistent continuations, prunes invalid tokens, and actively completes rare identifiers during autoregressive decoding (Wei et al., 2023).
 The workflow iterates token by token: the LLM supplies a probability list for the next token, the engine eliminates options statically infeasible, and occasionally inserts a full identifier when only one continuation type-checks, after which generation resumes.
-Evaluated on single-hunk Java bugs from Defects4J 1.2/2.0, Repilot fixes 66 and 50 bugs, respectively—27 % and 47 % more than prior state-of-the-art APR tools—and sustains a 66 % compilation rate at 30 samples per bug.
+Evaluated on single-hunk Java bugs from Defects4J 1.2/2.0, Repilot fixes 66 and 50 bugs, respectively, 27% and 47% more than prior state-of-the-art repair tools, and sustains a 59% compilation rate at 5000 samples per bug.
 Because static analysis is consulted for every partial program, Repilot delivers markedly higher validity and unique fixes but incurs roughly 7% overhead for CodeT5-large (Wang et al., 2021) and negligible overhead for InCoder-6.7B (Fried et al., 2022).
@@
 However, scalability to multi-file projects and languages beyond Java remains an open question.
-FormHe combines an ASP solver with a fine-tuned LLM for student assignments (Brancas et al., 2025).
-The solver first computes Minimal Strongly Inconsistent Correction Subsets and other hints, identifying the rules that violate the reference specification.
-These hints form a compact prompt for the LLM, which rewrites only the suspicious rules.
-If the repaired program still produces extra or missing answer sets, the solver extracts new MSICSs (Kullmann, 2009) and the loop repeats.
-In classroom data, the tool pinpointed faulty lines in $94\%$ of submissions and produced fully correct repairs for $58\%$ of the incorrect cases.
+SAN2PATCH is an automated vulnerability repair system that uses only sanitizer logs and source code, running a four-stage Tree of Thought LLM pipeline (Vulnerability Comprehension, Where-To-Fix, How-To-Fix, Candidate Generation) with AST-based code context retrieval and strict patch validation to synthesize and verify patches (Kim et al., 2025).
+On the VulnLoc (Shen et al., 2021) benchmark, SAN2PATCH with gpt-4o correctly repairs 31 of 39 vulnerabilities, achieving a 79.5% success rate.
+For AAG-in-the-loop vulnerability repair, the key insight is that careful prompt structuring and multi-stage LLM reasoning over lightweight runtime artifacts like sanitizer logs can replace heavy program analysis and manual localization while still producing functionally correct patches at scale.
 PredicateFix targets security alerts raised by CodeQL (De Moor et al., 2007) and GoInsight.
 Their analyzer inspects the rule predicates behind an alert, derives a bridging predicate that links the defect to missing sanitization, and then retrieves key examples that satisfy that predicate (Xiao et al., 2025).
-These examples, along with the original alert, are incorporated into a retrieval-augmented prompt, guiding GPT-4 or other models to generate a patch.
+These examples, along with the original alert, are incorporated into a retrieval-augmented prompt that guides GPT-4 or other models to generate a patch.
 After each proposal, the analyzer reruns; unresolved alerts trigger another round while already-fixed code is frozen.
@@
 Future work should further cut overhead by reusing partial analysis results.
-4.5. Discussion and Outlook
-Procedural workflows provide predictable cost and reproducible behavior because the script predetermines every retrieval step, test execution, and human intervention.
-Results from ChatRepair, ThinkRepair, Agentless, and KGCompass indicate that repair accuracy rises as the feedback given to a frozen model becomes more precise; structured traces, knowledge-graph context, and reviewer comments offer clearer guidance than raw assertion text.
+Overview Procedural workflows provide predictable cost and reproducible behavior because the script predetermines every retrieval step, test execution, and human intervention.
+Results from ChatRepair, ThinkRepair, Agentless, and KGCompass indicate that repair accuracy increases as the feedback provided to a frozen model becomes more precise; structured traces, knowledge graph context, and reviewer comments offer clearer guidance than raw assertion text.
 Future work can improve efficiency by prioritizing feedback according to its information value, so that prompts contain only the most helpful facts, and by allowing the loop to increase its analytical depth only when simpler cues are insufficient.
@@
 Within this category, autonomy rises along three tiers.
-Tool-augmented agents place one model at the center and let LLMs call repository utilities, build or test commands, and even perform graphical actions, while the outer loop stays static (Yang et al., 2024a, b; Wang et al., 2024; Bouzenia et al., 2024; Zhang et al., 2024b; Luo et al., 2025; Liu et al., 2025).
+Tool-augmented agents place one model at the center and let LLMs call repository utilities, build or test commands, and even perform graphical actions, while the outer loop stays static (Yang et al., 2024a, [n.
+d.]; Wang et al., 2024; Bouzenia et al., 2025; Zhang et al., 2024b; Luo et al., 2025a; Liu et al., 2025).
 LLM-as-judges detach generation from evaluation by adding a critic LLM that scores or ranks patches, guiding an external search without executing code (He et al., 2025; Zhou et al., 2024c; Li et al., 2024b; Yadavally et al., 2025).
-Self-controlled systems hand one or more LLMs full authority over planning, delegation, and termination, enabling them to tackle function-level or multi-file bugs at the cost of heavier coordination (Tao et al., 2024; Antoniades et al., 2024; Örwall, 2024; Su et al., 2025).
+Self-controlled systems hand one or more LLMs full authority over planning, delegation, and termination, enabling them to tackle function-level or multi-file bugs at the cost of heavier coordination (Tao et al., 2024; Antoniades et al., [n.
+d.]; Örwall, 2024; Hongjin et al., 2025).
 The remainder of this section examines these three paradigms in depth.
@@
 Tool-augmented agents place a frozen LLM at the center of a fixed perceive-think-act loop and expose a curated toolbox to that loop.
-The skeleton which “observe the state, reason, select a tool, apply it, repeat” is handwritten and never reordered at run time; the LLMs are limited to deciding which provided tool to call or whether to stop.
-Systems that follow this pattern include SWE-Agent (Yang et al., 2024a), SWE-Agent M (Yang et al., 2024b), OpenHands (Wang et al., 2024), RepairAgent (Bouzenia et al., 2024), AutoCodeRover (Zhang et al., 2024b), LANTERN (Luo et al., 2025), and VulDebugger (Liu et al., 2025), all of which improve over one-shot prompting by letting LLMs choose among grep, compilation, testing, translation, or debugging actions while the outer control flow stays static.
+The skeleton, which “observe the state, reason, select a tool, apply it, repeat” is handwritten and never reordered at run time; the LLMs are limited to deciding which provided tool to call or whether to stop.
+Systems that follow this pattern include SWE-Agent (Yang et al., 2024a), SWE-Agent M (Yang et al., [n.
+d.]), OpenHands (Wang et al., 2024), RepairAgent (Bouzenia et al., 2025), AutoCodeRover (Zhang et al., 2024b), LANTERN (Luo et al., 2025a), and VulDebugger (Liu et al., 2025), all of which improve over one-shot prompting by letting LLMs choose among grep, compilation, testing, translation, or debugging actions while the outer control flow stays static.
 SWE-Agent (Yang et al., 2024a) uses GPT-4 Turbo in a ReAct (Yao et al., 2023) loop that interleaves git_grep, targeted file edits, and unit-test runs to drive repair without any predefined call order.
@@
 Moreover, SWE-Agent cannot invent brand-new actions at run time; it can only select from the predefined toolkit registered by the authors.
-SWE-Agent M extends SWE-Agent by adding a build_and_screenshot tool that compiles the project, launches a headless browser, captures the rendered UI, and feeds the image to GPT-4V for inspection (Yang et al., 2024b).
+SWE-Agent M extends SWE-Agent by adding a build_and_screenshot tool that compiles the project, launches a headless browser, captures the rendered UI, and feeds the image to GPT-4V for inspection (Yang et al., [n.
+d.]).
 On their built SWE-bench Multimodal benchmark, it increases the fix rate by 2.3 percentage points over the original agent while maintaining the same reasoning loop, indicating that lightweight visual cues can help resolve issues rooted in front-end behavior without redesigning the core architecture.
@@
 Effectiveness drops when tests are weak or the project lacks a clear modular structure.
-RepairAgent wraps an LLM inside a finite-state controller that exposes tools for reading files, running tests, compiling code, and editing sources (Bouzenia et al., 2024).
+RepairAgent wraps an LLM inside a finite-state controller that exposes tools for reading files, running tests, compiling code, and editing sources (Bouzenia et al., 2025).
 The model chooses actions, observes outputs, and loops until tests pass or progress stalls, reaching 164 fixes on the Defects4J v1.2/2.0 benchmark.
@@
 Long exploratory sessions consume many tokens, and careful guardrails are required to avoid infinite loops.
-LANTERN tackles low-resource languages by translating buggy code into a language where the LLM is stronger, repairing it there, and translating the patch back (Luo et al., 2025).
+LANTERN tackles low-resource languages by translating buggy code into a language where the LLM is stronger, repairing it there, and translating the patch back (Luo et al., 2025a).
 A decision module picks the intermediate language, and iterative testing guides further rounds of translation and refinement.
@@
 In summary, tool-augmented agents demonstrate that granting an LLM selective tool access significantly enhances fix rates compared to procedural scripts.
-However, every additional reasoning step and external call still follows the same static loop.
+However, every additional reasoning step and external call still follow the same static loop.
 Future work should consider caching tool outputs, incorporating lightweight heuristics, or adaptively adjusting the reasoning budget to maintain flexibility while reducing the efficiency gap with purely scripted pipelines.
@@
 (He et al., 2025), positions them as scalable proxies for human reviewers when assessing whether a change fixes a defect or eliminates a vulnerability.
-Concrete judge modules now appear in 4 emerging lines of work: differential patch testing for benchmark validation (Wang et al., 2025), automatic patch correctness assessment (Zhou et al., 2024c), security-fix isolation (Li et al., 2024b), and reference-based execution-free critique (Yadavally et al., 2025).
-Each line integrates an LLM critic into an APR pipeline to filter, rank, or label patches.
-LLM4PatchCorrect uses in-context learning with an LLM for code to judge whether a candidate patch is correct by jointly processing bug descriptions, failing tests, execution traces, test coverage, and previously labeled patches (Zhou et al., 2024c).
-Using StarCoder as the backbone and evaluated on public automatic patch correctness assessment (APCA) datasets (Defects4J (Just et al., 2014) and BEARS (Madeiral et al., 2019)), it achieved 84.4% accuracy and an 86.5% F1 score without requiring any labels from the target repair tool.
-In this way, LLM4PatchCorrect can reduce the manual labeling effort when building a model to automatically assess the correctness of generated patches of new APR tools.
-Both the training and test bugs originate from a limited set of projects, so the tool may not generalize well to other software.
-VulSifter identifies the lines in a security-related commit that remove a vulnerability while ignoring cosmetic edits (Li et al., 2024b).
-It feeds the full diff and vulnerability description into an LLM, then applies handcrafted heuristics to discard incidental changes before accepting the patch.
-After scanning 5.3 million commits from 127 thousand projects, the workflow isolates 11,632 verified security fixes, and models trained on this purified data generalize better than those trained on raw commits.
-The LLM-rule hybrid yields high precision, but because it makes decisions after the commit, VulSifter cannot issue real-time alerts during an active repair loop.
-Execution-Free Critic supplies instant feedback by comparing a candidate patch with a known ground-truth fix without running code or tests (Yadavally et al., 2025).
-The critic LLM inspects each modified region, reasons about type safety and semantics, and predicts whether the edited program will compile and behave as the reference solution.
-On a repository-level benchmark, it matched actual test outcomes in 84.8% of cases and achieved a token-level judgment F1 of 0.916, surpassing static metrics and prior heuristics by up to 72%.
-Its first advantage is speed: removing build and test execution slashes evaluation latency in multi-step repair loops.
-Second, token-level commentary highlights exactly which parts of the candidate diverge from the oracle, giving search algorithms fine-grained guidance.
-The method, however, presupposes access to a correct patch, so it mainly fits evaluation or teacher-student training scenarios rather than real-world debugging.
-Performance also drops when large, cross-file patches overflow the model context, exposing the usual length limits of transformer critics.
-PatchDiff (Wang et al., 2025) pioneers call-trace-guided differential patch testing: it identifies functions touched by a candidate patch, prunes surrounding context, and prompts GPT-4o-mini to generate unit tests that pass under the oracle patch but fail under the candidate, thereby acting as an automated referee.
-Evaluated on the 500-task SWE-bench Verified benchmark and three leading agents (CodeStory (CodeStoryAI contributors, 2025), LearnByInteract (Su et al., 2025), and OpenHands (Wang et al., 2024)), running all developer tests already invalidates 7.8% of “plausible” patches.
-In contrast, PatchDiff’s tests expose behavioral divergence in another 29.6%, and manual inspection confirms 28.6% of those as incorrect, inflating leaderboard scores by approximately 6.4 percentage points.
-The study reveals that most false positives originate from weak test suites and subtle “similar-yet-divergent” implementations, indicating that LLM-driven differential testing can rapidly identify semantic errors without requiring the full execution of the regression suite.
+In our taxonomy, such judge components fall under the agentic paradigm whenever their scores gate which candidates or search branches are pursued next or when to stop searching.
+Concrete judge modules now appear in 4 emerging lines of work: differential patch testing for benchmark validation (Wang et al., 2025), tree-search-guided patch evaluation in iterative repair loops (Hu et al., 2025), dual abstention-and-validation funnels (Cambronero et al., 2025), and patch validation (Ruan et al., 2025).
+TSAPR instantiates a tree-search APR framework in which an LLM collaborates with Monte Carlo Tree Search to iteratively select, refine, and evaluate candidate patches instead of following a single-path trial-and-error loop (Hu et al., 2025).
+In this framework, the LLM plays a dual role as generator and judge: it produces new patches with chain-of-thought reasoning, then reuses the same model as an LLM-as-Judge to score failing candidates based on buggy context, test results, execution feedback, and self-reflection traces, while falling back to a Test-as-Judge policy when test suites provide dense coverage.
+Empirically, TSAPR repairs 201 out of 835 bugs on Defects4J (Just et al., 2014) under a pass@16 budget and 27 out of 79 vulnerabilities on VUL4J (Bui et al., 2022), surpassing prior APR and AVR baselines under substantially smaller patch budgets.
+These results suggest that combining global search with self-evaluating LLM critics can turn sparse test outcomes into informative rewards, strengthening LLM-as-Judge for complex multi-location and security-sensitive repairs.
+Abstain and Validate introduces a dual LLM policy layer on top of an industrial agentic APR system that learns when to attempt a repair and when to surface a candidate patch to developers, aiming to reduce noise and wasted compute at the repository level during program repair (Cambronero et al., 2025).
+As LLM-as-Judge components, the bug abstention model reads the bug report and predicts the success probability of the underlying agent, deciding to skip low probability bugs.
+In contrast, the patch validation model generates a textual fix specification and then judges each candidate patch and its trajectory, assigning a correctness score that filters patches even after build and regression tests pass.
+On 2 additional sets of 198 null pointer exceptions and 50 sanitizer bugs from Google’s codebase, the dual policies raise filtered fail-to-pass@1 on human bugs from 11.29% to 53% and improve filtered accept@1 by up to 24pp on NPE bugs.
+These results show that wrapping an agentic repair loop with front-end abstention and back-end LLM-as-Judge validation can sharply increase the precision and perceived reliability of APR in production.
+SpecRover builds on AutoCodeRover by coupling iterative specification inference with a multi-agent LLM workflow that searches the codebase, synthesizes function-level intent summaries, and uses them to drive patch generation for GitHub issues on SWE-bench (Ruan et al., 2025).
+In its review stage, a dedicated LLM reviewer agent acts as an LLM-as-Judge, executing the reproducer test on the original and patched programs, inspecting inferred specifications, tests, and issue descriptions, and then either rejecting the patch with natural-language feedback or approving it and triggering regression testing.
+Empirically, SpecRover solves 19.3% of issues on the full SWE-bench and 31.0% on SWE-bench Lite at an average cost of \$0.65 per issue.
+These results suggest that positioning an LLM as a high-level reviewer that reconciles specifications, tests, and natural language intent can substantially strengthen LLM-as-Judge, turning it into a calibrated confidence filter and explanation generator rather than a standalone oracle.
+LLM-as-Judge components recast patch evaluation as a learned decision problem, letting repair systems trade broader exploration for higher precision at acceptance time.
+Across tree search, abstention and validation policies, and specification-guided review, they already improve solve rates, suppress noisy patches that merely satisfy weak tests, and provide textual justifications that help developers understand decisions.
+Remaining challenges include calibration, robustness to spurious cues, and alignment with stronger oracles, so future work should more closely couple judges with static analysis and regression testing, learn explicit abstention behavior, and treat these critics as conservative gatekeepers rather than standalone authorities.
 7.3. Self-Controlled System
@@
 One or more LLMs generate or revise the global plan at run time, spawn new sub-tasks when needed, and decide when to terminate, so the model, not the script, controls the entire outer loop.
-Representative examples are MAGIS (Tao et al., 2024), which distributes work to four specialized roles; SWE-Search (Antoniades et al., 2024; Örwall, 2024), which couples LLMs with Monte Carlo tree search to debate and apply patches; and Learn-by-Interact (Su et al., 2025), which lets a single model explore an environment, refine its strategy, and distill the trajectory into reusable instructions.
+Representative examples are MAGIS (Tao et al., 2024), which distributes work to four specialized roles; SWE-Search (Antoniades et al., [n.
+d.]; Örwall, 2024), which couples LLMs with Monte Carlo tree search to debate and apply patches; and Learn-by-Interact (Hongjin et al., 2025), which lets a single model explore an environment, refine its strategy, and distill the trajectory into reusable instructions.
 With full control, these agents can tackle multi-file or cross-module defects but incur higher coordination costs and safety risks.
@@
 The price is heavy coordination overhead and modest absolute success, leaving many hard bugs unresolved.
-SWE-Search combines Monte Carlo Tree Search with a multi-agent control loop in which GPT-4o iteratively searches the repository, plans edits, and debates candidate patches before applying a fix (Antoniades et al., 2024; Örwall, 2024).
+SWE-Search combines Monte Carlo Tree Search with a multi-agent control loop in which GPT-4o iteratively searches the repository, plans edits, and debates candidate patches before applying a fix (Antoniades et al., [n.
+d.]; Örwall, 2024).
 On SWE-bench Lite, the system raises the solve rate to 31% with GPT-4o and delivers a 23% relative gain across five LLMs, while the Claude 3.5 Sonnet achieves 39% fixes at 2.7 issues per dollar.
 However, the search depth scales with inference-time computation and has been noted as “probably pretty expensive,” pointing to future work on cost-aware depth control and quicker repository traversal.
-Learn-by-Interact is a self-controlled agentic framework that lets an LLM explore a new software or web environment, plan its action sequence, execute those actions, and then retrospectively distill the trajectory into reusable instructions—so the model, not a script, decides every step of the outer loop, satisfying the 5.3 criterion for full autonomy (Su et al., 2025).
-They fine-tune a Codestral-22B backbone on the synthesized trajectories and also demonstrate a training-free variant in which Claude-3.5 Sonnet retrieves the distilled instructions at run time to guide its decisions.
-On SWE-bench and other benchmarks, Learn-by-Interact improves in-context learning accuracy by up to 12.2% and supervised fine-tuning performance by 19.5% over best baselines.
-However, it relies on detailed and accurate environmental documentation; when documentation is sparse or rapidly changing, the data-synthesis phase can yield low-quality trajectories and incur substantial computational overhead.
-Self-controlled agents, therefore, deliver the highest flexibility and already outperform simpler agents on public benchmarks.
-Three lessons emerge.
+Learn-by-Interact is a self-controlled agentic framework that lets an LLM explore a new software or web environment, plan its action sequence, execute those actions, and retrospectively distill the trajectory into reusable instructions, so the LLM, rather than a fixed script, decides every step of the outer loop (Hongjin et al., 2025).
+The authors fine-tune a Codestral-22B backbone on synthesized trajectories and also provide a training-free variant where Claude-3.5 Sonnet retrieves the distilled instructions at run time.
+On SWE-bench, Claude-3.5 Sonnet attains 60.0% pass@1 and improves pass@1 by up to 12.2pp over the baseline.
+The approach depends on accurate and up-to-date environment documentation; when documentation is sparse or changing, the data-synthesis phase can yield low-quality trajectories and incur substantial computational overhead.
+Self-controlled agents outperform simpler agents on public benchmarks and emerge with three lessons.
 First, explicit role separation or self-generated sub-goals narrows the search space without sacrificing autonomy.
-Second, stable feedback signals such as full test suites or critical debates prevent blind exploration.
-Third, careful cost management remains crucial because token budgets, tool latency, and coordination overhead can offset accuracy gains.
+Second, stable feedback signals, such as full test suites or critical debates, prevent blind exploration.
+Third, careful cost management remains crucial because token budgets, tool latency, and coordination overhead can offset gains in accuracy.
 Continued progress on adaptive stopping rules, incremental context caching, and lightweight self-reflection is essential before these fully autonomous agents can meet the throughput and reliability demands of continuous integration.
-5.4. Discussion and Outlook
-Future research in agentic repair should emphasize cost-aware planning that caches intermediate results, prunes unproductive branches, and calls judgment modules only when their expected benefit justifies the extra computational resources, thereby reducing latency.
+Overview Future research in agentic repair should emphasize cost-aware planning that caches intermediate results, prunes unproductive branches, and calls judgment modules only when their expected benefit justifies the extra computational resources, thereby reducing latency.
 Robust context management is also essential: agents must retain only the most relevant code fragments, tests, and analysis artifacts so that large repositories remain within the model’s context window without losing critical detail.
@@
 Sustained attention to these directions can elevate current prototypes into dependable tools for everyday software engineering practice.
-Table 3. Detailed Information of LLM-based APR Systems
+Having surveyed fine-tuning, prompting, procedural pipelines, and agentic frameworks, we now synthesize how the control paradigms and augmentation layers jointly organize the design space and answer RQ2.
+Figure 6. Alignment between LLM-based repair paradigms and typical deployment scenarios.
+Answer to RQ2 The LLM-based software repair design space is organized into four control paradigms with retrieval and analysis augmentation as cross-cutting layers.
+Fine-tuning aligns a model to recurring repair patterns and works well for single-hunk or domain-specific defects.
+Prompting enables rapid fixes for single-function bugs without training, but is sensitive to prompt design and context limits.
+Procedural pipelines script generates and tests or retrieves and analyzes loops to deliver deterministic behavior on harder tasks.
+Agentic frameworks plan, call tools, and coordinate multi-step changes, which extend to multi-hunk and cross-file repository-level repairs.
+Figure 6 further maps these paradigms to representative deployment scenarios, including educational assignments, function-level benchmarks, repository-level issue suites, security vulnerabilities, and industrial CI workflows, indicating where each paradigm tends to be a primary or secondary fit.
+Mapping the representation works to these cells makes the control authority, iteration budget, and external-tool reliance explicit and ties each cell to the defect scopes it is best suited to handle.
+8. Trends in Benchmark Repair Success Rates
+This section consolidates published success rates for widely used software repair benchmarks to illustrate how performance evolves across paradigms.
+We report each system’s headline metric as initially published and record the evaluation assumptions alongside the score, rather than simply providing a single result.
+This maintains fidelity to the original evaluation setups and makes explicit differences in fault localization, scope, and pass@$k$.
+We cover four widely-used benchmarks: Defects4J (v1.2 and v2.0), HumanEval-Java, SWE-bench Verified, and SWE-bench Lite.
+For each benchmark, we select representative systems that cover four paradigms and list their published scores along with the original metric and assumptions.
+We use pass@k to denote the number of candidate patches for which at least one of k independent samples yields a patch that passes the test suites.
+When a wall-clock or token budget is enforced, we state it explicitly, for example, pass@10 within 5 hours.
+Table 5. Defects4J trend snapshot.
+Year Paradigm System Metric Notes
+2022 Prompting (zero-shot) AlphaRepair top@$k$ within 5 hours: 74/109 Perfect fault localization.
+2023 Prompting (few-shot) Xia et al. pass@200: 99/255 Perfect fault localization.
+2023 Full fine-tuning Huang et al. pass@5: 82/107 Perfect fault localization.
+2023 Full fine-tuning Jiang et al. pass@10: 31/138 Perfect fault localization.
+2023 AAG-in-the-loop Repilot pass@5000: 116/273 Single-hunk Java bugs.
+2023 Knowledge distillation KNOD pass@1000: 122/837 Perfect fault localization.
+2024 AAG-enhanced prompting D4C pass@10: 180/437 Single-function setting, no FL hints.
+2024 Test-in-the-loop ChatRepair pass@500: 162/337 Single-function setting.
+2024 Tool-augmented agent RepairAgent pass@1: 164/835 Repository-level execution with tools.
+2024 Distillation (non-AR student) NARRepair pass@1: 110/127 Perfect fault localization.
+2025 LLM-as-Judges TSAPR pass@16: 201/835 Perfect fault localization.
+On Defects4J (Table 5), the field evolves from early single-turn prompting with perfect fault localization toward analysis-augmented prompting, distillation, and procedural loops that reuse failing assertions or traces.
+Solved-bug counts rise substantially even as some assumptions are relaxed, for example moving from single-hunk Java snippets to single-function or repository-level settings.
+Table 6. HumanEval-Java trend snapshot.
+Year Paradigm System Metric Notes
+2023 PEFT RepairLLaMA pass@10: 67.28% Perfect fault localization.
+2024 PEFT Li et al. pass@10: 68.10% Perfect fault localization.
+2024 PEFT (federated) Luo et al. pass@10: 76.88% Privacy-preserving training.
+2025 PEFT (iterative) Ruiz et al. pass@10: 78.53% Perfect fault localization.
+2025 PEFT (multi-objective) MORepair pass@10: 77.90% Rationale-aware training.
+2024 Context-enriched FT NTR pass@100: 83.44% Template guidance, perfect FL.
+2024 Test-in-the-loop ContrastRepair pass@40: 84.05% Contrastive test pairs.
+2025 Analysis-augmented prompting TracePrompt accuracy: 71.30% Concise execution traces.
+On HumanEval-Java (Table 6), parameter-efficient fine-tuning already delivers pass@10 near eighty percent.
+Context- and template-enriched fine-tuning, analysis-augmented prompts with concise traces, and conversational test-in-the-loop loops push headline scores higher at larger k.
+These results suggest that most low-hanging fruit on this benchmark now lies in better context selection and decoding strategies rather than in further scaling model size.
+Table 7. SWE-bench trend snapshot. Verified and Lite variants are reported separately.
+Year Variant Paradigm System Metric
+2024 Lite Tool-augmented agent SWE-Agent pass@1: 18.00%
+2024 Lite Tool-augmented agent OpenHands pass@1: 26.00%
+2024 Lite Tool-augmented agent AutoCodeRover pass@1: 19.00%
+2024 Lite RAG-in-the-loop Agentless pass@1: 32.00%
+2024 Lite Self-controlled agent SWE-Search pass@1: 39.00%
+2025 Lite RAG-in-the-loop KGCompass pass@1: 46.00%
+2025 Lite LLM-as-Judges TSAPR pass@32: 54.67%
+2025 Lite LLM-as-Judges SpecRover pass@1: 31.00%
+2024 Verified Human-in-the-loop HULA pass@1: 31.00%
+2025 Verified RLFT + procedural SWE-RL pass@1: 41.00%
+On SWE-bench (Table 7), repository-level repair benefits most from richer control and context.
+Tool-augmented agents achieve the first gains on SWE-bench Lite, but RAG-in-the-loop pipelines and knowledge-graph-guided retrieval drive pass@1 into the mid-forties by supplying focused repository context between generations.
+On SWE-bench Verified, human-in-the-loop pipelines and reinforcement-learning-enhanced procedural workflows raise success rates.
+Overall, progress on SWE-bench is closely tied to better repository modeling and feedback rather than to single-step prompting.
+Heterogeneous metrics and sampling budgets further obscure whether improvements arise from better models, from more generous pass@k, or from stronger evaluation setups.
+Table 7 makes protocol differences explicit, supports more transparent comparison across systems, and highlights where bias is most likely to appear when interpreting published repair success rates.
+Answer to RQ3 Across the benchmarks we survey, reported repair rates are shaped as much by protocol design as by model quality.
+On Defects4J and HumanEval-Java, many systems evaluate on different subsets of bugs and mix pass@$k$ with accuracy-style metrics, and most headline numbers still assume perfect fault localization rather than end-to-end repair.
+On SWE-bench, recent agentic and reinforcement-learning-enhanced systems are converging on repository-level, end-to-end settings with fixed pass@1 budgets, but vulnerability repair remains fragmented across small, non-overlapping datasets, so scores there are substantially harder to compare.
+These observations suggest three concrete priorities for future evaluation.
+First, results should be reported on the full version of each benchmark, or with clearly marked subsets and explicit denominators to make the results comparable.
+Second, studies should standardize metrics, for example, always reporting pass@1 and one shared larger $k$ without other sampling or wall-clock budget, and leveraging end-to-end settings without perfect fault-localization hints.
+Third, the vulnerability-repair community would benefit from a consolidated benchmark in the spirit of SWE-bench, with curated repository-level defects and strong oracles; until such a suite exists, improvements on narrow vulnerability datasets should be treated as evidence of local progress rather than as general advances in LLM-based repair.
+9. Related Work
+Earlier surveys on LLM-based software repair focus on predate LLM-based repair or mention LLMs only briefly, so they do not analyze LLM-centered control structures.
+Huang et al.
+(Huang et al., 2023b) updated the traditional landscape with a four-way taxonomy over search, constraint, template, and learning techniques, but relegated LLMs to a short remark about future potential.
+Zhang et al.
+(Zhang et al., 2023a) presented a thorough synthesis of learning-based software repair, framing the workflow around localization, generation, ranking, and validation; because their literature search concluded before mid-2022, large-scale LLMs such as Codex receive only brief mention, and prompt-driven repair workflows are not yet covered in depth.
+Winter et al.
+(Winter et al., 2022) shifted focus to developer perceptions and usability, without a technical taxonomy.
+Dikici et al.
+(Dikici and Bilgin, 2025) grouped tools into template, machine-learning, and deep-learning categories, implicitly folding GPT-style systems into the deep-learning bin without analyzing their distinctive traits or their impact on the repair loop.
+More recent work concentrates on LLMs and related evaluation issues.
+Anand et al.
+(Anand et al., 2024) explicitly surveyed LLM-based software repair and code generation but covered only a small set of software repair papers and did not propose a dedicated taxonomy.
+Haque et al.
+(Haque et al., 2025a) conducted a systematic literature review on parameter-efficient fine-tuning for LLMs, covering software repair but focusing on fine-tuning methods rather than repair workflows.
+He et al.
+(He et al., 2025) examined the emerging LLM-as-judge paradigm for code evaluation, offering insights into assessing LLM-generated patches while omitting a repair-oriented taxonomy.
+Zhou et al.
+(Zhou et al., 2024a) published a review that catalogs work on vulnerability detection and repair and discusses adaptation techniques and dataset usage, but its analysis is limited to security-related bugs, and it treats retrieval-augmented and agentic methods mainly as future directions.
+Zhang et al.
+conducted a literature review focused on LLM-based software repair (Zhang et al., 2024a), collecting a broader corpus of LLM-based repair studies and classifying them by LLM utilization modes, without explicitly modeling control authority or augmentation layers.
+Zhang et al.
+provide a critical review that uses ChatGPT as a case study on a contamination-resistant benchmark, examining prompt design and data-contamination risks (Zhang et al., 2023b).
+Table 8 summarizes these surveys and their coverage of LLM-based techniques.
+Table 8. Recent surveys on automated program repair and their coverage of LLM-based techniques.
+Survey Venue Year LLM Coverage Classification Scheme Key Limitations
+Winter et al. (Winter et al., 2022) IEEE TSE 2022 None Human-factor perspective Lacks technical classification; predates LLM era.
+Huang et al. (Huang et al., 2023b) arXiv 2023 Mentioned as future work Search, constraint, template, learning No category devoted to LLM repair.
+Zhang et al. (Zhang et al., 2023a) ACM TOSEM 2023 Moderate Pipeline stages Excludes non-ML approaches and prompt-based repair.
+Zhang et al.
+(Zhang et al., 2023b) arXiv 2023 Strong for ChatGPT case ChatGPT-centered critical review and prompt variants Single-model case study; not a general software repair taxonomy or broad LLM-based software repair survey.
+Anand et al. (Anand et al., 2024) arXiv 2024 Strong LLM-based software repair and code generation software repair coverage limited; no fine-grained taxonomy.
+Zhang et al.
+(Zhang et al., 2024a) arXiv 2024 Strong LLM utilization modes and repair scenarios Emphasizes how LLMs are used rather than control paradigms; does not explicitly model retrieval- or analysis-augmented repair; limited cell-wise comparison of evaluation practices.
+Dikici et al. (Dikici and Bilgin, 2025) KIS 2025 Partial Template, ML, deep learning LLM tools subsumed under deep learning.
+Haque et al. (Haque et al., 2025a) arXiv 2025 Moderate Fine-tuning strategy taxonomy Focused on fine-tuning methods, not repair workflows.
+He et al. (He et al., 2025) arXiv 2025 Evaluation only Evaluation role taxonomy Covers judging, not generation or workflow design.
+Zhou et al. (Zhou et al., 2024a) ACM TOSEM 2025 Strong Security-oriented vulnerability repair Focuses only on vulnerabilities; no agentic approaches
+Across these studies, three limitations remain.
+First, most taxonomies either predate LLM-based repair or retain a pipeline- or model-centric view and therefore do not treat prompt-driven, retrieval-augmented, procedural, and agentic workflows as first-class control paradigms tied to who steers the repair loop and whether the base LLM is adapted.
+Second, LLM-based works are often omitted or folded into broad deep-learning bins, which obscures how foundation models reshape software repair and how external tools, retrieval modules, and analysis components are orchestrated around them.
+Third, evaluation challenges specific to LLMs, including data leakage, weak test suites, heterogeneous pass@k settings, and differing fault-localization assumptions, are rarely treated systematically or aligned across benchmarks.
+We complement these surveys by offering a control-paradigm-centered taxonomy defined by parameter adaptation and control authority with explicit retrieval- and analysis-augmentation layers, by mapping 62 recent systems into this design space together with their defect scope and deployment scenario, and by consolidating benchmark, metric, and fault-localization assumptions into comparative tables, supported by a versioned, auditable artifact that exposes screening and coding decisions for future reuse.
+10. Open Challenges
+Despite rapid progress, LLM-based software repair still faces several open problems that span the four paradigms.
+Training and deployment cost.
+Fine-tuning larger LLMs requires many GPUs and long wall-clock hours.
+Even token-efficient adapters still process millions of tokens.
+During inference, procedural and especially agentic workflows frequently call the LLM, invoke external tools, and rerun tests.
+Each round consumes a significant amount of tokens; running the loop locally demands high-memory GPUs, while sending it to a commercial API incurs substantial token charges.
+These long, costly sessions make real-world deployment of repair systems challenging.
+Evaluation reliability.
+Benchmarks may overlap with pretraining data, which causes data leakage and inflated success rates.
+Weak test suites let a patch pass if all test succeeds, and underspecified test suites can miss semantic faults of LLM-generated patches.
+This leads to patch overfitting: the LLM produces a plausible patch that passes every available test but still contains hidden defects.
+Stricter suites, extra hidden tests, and leakage-excluded datasets frequently reveal accuracy drops.
+Some studies attempt to compensate by having experts manually judge patch correctness; however, manual review introduces reviewer bias and does not scale due to the cost of human verification.
+Workflow robustness.
+Minor changes to prompt wording, sampling temperature, or random seeds can alter the generated patch and turn a successful repair into a failure.
+In agentic workflows, this randomness accumulates across steps, so two runs of the same system can follow different tool sequences and yield different results.
+Language models may also produce invalid commands, ignore provided rules, or loop for too long, making the total running time unpredictable.
+These issues make it hard to guarantee repeatable outcomes or keep running costs within fixed limits in real-world pipelines.
+Retrieval and context management.
+Most RAG pipelines still rely on embedding-based similarity search, which often misses some of the key files or functions and gives no recall guarantee.
+Even when relevant snippets are found, the added context can push the prompt beyond the LLM’s context limitation, limiting the effectiveness of RAG.
+Recent adaptive memory systems, such as A-MEM, store each interaction as a structured note, maintain a link graph, and supply only a compact top-$k$ context, thereby improving recall while staying within the token budget (Xu et al., 2025b).
+AAG pipelines generate diagnostics and traces directly from program analysis; however, current LLMs treat this structured data as plain text and fail to meaningfully exploit its logical structure.
+Security and trust.
+LLM-generated patches often change more code than needed: they may pass the available tests yet involve new latent bugs, break style guides, or slow the program.
+Calls to commercial APIs can leak proprietary codes and documents, and agents that bypass static checks may leave vulnerabilities unfixed.
+In addition, LLMs usually provide a natural-language rationale for each patch; if this explanation is incomplete or inconsistent with the actual code change, it can mislead reviewers and hinder debugging.
+A central open problem is how to verify both the patch and its rationale quickly and accurately, ideally before deployment, using lightweight static analysis, selective dynamic checks, and automated rollback when risks are detected.
+Answer to RQ4 Across paradigms, four bottlenecks stand out for semantic correctness beyond tests and for repository-level multi-hunk repair: (i) limited and often shallow test suites, (ii) brittle and expensive multi-step workflows whose time and token usage grow quickly with project size, (iii) retrieval and analysis pipelines that still supply noisy or incomplete context, and (iv) weak assurance mechanisms for the safety and trustworthiness of accepted patches.
+We see three concrete opportunities.
+First, build benchmarks and test suites that explicitly target large-scale, multi-hunk changes, with stronger oracles, hidden tests, and contamination-aware splits, so that progress reflects real semantic correctness rather than overfitting narrow unit suites.
+Second, design cost-aware repair policies that treat success per unit budget as an objective, by reusing intermediate analysis results, capping the number of LLM calls, and falling back to simpler scripted pipelines when a bug appears easy, thereby controlling latency and variability on large repositories.
+Third, separate generation from evaluation more sharply by combining analysis- and test-based modules that gate patch acceptance and rationales, so that final decisions do not rely solely on larger LLMs.
+Table 9. Detailed Information of LLM-based Software Repair Systems
 Sub-Paradigm System Training Dataset Representative Result Assumption
 Base Model Benchmark Metric Result
-Fine-Tuning Full Fine-Tuning Huang (2023) Recoder dataset CodeT5 Defects4J v1.2/2.0 pass@5 82/107 Perfect Fault Localization
-Jiang (2023) Recoder dataset Incoder 7B Defects4J v1.2/2.0 pass@10 31/138 Perfect Fault Localization
+Fine-Tuning
+Full Fine-Tuning
+Huang et al. Recoder dataset CodeT5 Defects4J v1.2/2.0 pass@5 82/107 Perfect Fault Localization
+Jiang et al. Recoder dataset Incoder 7B Defects4J v1.2/2.0 pass@10 31/138 Perfect Fault Localization
 VulMaster CVEFixes, Big-Vul CodeT5 CVEFixes full match 20% train/test split, requires CVE type
 RepairCAT APPS StarCoder-1B APR Competition 2024 pass@1000 14%
-Fine-Tuning PEFT RepairLlama Megadiff CodeLlama 7B HumanEval-Java pass@10 67.28% Perfect Fault Localization
+Fine-Tuning
+PEFT
+RepairLLaMA Megadiff CodeLlama 7B HumanEval-Java pass@10 67.28% Perfect Fault Localization
 MORepair TutorLLMCode CodeLlama 13B HumanEval-Java pass@10 77.90%
-Luo (2024) TutorCode CodeLlama 13B HumanEval-Java pass@10 76.88%
-Li (2024) APR Instruction DeepSeek-Coder 7B HumanEval-Java pass@10 68.10% Perfect Fault Localization
-Ruiz (2025) Recoder dataset DeepSeek-Coder 7B HumanEval-Java pass@10 78.53% Perfect Fault Localization
-Fine-Tuning Knowledge Distilling KNOD GitHub / Defects4J v1.2/2.0 pass@1000 122/837 Perfect Fault Localization
-DistiLRR MBXP CodeLlama-7B HumanEval-Perl pass@1 42.10% train/test split
+Luo et al. TutorCode CodeLlama 13B HumanEval-Java pass@10 76.88%
+Li et al. APR-Instruction DeepSeek-Coder 7B HumanEval-Java pass@10 68.10% Perfect Fault Localization
+Ruiz et al. Recoder dataset DeepSeek-Coder 7B HumanEval-Java pass@10 78.53% Perfect Fault Localization
+Fine-Tuning
+Knowledge Distillation
+KNOD GitHub / Defects4J v1.2/2.0 pass@1000 122/837 Perfect Fault Localization
+DistiLRR MBPP CodeLlama-7B HumanEval-Perl pass@1 42.10% train/test split
 NARRepair selfAPR CodeT5 Defects4J v1.2/2.0 pass@1 110/127 Perfect Fault Localization
-Fine-Tuning RLFT RePair CodeNet4Repair StarCoder-16B CodeNet4Repair pass@5 65.66% train/test split
+Fine-Tuning
+RLFT
+RePair CodeNet4Repair StarCoder-16B CodeNet4Repair pass@5 65.66% train/test split
 SecRepair InstructVul CodeGen2-7B InstructVul F1 82 train/test split
@@
 AdaPatcher ACPR CodeLlama-7B ACPR pass@1 67.57% train/test split
-Smart-LLaMA-DPO built trainset Llama3.1 8B test split F1 83.4 train/test split
-Fine-Tuning Context-Enriched TraceFixer CodeNet CodeT5-small CodeNet pass@10 87%
-InferFix InferredBugs CodeX 12B InferredBugs pass@1 76.80% train/test split
+Vul-R2 PrimeVul Qwen-14B-Instruct-1M PrimeVul pass@10 24.83 train/test split
+Fine-Tuning
+Context-Enriched
+TraceFixer CodeNet CodeT5-small CodeNet pass@10 87%
+InferFix InferredBugs Codex-12B InferredBugs pass@1 76.80% train/test split
 PyTy PyTyDefects CodeT5 PyTyDefects full match 54.40% train/test split
 NTR TRANSFER CodeLlama-70B HumanEval-Java pass@100 83.44% Perfect Fault Localization
-Prompting Zero-Shot AlphaRepair CodeBERT Defects4J v1.2/2.0 top@k (5h) 74/473 Perfect Fault Localization
-Prenner (2022) CodeX-175B Quixbugs pass@1 37/80 Perfect Fault Localization
-Fan (2023) CodeX-175B LMDefects pass@5 46/113
-Tian (2023) GPT-3.5 Turbo LeetCode pass@5 105/120
-Prompting Few-Shot Xia (2023) CodeX 12B Defects4J v1.2 pass@600 99/255 Perfect Fault Localization
-Gao (2023) CodeX-175B B2F small full match 32.25%
-Ahmed (2023) CodeX-175B MODIT pass@30 31.80%
+Prompting
+Zero-Shot
+AlphaRepair CodeBERT Defects4J v1.2/2.0 top@k (5h) 74/109 Perfect Fault Localization
+Prenner et al. Codex-175B Quixbugs pass@1 37/80 Perfect Fault Localization
+Fan et al. Codex-175B LMDefects pass@5 46/113
+Tian et al. GPT-3.5 Turbo LeetCode pass@5 105/120
+Prompting
+Few-Shot
+Xia et al. Codex 12B Defects4J v1.2 pass@200 99/255 Perfect Fault Localization
+Gao et al. Codex-175B B2F small full match 32.25%
+Ahmed et al. Codex-175B MODIT pass@30 31.80%
 CEDAR Codex-175B ATLAS full match 76.55%
-Prompting + RAG T-RAP CodeT5 Big-Fixes pass@10 25.80% Perfect Fault Localization
+Prompting
++ RAG
+Ehsani et al. Llama-3.3 BugsInPy pass@5 79.62% Perfect Fault Localization
 RLCE GPT-4 RepoBugs pass@1 81.45%
 DsRepair DeepSeek-Coder V2 DS-1000 pass@1 29.18%
-Prompting + AAG D4C GPT-4 Defects4J v1.2/2.0 pass@10 180/437 single-function bugs
+Prompting
++ AAG
+D4C GPT-4 Defects4J v1.2/2.0 pass@10 180/437 single-function bugs
 Appatch Claude-3.5 Sonnet Zero-Day F1 36.46
 TracePrompt GPT-4 HumanEval-Java accuracy 71.30%
-Procedural Test-in-the-Loop ChatRepair GPT-3.5 Defects4J v1.2/2.0 pass@500 162/337 single-function bugs
+Procedural
+Test-in-the-Loop
+ChatRepair GPT-3.5 Defects4J v1.2/2.0 pass@500 162/337 single-function bugs
 ThinkRepair GPT-3.5 Defects4J v1.2/2.0 pass@25 98/255 single-function bugs
@@
 ContrastRepair GPT-3.5 HumanEval-Java pass@40 84.05%
-Procedural Human-in-the-Loop CREF GPT-4 TutorCode avg@5 76.60% requires tutor guidance
+Procedural
+Human-in-the-Loop
+CREF GPT-4 TutorCode avg@5 76.60% requires tutor guidance
 HULA GPT-4 SWE-bench Verified pass@1 31% requires human feedback
 DRCodePilot CodeT5+ Flink full match 15.27% requires design-rationale logs
-Procedural RAG-in-the-Loop Agentless GPT-4o SWE-bench Lite pass@1 32%
-RAGFix Llama3 70B HumanEvalFix pass@1 78%
+Procedural
+RAG-in-the-Loop
+Agentless GPT-4o SWE-bench Lite pass@1 32%
+PATCH GPT-4 BFP pass@5 39.81%
 KGCompass Claude-3.5 Sonnet SWE-bench Lite pass@1 46%
-Procedural AAG-in-the-Loop Repilot Incoder 7B Defects4J v1.2/2.0 pass@5000 116/273 single-hunk bugs
-FormHe Gemma 2B Student Submissions pass@1 94.20%
-PredicateFix Claude-3.5 Sonnet CVE-2020-13946 pass@1 81/117
+Procedural
+AAG-in-the-Loop
+Repilot Incoder 7B Defects4J v1.2/2.0 pass@5000 116/273 single-hunk bugs
+SAN2PATCH GPT-4o VulnLoc pass@5 79.49%
+PredicateFix Claude-3.5 Sonnet CVE pass@1 81/117
 LLM4CVE Llama3 70B CVEFixes CodeBLEU +20%
-Agentic Tool-Augmented SWE-Agent GPT-4 Turbo SWE-bench Lite pass@1 18.00%
+Agentic
+Tool-Augmented
+SWE-Agent GPT-4 Turbo SWE-bench Lite pass@1 18.00%
 SWE-Agent M GPT-4o SWE-bench Multimodal pass@1 12.20%
@@
 VulDebugger GPT-4 50 of ExtractFix and ARVO pass@1 60%
-Agentic LLM-as-Judges VulSifter GPT-4 CleanVul F1 82%
-LLM4PatchCorrect StarCoder-7B Defects4J v1.2/2.0 AUC 76.9
-Execution-free Critic Claude-3 Opus SWE-bench F1 91.60%
-PatchDiff GPT-4o-mini SWE-bench-Verified accuracy 28.60%
-Agentic Self-Controlled System MAGIS GPT-4 SWE-bench Lite pass@1 16.67%
+Agentic
+LLM-as-Judges
+Abstain and Validate Gemini 2.0 Flash Google accept@1 62%
+TSAPR GPT-3.5 Defects4J v1.2/2.0 pass@16 24.07%
+SpecRover Claude-3.5 Sonnet SWE-bench Lite pass@1 31.00%
+Agentic
+Self-Controlled System
+MAGIS GPT-4 SWE-bench Lite pass@1 16.67%
 SWE-Search Claude 3.5 Sonnet SWE-bench Lite pass@1 39.00%
 Learn-by-Interact Claude-3.5 Sonnet SWE-bench pass@1 60.00%
-6. Open Challenges
-Despite rapid progress, LLM-based APR still faces several open problems that cover the four paradigms.
-Training and deployment cost.
-Fine-tuning large checkpoints needs many GPUs, long wall-clock hours, and sizable energy budgets.
-Even token-efficient adapters must still process millions of tokens.
-During inference, procedural and especially agentic workflows frequently call the LLM, invoke external tools, and rerun tests.
-Each round consumes thousands of tokens; running the loop locally demands high-memory GPUs, while sending it to a commercial LLM incurs substantial token charges.
-These long, costly sessions often exceed the few-minute latency and budget limits typical of continuous-integration pipelines, which makes real-world deployment challenging.
-Evaluation reliability.
-Benchmarks often overlap with pre-training data, which causes data leakage and inflated scores.
-Weak oracles let a patch pass if any test succeeds, and underspecified test suites can miss semantic faults.
-This leads to patch overfitting: the model produces a plausible patch that passes every available test but still contains hidden errors.
-Stricter suites, extra hidden tests, and leakage-excluded datasets frequently reveal large accuracy drops.
-Some studies attempt to compensate by having experts manually judge patch correctness; however, manual review introduces reviewer bias and does not scale due to the cost of human verification.
-Workflow robustness.
-The repair process is highly sensitive to multiple factors.
-Minor changes in prompt wording, sampling temperature, or random seeds can change the generated patch and may turn a successful repair into a failure.
-In agentic workflows, this randomness accumulates across steps, so two runs of the same system can follow different tool sequences and yield different results.
-Language models may also produce invalid commands, ignore provided rules, or loop too long, which makes the total running time unpredictable.
-These issues make it hard to guarantee repeatable outcomes or keep running costs within fixed limits in real-world pipelines.
-Retrieval and context management.
-Most RAG pipelines still rely on embedding-based similarity search, which often misses some of the key files or functions and gives no recall guarantee.
-Even when relevant snippets are found, the added context can push the prompt beyond the LLM’s context limitation, limiting the effectiveness of RAG.
-Recent adaptive-memory systems, A-MEM, store each interaction as a structured note, maintain a link graph, and supply only a compact top-$k$ context, thereby improving recall while staying within the token budget (Xu et al., 2025).
-AAG pipelines generate diagnostics and traces directly from program analysis; however, current LLMs treat this structured data as plain text and fail to exploit its logical structure in a meaningful way.
-Security and trust.
-LLM-generated patches often change more code than needed: they may pass the available tests yet hide latent bugs, break style guides, or slow the program.
-Calls to commercial APIs can leak proprietary code, and agents that bypass static checks may leave vulnerabilities unfixed.
-In addition, LLMs usually provide a natural-language rationale for each patch; if this explanation is incomplete or inconsistent with the actual code change, it can mislead reviewers and hinder debugging.
-A central open problem is how to verify both the patch and its rationale quickly and accurately, ideally before deployment, using lightweight static analysis, selective dynamic checks, and automated rollback when risks are detected.
+11. Threats to Validity
+11.1. Internal validity
+Study identification, selection, and coding threaten internal validity; we mitigate these threats with a structured pipeline.
+We search multiple digital libraries with calibrated queries and citation chasing, apply explicit inclusion and exclusion criteria, and use a unified filtering process reviewed by a coauthor with extensive program repair experience to ensure consistent treatment of borderline cases.
+All search scripts and results are released in our replication package, enabling others to inspect our decisions and recompute statistics with the same notebooks.
+11.2. External validity
+Narrow corpora and artificial evaluation settings threaten external validity.
+Representation intentionally focuses on work published in tier-1 SE, AI, and security venues, and on studies that report quantitative results on large-scale benchmarks.
+As a result, tools evaluated only on proprietary data fall outside our scope.
+We report headline scores as published initially and do not normalize pass@k or budgets across studies.
+We also do not perform a joint reevaluation.
+This preserves faithfulness to prior work but limits cross-paper comparability.
+To mitigate this limitation, we annotate benchmark scope, metric definitions, oracle strength, and fault localization assumptions alongside each reported score to reduce misleading.
+11.3. Construct validity
+Ambiguous definitions and measurements threaten construct validity.
+We mitigate these issues by operationalizing constructs precisely and enforcing a stable coding scheme.
+We define the four repair paradigms and cross-cutting augmentation layers by workflow control and accessible information, apply them consistently across all studies, and discuss borderline cases to reach a shared decision.
 12. Conclusion
-LLM-based APR has progressed from early prototypes to a diverse ecosystem that includes fine-tuned models, prompt-centric methods, scripted pipelines, and autonomous agents.
-Each paradigm balances accuracy, cost, and engineering effort in a distinct manner, and no single approach prevails across all repair scenarios.
-Future gains are more likely to come from disciplined evaluation, smarter context selection, deeper integration with existing developer workflows, and stronger safeguards that keep patches correct, secure, and maintainable.
-Comprehensive Overview.
-Table 3 consolidates 63 representative LLM-based APR systems, aligning each with its paradigm, training corpus, base model, benchmark, and headline result.
-This overview enables readers to see how design choices impact repair performance easily and sets the stage for the forward-looking discussion that follows.
-LLM-based APR: A Turning Point.
-Program repair has moved beyond single-shot prompts that fix simple bugs.
-Modern agentic systems can scan real repositories, interpret failing tests, and submit production-ready patches.
-These systems already significantly surpass traditional techniques on established benchmarks.
-They also open the door to repairing cross-file, multilingual, and security-critical defects that prior methods rarely addressed.
-LLM-based APR: A Path Forward.
-Turning promising prototypes into dependable development infrastructure will require several advances.
-Benchmarks must be free of data leakage and patch over-fitting; training pipelines must respect practical resource limits; and retrieval engines must reliably surface the most relevant context.
-Agentic planners need to reason efficiently and stay within predictable time and money costs.
-Meeting these challenges will elevate LLM-based APR from an emerging research topic to a routine part of modern software engineering.
+LLM-based software repair has evolved from early single-prompt experiments to a set of distinct paradigms that combine different forms of control over the repair loop with different degrees of model adaptation.
+This survey provides a structured view of the field by proposing a taxonomy based on control authority and parameter adaptation, with retrieval and analysis augmentation as orthogonal layers; mapping 62 representative systems into this framework; and consolidating their benchmarks, evaluation assumptions, and pass@k definitions into a common set of tables.
+Taken together, the taxonomy, corpus, and evaluation summary clarify the tradeoffs among fine-tuning, prompting, procedural pipelines, and agentic frameworks, and show how design choices and evaluation protocols jointly shape reported repair success.
+The synthesis points to the most important directions for future work: stronger, less contaminated benchmarks; more transparent, comparable reporting of pass@k and resource budgets; more selective, repository-aware retrieval and analysis pipelines; and cost-aware planning for multi-step workflows.
+Addressing these issues is central to turning LLM-based repair from promising prototypes into dependable infrastructure in everyday software engineering practice.